### v1

# Comparison Principles for Stochastic Volterra equations

Abstract.
In this work, we establish a comparison principle for stochastic Volterra equations with respect to the initial condition and the drift $b$ applicable to a wide class of Volterra kernels and driving forces $g$ that may be singular in zero. For completely monotone Volterra kernels such a result holds without any further restrictions, while for not completely monotone kernels it is shown that such a principle fails unless the drift is additionally monotone. As a side-product of our results, we also complement the literature on the weak existence of continuous nonnegative solutions. This covers the rough Cox-Ingersoll-Ross process with singular initial conditions.
Key words and phrases:
stochastic Volterra process; comparison principle; rough Cox-Ingersoll-Ross process; splitting method
O.C. is funded by the Irish Research Council grant GOIPG/2023/3129 and was previously supported by the ERASMUS Mobility project during which part of this research was carried out.
1. Introduction
Let $B$ be a standard one-dimensional Brownian motion, $0\neq K\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ a Volterra convolution kernel, $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ continuous with linear growth in the spatial variable, and $g\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. In this work, we study stochastic Volterra equations (SVEs) of the form
(1) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}B_{s}.$
In many applications, one assumes that the Volterra kernel is given by the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,2)$. Such equations have gained increased attention in mathematical Finance for the modelling of rough volatility, see [8, 14, 15, 16, 17, 23]. The weak existence of a solution was established in [4] whenever $b,\sigma$ are continuous with linear growth, $g$ constant and $K$ satisfies an additional regularity condition given below. An extension to processes with jumps and more general $g$ is discussed in [1]. When $b$ and $\sigma$ are Lipschitz continuous, the existence of a unique strong solution was shown in [2, Theorem A.1], see also [9, 31]. For non-Lipschitz coefficients, the pathwise uniqueness of solutions is more subtle and yet not fully understood. For the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$, the method proposed in [30] gives the existence and uniqueness of strong solutions whenever $\sigma$ is $\eta$-Hölder continuous and $\alpha\eta>\frac{1}{2}$. An extension of this method towards general drifts and non-convolution equations was recently discussed in [32]. Most recently, in [26], the author proves the weak existence and uniqueness in law for (1) provided that $b,\sigma$ are uniformly continuous, $\sigma$ is non-degenerate, and some technical condition on the modulus of continuity holds. In contrast to the existing literature, see also [32, 33], in this work we focus on continuous solutions where $g$ is not regular but belongs to $g\in L_{\mathrm{loc}}^{p}(\mathbb{R}_{+})\cap C((0,\infty))$ with an additional growth condition at the origin. Since such a case is yet not covered by the literature, we first provide in Theorem 2.6 a general result on the weak existence of continuous solutions for continuous coefficients with linear growth. In the main part of this work, we study comparison principles for solutions of (1). More precisely, given initial data $(g_{i},K,b_{i},\sigma)$ with $i=1,2$, we provide sufficient conditions such that the corresponding solutions $X^{1},X^{2}$ satisfy
$\mathbb{P}[X_{t}^{1}\leq X_{t}^{2},\ t\geq 0]=1.$
While such comparison principles are well-known for classical ordinary differential equations (see [38]) and for Markovian stochastic equations (see [27] for the continuous case and [21] for the case of continuous-state branching processes) much less is known for their Volterra counterparts. For instance, comparison principles for deterministic Volterra equations have been studied in [25, Chapter 12.1], while the first results on stochastic Volterra equations were obtained in [18, 36] for general drifts and restrictive assumptions on the diffusion which exclude our convolution setting (1). Based on the splitting method, Alfonsi was studied in [6] a numerical approximation of stochastic Volterra equations applicable to $g(t)=x$ constant, globally Lipschitz continuous coefficients $b,\sigma$, and $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. The latter work includes, up to our knowledge, the first general comparison principle for equations of the form (1). This splitting method has been also recently applied in [7] to prove the existence and uniqueness of nonnegative solutions of stochastic Volterra equations with jumps for non-Lipschitz continuous coefficients and regular Volterra kernels $K\in C^{2}(\mathbb{R}_{+})$. The reader interested in other numerical schemes for stochastic Volterra processes may consult [3, 5] and the references therein. In this work, we first extend Alfonsi’s numerical approximation [6] to non-constant $g$ with a particular focus on possibly singular functions, and time-dependent coefficients. Based on this approximation, we obtain a comparison principle that covers such cases including not necessarily Lipschitz continuous. Afterwards, we focus on kernels $K$ that are not completely monotone hence going beyond the framework discussed in [6]. In this generality, it is easy to see that for non-constant $g$ the comparison principle may fail (see the fractional Ornstein-Uhlenbeck process discussed in Example 2.5) unless further conditions on the driving function $g$ are imposed. The latter motivates the following definition.
Definition 1.1.
Let $g_{1},g_{2}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C((0,\infty))$ be such that
$g_{i}(t)=\widetilde{g}_{i}(t)+\int_{0}^{t}K(t-s)h_{i}(s)\,\mathrm{d}s,\qquad i=1,2,$
where $\widetilde{g_{1}},\widetilde{g}_{2}\in C((0,\infty))$ and $h_{1},h_{2}\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$. We call $(g_{1},g_{2})$ admissible, if they are of the above form, $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing on $(0,\infty)$, and $h_{2}-h_{1}\geq 0$.
This class of admissible functions is sufficiently rich to cover the most frequently used examples from applications including fractional kernels as illustrated below.
Example 1.2.
Given the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$. Define $g_{i}(t)=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}$ with $\gamma>\alpha$ and $x_{1}\leq x_{2}$. Then $g_{1},g_{2}$ are admissible with
  1. (i)
    If $\gamma\in(\alpha,1)$, then $\widetilde{g}_{i}(t)=0$ and $h_{i}(t)=x_{i}\frac{t^{(\gamma-\alpha)-1}}{\Gamma(\gamma-\alpha)}$.
  2. (ii)
    If $\gamma\geq 1$, then $\widetilde{g}_{i}(t)=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}$ and $h_{i}(t)=0$.
To illustrate our results, we consider the particular case of the fractional Ornstein-Uhlenbeck process given by
(2) $X_{t}=\frac{t^{\gamma-1}}{\Gamma(\gamma)}x+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\left(b+\beta X_{s}\right)\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\,\mathrm{d}B_{s}.$
where $x,b,\beta\in\mathbb{R}$, $\sigma,\gamma>0$, and $\alpha\in(1/2,1)$. The precise formulation is given in Example 2.5. Furthermore, based on our weak existence result obtained in Theorem 2.6 combined with the comparison principle, in Theorem 2.8 we obtain the weak existence of nonnegative weak solutions of (1). In particular, in Example 2.9, we construct a weak nonnegative solution for the rough Cox-Ingersoll-Ross process
$X_{t}=g(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\left(b+\beta X_{s}\right)\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\sqrt{X_{s}}\,\mathrm{d}B_{s},\qquad t>0.$
for a large class of functions $g$ that may be singular in $t=0$. The latter extends [2] where $g$ was assumed to be regular in zero, and provides a simpler condition on the nonnegativity of the solutions that does not require bounds on the increments of the resolvent of the first kind. The first key obstacle when dealing with stochastic Volterra equations stems from the absence of the semimartingale property. Indeed, while the classical methods described in [27, Chapter VI] are deeply based on the Itô formula, in our setting an analogous approach is only applicable when $K$ is sufficiently regular in the sense that $K\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$. The latter excludes, e.g., the Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,3/2)$. To address this issue, we use the splitting method proposed in [6] to approximate the solution via solutions of classical stochastic differential equations. Observing that the approximation satisfies the comparison principle, we may then pass to the limit. This approach covers $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. In particular, it covers the fractional kernel with $\alpha\in(1/2,1)$. On the other side, if $K$ is neither non-negativity preserving nor completely monotone, a comparison principle is, in general, not valid as demonstrated in Example 3.1. The latter is, e.g., the case for the fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>1$. Such a failure of the comparison principle is not surprising since solutions of (1) are typically not Markov processes and hence it is not clear if the past trajectory entering the equation via $b,\sigma$ violates the ordering of solutions. To exclude such a possibility we suppose that the drift is additionally monotone. Then, we first consider the case of regular kernels $K\in W_{\mathrm{loc}}^{1,2}(\mathbb{R}_{+})$ for which a semimartingale representation is available and a modification of the classical comparison principle for SDEs can be used to deduce the comparison principle. Afterwards, we extend these results to general $K\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C^{1}((0,\infty))$ with $K^{\prime}\geq 0$ or $K^{\prime}\leq 0$ by approximation. This work is organised as follows. In Section 2 we treat the case of completely monotone kernels, where we first prove a comparison principle for Lipschitz continuous $b,\sigma$. Then we prove the existence of continuous weak solutions for continuous $b,\sigma$ with at most linear growth, and finally derive the corresponding comparison principle in this case. As an application, we conclude Section 2 with a result on the weak existence of continuous nonnegative solutions of (1). Afterwards, in section 3 we study similar results for the case of not completely monotone kernels. Some technical results on the existence, uniqueness, and sample path regularity of solutions to stochastic Volterra equations are collected in Appendix A, while the convergence of the splitting method is studied in Appendix B.
2. Non-negativity preserving case
In this section, we study the case where the Volterra kernel preserves nonnegativity as recently introduced and studied in [6].
Definition 2.1.
A Volterra kernel $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ preserves non-negativity if it satisfies one of the following conditions:
  1. (a)
    $K$ is completely monotone with $K(0_{+})\coloneqq\lim_{t\downarrow 0}K(t)=+\infty$.
  2. (b)
    $K$ is nonincreasing, $K(0_{+})\in(0,\infty)$ and for each $N\in\mathbb{N}$, and all $x_{1},\dots,x_{N}\in\mathbb{R}$ and $0\leq t_{1}<\dots<t_{N}$ it holds that
    (3) $\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})\geq 0,\ \ \forall k\in\{1,\dots,N\}\Longrightarrow\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}x_{k}K(t-t_{k})\geq 0,\quad\forall t\geq 0.$
Here we use the convention $\sum_{\varnothing}\coloneqq 0$. Such a condition allows us to show that the Volterra kernel preserves the order structure subject to the discretisation of the process via the splitting method. Let us assume that the input data $(g,K,b,\sigma)$ of (1) satisfies throughout the entire paper the following regularity condition.
Assumption 2.2.
There exist $C>0$, $\xi\in[0,1]$ such that
$|b(t,x)|+|\sigma(t,x)|\leq(1+|x|)^{\xi}$
for a.a. $t\geq 0$ and $x\in\mathbb{R}$. Moreover, one of the following two cases holds:
  1. (i)
    $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$ or $\xi=0$ with $g\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. Moreover, $K\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and there exist $\gamma\in(0,1/2]$ and for each $T>0$ a constant $c(T)>0$ such that
    $\|K\|_{L^{2}([0,h])}+\|K(\cdot+h)-K(\cdot)\|_{L^{2}([0,T])}\leq c(T)h^{\gamma},\quad h\in[0,1].$
  2. (ii)
    $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$ for some $q\in(2,\infty)$, $\xi\in(0,1]$, there exists $\eta>0$ such that $K\in L^{2+\eta}_{\mathrm{loc}}(\mathbb{R}_{+})$ and there exist $\gamma\in(0,1/2]$ with and for each $T>0$ a constant $c(T,\eta)$ such that for $h\in[0,1]$
    $\|K\|_{L^{2+\eta}([0,h])}+\|K(\cdot+h)-K(\cdot)\|_{L^{2+\eta}([0,T])}\leq c(T,\eta)h^{\gamma}.$
    Moreover, the value $q$ satisfies
    (4) $q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}.$
The above assumption allows us to assume without loss of generality that for any solution $X$ of (1), $X-g$ is continuous, c.f. Proposition A.2. Note that such a condition is satisfied by the fractional Riemann-Liouville kernel as demonstrated in the next example.
Example 2.3.
Let $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1]$. Then $2+\eta<\frac{1}{1-\alpha}$, $\gamma=\alpha-1+\frac{1}{2+\eta}$. Hence (4) takes the particular form $q>2\xi\frac{1+1/\eta}{\alpha-1/2}$. By letting $\eta\nearrow\frac{2\alpha-1}{1-\alpha}$, we see that this condition is satisfied whenever
$q>\frac{\xi\alpha}{\left(\alpha-\frac{1}{2}\right)^{2}}.$
The following is our first result for non-negativity preserving kernels.
Theorem 2.4.
Suppose that $(g_{i},K,b_{i},\sigma)$ with $i=1,2$ satisfy assumption 2.2, that $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous and preserves non-negativity, and let $g_{1},g_{2}$ be admissible such that there exists $\delta>0$ with
(5) $\int_{0}^{T}\overline{g}_{i}(t)^{2+\delta}(1+h_{i}(t)^{2})\,\mathrm{d}t$ $<\infty,\qquad i=1,2$
where $\overline{g}_{i}(t)=\sup_{s\in[t,T]}|g_{i}(s)|$. Let $b_{1},b_{2},\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ be measurable and there exists $C>0$ such that for a.a. $t\geq 0$ and $x,y\in\mathbb{R}$,
(6) $|b_{1}(t,x)-b_{1}(t,y)|+|b_{2}(t,x)-b_{2}(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq C|x-y|.$
Finally, assume that the ordering condition $b_{1}(t,x)\leq b_{2}(t,x)$ holds. Let $X^{1},X^{2}$ be the unique strong solutions of
(7) $X_{t}^{i}=g_{i}(t)+\int_{0}^{t}K(t-s)b_{i}(s,X^{i}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X^{i}_{s})\,\mathrm{d}B_{s}$
for $i=1,2$. Then $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\,t>0]=1$.
Proof.
First note that equation (7) has a unique strong solutions $X^{1},X^{2}$ with sample paths in $L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ due to Proposition A.1. Moreover, given Assumption 2.2, it follows from Proposition A.2 that $X^{1}-g_{1}$, $X^{2}-g_{2}$ have a modification with continuous sample paths. Hence $X^{1},X^{2}$ are continuous on $(0,\infty)$. Finally, define
$\widetilde{b}_{i}(t,x)=b_{i}(t,x)+h_{i}(t),\quad t\in[0,T].$
Hence $\widetilde{b}_{1}\leq\widetilde{b}_{2}$ and (6) still holds with the same $C$, while the linear growth constant becomes $|\widetilde{b}_{1}(t,x)|+|\widetilde{b}_{2}(t,x)|\leq\widetilde{C}(t)(1+|x|)$ with $\widetilde{C}(t)=C+|h_{1}(t)|+|h_{2}(t)|$, so that $\widetilde{C}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. The proof is divided into three steps. Step 1: We introduce a similar splitting method to Alfonsi [6], but now applied for non-constant possibly singular at zero $g$ and time-dependent coefficients. Suppose first that $K$ is continuous with $K(0_{+})>0$. Let $T>0$, $N\in\mathbb{N}$ be fixed and $t_{k}=kT/N,$ $k\in\{0,1,\dots,N\},$ be an equidistant grid of $[0,T]$ with step size $T/N.$ Recursively, we define the process $(\widehat{X}^{i}_{t})_{t\in(0,T]}$ by setting
$\widehat{X}^{i}_{t}$ $=\widetilde{g}_{i}(t)+\sum_{k=1}^{N}\mathbbm{1}_{[t_{k},T]}(t)\left(K(t-t_{k})\int_{t_{k-1}}^{t_{k}}[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}]\right)$
where $(\xi^{i,k}_{t})_{t\in[t_{k-1},t_{k})}$, $k\in\{1,\dots,N\},$ denotes the strong solution of the auxiliary SDE
$\xi^{i,k}_{t}=\widehat{X}_{t_{k}-}^{i}+\int_{t_{k-1}}^{t}K(0_{+})[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}].$
Under the given conditions, it follows from Proposition B.2 that
(8) $\mathbb{E}[|\widehat{X}^{i}_{t}-X^{i}_{t}|^{2}]\longrightarrow 0,\qquad N\longrightarrow\infty.$
Thus it suffices to prove that $\widehat{X}_{t}^{1}\leq\widehat{X}_{t}^{2}$ a.s. for $t\in(0,T]$. Step 2: We consider the case where $K$ is continuous and non-negativity preserving with $K(0_{+})\in(0,\infty).$ Let $T>0$ be fixed. For $N\in\mathbb{N}$, we define the corresponding approximations $\widehat{X}^{1},\widehat{X}^{2}$ and denote by $\xi^{1},\xi^{2}$ the auxiliary processes that arise in their construction. Note that we have by continuity of $K$ and $g$ we have
$\xi^{i,k}_{t_{k}-}$ $=\widetilde{g}(t_{k})+\sum_{\ell=1}^{k-1}K(t_{k}-t_{\ell})\int_{t_{\ell-1}}^{t_{\ell}}[\widetilde{b}_{i}(s,\xi^{i,\ell}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,\ell}_{s})\,\mathrm{d}B_{s}]$
$\quad+K(t_{k}-t_{k})\int_{t_{k-1}}^{t_{k}}[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}]=\widehat{X}^{i}_{t_{k}}$
and similarly
$\widehat{X}^{i}_{t_{k}}-\widehat{X}^{i}_{t_{k}-}$
$\quad=\widetilde{g}_{i}(t_{k})+\sum_{\ell=1}^{k}K(t_{k}-t_{\ell})\int_{t_{\ell-1}}^{t_{\ell}}[\widetilde{b}_{i}(s,\xi^{i,\ell}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,\ell}_{s})\,\mathrm{d}B_{s}]$
$\qquad-\widetilde{g}_{i}(t_{k})-\sum_{\ell=1}^{k-1}K(t_{k}-t_{\ell})\int_{t_{\ell-1}}^{t_{\ell}}[\widetilde{b}_{i}(s,\xi^{i,\ell}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,\ell}_{s})\,\mathrm{d}B_{s}]$
$=K(0_{+})\int_{t_{k-1}}^{t_{k}}[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}].$
Hence we can represent them in compact form as
$\widehat{X}^{i}_{t}=\widetilde{g}_{i}(t)+\sum_{\ell=1}^{N}\mathbbm{1}_{[t_{\ell},T]}(t)\frac{\widehat{X}^{i}_{t_{\ell}}-\widehat{X}^{i}_{t_{\ell}-}}{K(0_{+})}K(t-t_{\ell}),\qquad t\in(0,T].$
We show by induction on $k$ that $\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t}$ on $(0,t_{k}]$. For $k=1,$ we have per assumption $\widehat{X}^{1}_{t}=\widetilde{g}_{1}(t)\leq\widetilde{g}_{2}(t)=\widehat{X}^{2}_{t}$ for $t\in(0,t_{1})$. Using the Comparison Principle for SDEs, see [27, Chapter VI, Theorem 1.1], we obtain $\widehat{X}^{1}_{t_{1}}=\xi_{t_{1}-}^{1,1}\leq\xi_{t_{1}-}^{2,1}=\widehat{X}_{t_{1}}^{2}$. Suppose that the result is true for some $k\geq 1$. Then, we have for all $j\in\{1,\dots,k\}$, by construction
$0\leq\widehat{X}_{t_{j}}^{2}-\widehat{X}_{t_{j}}^{1}=(\widetilde{g}_{2}-\widetilde{g}_{1})(t_{j})+\sum_{\ell=1}^{j}\frac{(\widehat{X}_{t_{\ell}}^{2}-\widehat{X}_{t_{\ell}-}^{2})-(\widehat{X}_{t_{\ell}}^{1}-\widehat{X}_{t_{\ell}-}^{1})}{K(0_{+})}K(t_{j}-t_{\ell}).$
Since $K$ preserves non-negativity and $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing, we obtain by Lemma B.3 that $\widehat{X}_{t}^{2}-\widehat{X}_{t}^{1}\geq 0$ holds for $t\in[t_{k},t_{k+1})$. Using again the comparison result for SDEs yields
$\widehat{X}_{t_{k+1}}^{1}=\xi^{1,k+1}_{t_{k+1}-}\leq\xi^{2,k+1}_{t_{k+1}-}=\widehat{X}^{2}_{t_{k+1}}.$
By induction, we get for $k=N$ that $\mathbb{P}[\widehat{X}^{1}_{T}\leq\widehat{X}^{2}_{T}]=1$. By (8) we conclude $\mathbb{P}[X^{1}_{T}\leq X^{2}_{T}]=1$. Since $T>0$ was arbitrary and $X$ has continuous sample paths on $(0,\infty)$, we obtain the desired comparison principle. Step 3: Let us now prove the assertion when $K(0_{+})=+\infty$ and $K$ is completely monotone. Using the Bernstein theorem, see [35, Theorem 4.8], we find a Borel measure $\mu$ on $\mathbb{R}_{+}$ such that
$K(t)=\int_{\mathbb{R}_{+}}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho),\quad t\in(0,\infty).$
Define, for $H>0$, the approximation $K^{H}\colon\mathbb{R}_{+}\longrightarrow\mathbb{R}_{+}$ by
$K^{H}(t)\coloneqq\int_{[0,H]}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho).$
Then $0\leq K^{H}(t)\leq K^{H}(0)$ and $|(K^{H})^{\prime}(t)|\leq HK^{H}(0)$. Hence, it satisfies Assumption 2.2 with $\gamma=1/2$ and therefore
$X^{i,H}_{t}=g_{i}(t)+\int_{0}^{t}K^{H}(t-s)b_{i}(s,X^{i,H}_{s})\,\mathrm{d}s+\int_{0}^{t}K^{H}(t-s)\sigma(s,X^{i,H}_{s})\,\mathrm{d}B_{s}$
admits a unique strong continuous solution for $i=1,2$. Since completely monotone kernels are non-negativity preserving and satisfy (3), see [6, Theorem 2.3], we may apply steps 1 and 2 to obtain $\mathbb{P}[X^{1,H}_{T}\leq X^{2,H}_{T}]=1$ for each $T>0$. Next we prove that
$\lim_{H\longrightarrow\infty}\mathbb{E}[|X^{i}_{T}-X^{i,H}_{T}|^{2}]=0$
which completes the proof. Indeed, for fixed $t>0$, we have by Jensen’s inequality and Itô’s isometry
$\mathbb{E}[|X^{i}_{t}-X_{t}^{i,H}|^{2}]$ $\lesssim\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[b(s,X^{i}_{s})^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t}K^{H}(t-s)^{2}\mathbb{E}[|b_{i}(s,X_{s})-b_{i}(s,X^{i,H}_{s})|^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[\sigma(s,X^{i}_{s})^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t}K^{H}(t-s)^{2}\mathbb{E}[|\sigma(s,X_{s})-\sigma(s,X^{i,H}_{s})|^{2}]\,\mathrm{d}s$
$\lesssim\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[(1+|X_{s}|)^{2\xi}]\,\mathrm{d}s$
$\qquad+\int_{0}^{t}K^{H}(t-s)^{2}\mathbb{E}[|X_{s}-X^{i,H}_{s}|^{2}]\,\mathrm{d}s$
where we have used the linear growth and Lipschitz condition. For the first term, we find by using Hölder’s inequality
$\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[(1+|X_{s}|)^{2\xi}]\,\mathrm{d}s$
$\quad\leq\left(\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s\right)^{\frac{2}{2+\eta}}\mathbb{E}\left[\left(\int_{0}^{T}(1+|X_{s}|)^{2\xi\left(1+\frac{2}{\eta}\right)}\,\mathrm{d}s\right)^{\frac{\eta}{2+\eta}}\right]$
$\quad\leq\left(\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s\right)^{\frac{2}{2+\eta}}\int_{0}^{T}\mathbb{E}\left[(1+|X_{s}|)^{2\xi\left(1+\frac{2}{\eta}\right)}\right]\,\mathrm{d}s$
where the second term is finite due to Assumption 2.2
$q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}\geq 2\xi\frac{1+\frac{1}{\eta}}{\frac{1}{2}+\frac{1}{2}\frac{\eta}{2+\eta}}=2\xi\left(1+\frac{2}{\eta}\right)$
and Proposition A.1. Applying the Volterra-type Gronwall inequality yields
$\mathbb{E}[|X_{t}-X^{i,H}_{t}|^{2}]\\ \leq\left(\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s\right)^{\frac{2}{2+\eta}}\left(\int_{0}^{T}\mathbb{E}\left[(1+|X_{s}|)^{2\xi\left(1+\frac{2}{\eta}\right)}\right]\mathrm{d}s\right)\left(1+\int_{0}^{t}R(s)\,\mathrm{d}s\right),$
where $R$ denotes the resolvent of the second kind of $c(K^{H})^{2}$, where $c$ is some large enough constant. Finally, the first term can be bounded by
$\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s$ $\leq\int_{0}^{T}\left(\int_{[H,\infty)}\mathrm{e}^{-\rho s}\,\mu(\mathrm{d}\rho)\right)^{\frac{\eta}{2+\eta}}\,\mathrm{d}s$
$\leq\int_{0}^{T}\mathrm{e}^{-sH(2+\eta)/2}K(s/2)^{2+\eta}\,\mathrm{d}s.$
By dominated convergence, the latter tends to zero as $H\longrightarrow\infty$, which proves the assertion. ∎
Note that the case where $g_{i}(t)=x_{i}$ for $i=1,2$ was already covered in [6]. Since the fractional Riemann-Liouville kernel satisfies Assumption 2.2 and is completely monotone, Theorem 2.4 can be applied to this case. Our next example shows that our comparison principle may also hold when $g_{1},g_{2}$ are not admissible.
Example 2.5.
Let $X^{1},X^{2}$ be fractional Ornstein-Uhlenbeck processes given as the unique strong solutions of (2), i.e.,
$X^{i}_{t}=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}(b_{i}+\beta X^{i}_{s})\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\,\mathrm{d}B_{s},$
where $b_{1}\leq b_{2}$, $\beta,\sigma\in\mathbb{R}$, $\alpha\in(1/2,1)$, $\gamma>0$, and $x_{1}\leq x_{2}$. If $\beta\geq 0$, then the comparison principle holds. If $\beta<0$, then the comparison principle holds if and only if $\gamma\geq\alpha$.
Proof.
It follows from [4] that $X^{i}$ is explicitly given by
$X_{t}^{i}=x_{i}t^{\gamma-1}E_{\alpha,\gamma}(\beta t^{\alpha})+\int_{0}^{t}s^{\alpha-1}E_{\alpha,\alpha}(\beta s^{\alpha})b_{i}\,\mathrm{d}s+\int_{0}^{t}(t-s)^{\alpha-1}E_{\alpha,\alpha}(\beta(t-s)^{\alpha})\,\mathrm{d}B_{s}$
where $E_{\alpha,\gamma}(z)=\sum_{n=0}^{\infty}\frac{z^{n}}{\Gamma(\alpha n+\gamma)}$ denotes the two-parameter Mittag-Leffler function. Hence
$X_{t}^{2}-X_{t}^{1}=(x_{2}-x_{1})t^{\gamma-1}E_{\alpha,\gamma}(\beta t^{\alpha})+\int_{0}^{t}s^{\alpha-1}E_{\alpha,\alpha}(\beta s^{\alpha})(b_{2}-b_{1})\,\mathrm{d}s.$
Since $E_{\alpha,\alpha}(\beta t^{\alpha})\geq 0$ for any $\beta,t$, the comparison principle is satisfied if and only if $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ for all $t\geq 0$. If $\beta\geq 0$, then $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ and nothing needs to be shown. Suppose that $\beta<0$. If $\gamma\geq\alpha$, then $t\longmapsto E_{\alpha,\gamma}(-t^{\alpha})$ is completely monotone, see [24, Section 4.9.2], and hence the comparison principle holds. In this case $g_{1},g_{2}$ are admissible, see Example 1.2. Conversely, suppose that $\gamma<\alpha$. Then $(0,\infty)\ni t\longmapsto t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ is integrable and its Laplace transform is given by $\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\mathrm{e}^{-st}\,\mathrm{d}s=\frac{s^{\alpha-\gamma}}{s^{\alpha}+|\beta|}$ for $s\geq 0$. Evaluating this in $s=0$, shows that $\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\,\mathrm{d}s=0$ and hence $(0,\infty)\ni t\longmapsto E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ needs to attain negative values. Consequently, the comparison principle fails to hold. ∎
Next, let us examine cases where the coefficients $b_{1},b_{2},\sigma$ are not Lipschitz continuous. Since, in this generality, the weak existence of solutions is not guaranteed from the literature, below we complement the latter by proving the weak existence of continuous solutions when the function $g$ may be singular in $t=0$. Our proof follows the typical tightness arguments, however, in contrast to existing arguments such as [33, 4], here we introduce an auxiliary weighted space of continuous functions to control the order of singularity of the function $g$.
Theorem 2.6.
Let $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ be continuous, nonincreasing, and not identically zero. If $(g,K,b,\sigma)$ satisfies Assumption 2.2 and $g$ is continuous on $(0,\infty)$ such that there exists $\delta\in(0,1/2)$ such that for every $T>0$
(9) $\sup_{t\in(0,T]}t^{\delta}|g(t)|<\infty,$
then the stochastic Volterra equation
(10) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}B_{s},\qquad t>0$
admits a continuous weak solution.
Proof.
Following [33, Lemma 3.6], we define for $f\in\{b,\sigma\}$ a sequence of functions
(11) $f_{n}(t,x)=\psi_{n}(x)\int_{\mathbb{R}}f(t,x-y)\varphi_{n}(y)\,\mathrm{d}y,\quad,n\in\mathbb{N},$
where $\psi_{n}\colon\mathbb{R}\longrightarrow[0,1]$ with $\operatorname{supp}\psi_{n}\subset[-(n+1),n+1]$ and $\psi_{n}(x)=1$ for $x\in[-n,n]$, where $\varphi(y)=\frac{1}{c_{n}}(1-y^{2})^{n}\mathbbm{1}_{[-1,1]}(y)$ with $c_{n}=\int_{[-1,1]}(1-y^{2})^{n}\,\mathrm{d}y$. Then $(f_{n})_{n\in\mathbb{N}}$ is a sequence of Lipschitz continuous functions in the second variable and $\lim_{n\longrightarrow\infty}f_{n}(t,\cdot)=f(t,\cdot)$ uniformly on compacts for all $t\geq 0$. Moreover, the following growth condition is satisfied uniformly in $n$,
(12) $|b_{n}(t,x)|+|\sigma_{n}(t,x)|\lesssim C(1+|x|)^{\xi},\quad\text{a.a. }t\geq 0,x\in\mathbb{R},$
Let $(X^{n})_{n\in\mathbb{N}}$ be the family of processes constructed for $(g,K,b_{n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. Thanks to Proposition A.2 and (12), one readily checks that for $T>0$ and all $s,t\in[0,T]$,
$\sup_{n\in\mathbb{N}}\mathbb{E}[{|(X_{t}^{n}-g(t))-(X_{s}^{n}-g(s))|^{p}}]\leq c|t-s|^{1+\theta}$
holds for some $\theta>0$, $p\geq 2$ and a constant $c$ that only depends on $C$, $T$, $K$, $p$. By applying Kolmogorov’s tightness criterion (see e.g. [34, Chapter XIII, Theorem 1.8]), we conclude that $(X^{n}-g)_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+})$ with limit denoted by ¯$\overline{\mathbb{P}}_{g}$. By abuse of notation, we denote this subsequence also by $(X^{n}-g)_{n\in\mathbb{N}}$. Let $C(\mathbb{R}_{+})_{0}\subset C(\mathbb{R}_{+})$ be the space of continuous functions that vanish at $t=0$. Since $\mathcal{L}(X^{n}-g)(C(\mathbb{R}_{+})_{0})=1$ it follows by the Portmanteau theorem that
$1=\limsup_{n\to\infty}\mathcal{L}(X^{n}-g)(C(\mathbb{R})_{0})\leq\overline{\mathbb{P}}_{g}(C(\mathbb{R}_{+})_{0}).$
Next, to prove weak convergence of $(X^{n})_{n\in\mathbb{N}}$, let $C_{\kappa}([0,T])$ with $\kappa>0$ be the Banach space of continuous functions $f\colon(0,T]\longrightarrow\mathbb{R}$ with finite norm $\|f\|_{\kappa,T}\coloneqq\sup_{t\in(0,T]}t^{\kappa}|f(t)|$ such that $\lim_{t\downarrow 0}t^{\kappa}f(t)$ exists. Note that $\iota\colon C_{\kappa}([0,T])\longrightarrow C([0,T])$ defined by $\iota(f)(t)=t^{\kappa}f(t)$ satisfies $\|\iota(f)\|_{[0,T]}=\|f\|_{\kappa,T}$, and hence is an isometric isomorphism, whence $C_{\kappa}([0,T])$ is a polish space. Finally, we let $C_{\kappa}(\mathbb{R}_{+})\coloneqq\cap_{k\in\mathbb{N}}C_{\kappa}([0,k])$ be equipped with the metric
$\varrho(f,g)\coloneqq\sum_{k=1}^{\infty}2^{-k}(\|f-g\|_{\kappa,k}\land 1).$
By assumption (9), $g\in C_{\kappa}(\mathbb{R}_{+})$ for $\kappa>\delta$. Hence $\Phi\colon C(\mathbb{R}_{+})_{0}\longrightarrow C_{\kappa}(\mathbb{R}_{+})$ given by $\Phi(f)=f+g$ defines a continuous mapping. Thus, $(\Phi(X^{n}-g))_{n\in\mathbb{N}}=(X^{n})_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+})$ with weak limit denoted by $\overline{\mathbb{P}}$. Note that ¯$\overline{\mathbb{P}}_{g},\overline{\mathbb{P}}$ are related by
(13) ¯$\overline{\mathbb{P}}_{g}\circ\Phi^{-1}=\overline{\mathbb{P}}.$
It remains to show that $\overline{\mathbb{P}}$ determines a weak solution of (10). Let $L$ be the resolvent of the first kind of $K$111i.e. a locally finite (positive) measure $L$ such that $L\ast K=K\ast L=1$. The latter exists since $K\neq 0$ is non-negative and non-increasing, see [25, Theorem 5.5.5].. Let $(M_{t}^{n})_{t\geq 0}$ be the continuous $\mathbb{P}$-martingale for the filtration $\mathbb{F}$ given by
$M_{t}^{n}=\int_{0}^{t}\sigma_{n}(s,X^{n}_{s})\,\mathrm{d}B_{s}=\int_{[0,t]}(X^{n}_{t-s}-g(t-s))\,L(\mathrm{d}s)-\int_{0}^{t}b_{n}(s,X^{n}_{s})\,\mathrm{d}s$
where the second equality follows from [4, Lemma 2.6]. Moreover, since (12) holds uniformly in $n\in\mathbb{N}$, it follows that $(M_{t}^{n})_{n\in\mathbb{N}}$ is uniformly integrable due to
$\sup_{n\in\mathbb{N}}\mathbb{E}\left[\sup_{t\in[0,T]}|M_{t}^{n}|^{2}\right]$ $\leq\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}\left[\sigma_{n}(s,X_{s}^{n})^{2}\right]\mathrm{d}s$
$\lesssim C\left(T+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}\left[|X_{s}^{n}|^{2}\right]\mathrm{d}s\right)<\infty.$
Define a continuous mapping $\mathcal{M}\colon C(\mathbb{R}_{+})_{0}\longrightarrow C(\mathbb{R}_{+})_{0}$ by
$\mathcal{M}_{t}(w)=\int_{[0,t]}w(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b(s,\Phi(w)(s))\,\mathrm{d}s,$
where the continuity is guaranteed by [25, Corollary 3.6.2 (iii)]. Since $X^{n}-g\in C(\mathbb{R}_{+})_{0}$, by continuous mapping theorem combined with (13), we obtain
$M^{n}=\mathcal{M}(X^{n}-g)\Longrightarrow\overline{\mathbb{P}}_{g}\circ\mathcal{M}^{-1}=\overline{\mathbb{P}}\circ(\mathcal{M}^{\Phi})^{-1}$
where $\mathcal{M}^{\Phi}\colon\Phi(C(\mathbb{R}_{+})_{0})\longrightarrow C(\mathbb{R}_{+})$ is given by
$\mathcal{M}_{t}^{\Phi}(w)\coloneqq\mathcal{M}_{t}(\Phi^{-1}(w))=\int_{[0,t]}(w(t-s)-g(t-s))\,L(\mathrm{d}s)-\int_{0}^{t}b(s,w(s))\,\mathrm{d}s.$
Next, we show that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale for the natural filtration generated by coordinate process on $C_{\kappa}(\mathbb{R}_{+})$, and determine its quadratic variation. The continuity follows from $\Phi(C(\mathbb{R}_{+})_{0})=\{f\in C_{\kappa}(\mathbb{R}_{+}):f-g\in C(\mathbb{R}_{+})_{0}\}$, [25, Corollary 3.6.2 (iii)] and ¯$\overline{\mathbb{P}}(\Phi(C(\mathbb{R}_{+})_{0}))=\overline{\mathbb{P}}_{g}(C(\mathbb{R}_{+})_{0})=1.$ Let $s<t$, $m\in\mathbb{N}$, $f\colon\mathbb{R}^{m}\longrightarrow\mathbb{R}$ bounded and continuous, and $0\leq s_{1}\leq\dots\leq s_{m}\leq s$. Let $(x(t))_{t\geq 0}$ denote the coordinate process $x(t):C_{\kappa}(\mathbb{R}_{+};\mathbb{R})\longrightarrow\mathbb{R},\ \ w\longmapsto w(t)$. Applying [11, Theorem 3.5] yields
(14) $\mathbb{E}^{\overline{\mathbb{P}}}\left[f(x(s_{1}),\dots,x(s_{m}))\left(\mathcal{M}_{t}^{\Phi}-\mathcal{M}_{s}^{\Phi}\right)\right]=\lim_{n\longrightarrow\infty}\mathbb{E}[f(X^{n}_{s_{1}},\dots,X^{n}_{s_{m}})(M_{t}^{n}-M_{s}^{n})]=0.$
A monotone class argument implies that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a $\overline{\mathbb{P}}$-martingale. To determine its quadratic variation, let us write
$\langle M^{n}\rangle=\int_{0}^{\cdot}\left(\sigma_{n}(s,X_{s}^{n})^{2}-\sigma(s,X_{s}^{n})^{2}\right)\,\mathrm{d}s+\int_{0}^{\cdot}\sigma(s,X_{s}^{n})^{2}\,\mathrm{d}s.$
Since $X^{n}\Longrightarrow\overline{\mathbb{P}}$, the continuous mapping theorem implies that
$\int_{0}^{\cdot}\sigma(s,X_{s}^{n})^{2}\,\mathrm{d}s\Longrightarrow\mathcal{L}_{\overline{\mathbb{P}}}\left(\int_{0}^{\cdot}\sigma(s,x(s))\,\mathrm{d}s\right).$
Thus, by Slutsky’s theorem, it suffices to show that the first term converges to zero in probability. Let $T>0$, then
$\ \int_{0}^{T}\mathbb{E}\left[\left|\sigma_{n}(s,X_{s}^{n})^{2}-\sigma(s,X_{s}^{n})^{2}\right|\right]\mathrm{d}s$
$\lesssim\left(\int_{0}^{T}\mathbb{E}\left[\sigma_{n}(s,X_{s}^{n})^{2}+\sigma(s,X_{s}^{n})^{2}\right]\mathrm{d}s\right)^{1/2}\left(\int_{0}^{T}\mathbb{E}\left[\left(\sigma_{n}(s,X_{s}^{n})-\sigma(s,X_{s}^{n})\right)^{2}\right]\mathrm{d}s\right)^{1/2}$
$\lesssim\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}[|X_{s}^{n}|^{2}]\,\mathrm{d}s\right)^{1/2}\left(\int_{0}^{T}\mathbb{E}\left[(\sigma_{n}(s,X_{s}^{n})-\sigma(s,X^{n}_{s}))^{2}\right]\mathrm{d}s\right)^{1/2}$
$\lesssim R^{-\lambda/2}\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}[|X_{s}^{n}|^{2}]\mathrm{d}s\right)^{1/2}\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}\left[|X_{s}^{n}|^{2+\lambda}\right]\mathrm{d}s\right)^{1/2}$
$\qquad+\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}[|X_{s}^{n}|^{2}]\,\mathrm{d}s\right)^{1/2}\left(\int_{0}^{T}\|\sigma_{n}(s,\cdot)-\sigma(s,\cdot)\|_{[-R,R]}^{2}\,\mathrm{d}s\right)^{1/2}$
where we have used (12) and the inequality
$\mathbb{E}[|\sigma_{n}(s,X^{n}_{s})-\sigma(s,X_{s}^{n})|^{2}]$ $\lesssim\mathbb{E}\left[\mathbbm{1}_{\{|X^{n}_{s}|>R\}}\left(|\sigma_{n}(s,X^{n}_{s})|^{2}+|\sigma(s,X_{s}^{n})|^{2}\right)\right]$
$\qquad\qquad+\|\sigma_{n}(s,\cdot)-\sigma(s,\cdot)\|_{[-R,R]}^{2}$
$\lesssim\frac{C}{R^{\lambda}}\left(1+\mathbb{E}[|X^{n}_{s}|^{2+\lambda}]\right)+\|\sigma_{n}(s,\cdot)-\sigma(s,\cdot)\|_{[-R,R]}^{2},$
where $\lambda=q-2>0$ if $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}),q\in(2,\infty),$ and $\lambda=1$ if $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$. Thus, letting first $n\longrightarrow\infty$ and then $R\longrightarrow\infty$, proves the asserted. Arguing as in (14) yields that $(\mathcal{M}^{\Phi})^{2}-\int_{0}^{\cdot}\sigma(s,x(s))^{2}\,\mathrm{d}s$ is an $\overline{\mathbb{P}}$-martingale. Consequently, we have shown that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale with quadratic variation
$\langle\mathcal{M}^{\Phi}\rangle_{t}=\int_{0}^{t}\sigma(s,x(s))^{2}\,\mathrm{d}s.$
By martingale representation theorem [34, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a Brownian motion $W$ and a predictable process $\alpha$ such that $\alpha_{s}^{2}=\sigma(s,\widehat{X}_{s})^{2}$ and $\mathcal{M}^{\Phi}_{t}=\int_{0}^{t}\alpha_{s}\,\mathrm{d}W_{s}$ where $\widehat{X}$ denotes the processes $x$ considered on the enlargement such that $\mathcal{L}(\widehat{X})=\mathcal{L}(x)$. Consequently, by [4, Lemma 2.6]
$\widehat{X}_{t}$ $=g(t)+\int_{0}^{t}K(t-s)b(s,\widehat{X}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,\widehat{X}_{s})\,\mathrm{d}W_{s}$
which completes the proof. ∎
Remark that $g$ given as in (9) automatically satisfies $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$ and hence condition (4) is satisfied whenever
$2\xi\delta<\frac{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}{1+\frac{1}{\eta}}.$
Next, we turn to the comparison principle beyond Lipschitz continuous coefficients. Since a comparison principle as formulated in Theorem 2.4 would imply pathwise uniqueness of (1), it is in this generality out of reach. However, by approximation, we show that for any two weak solutions $X^{1},X^{2}$ there exists a coupling $(\widehat{X}^{1},\widehat{X}^{2})$ on another filtered probability space on which the comparison principle holds.
Theorem 2.7.
Let $(g_{i},K,b_{i},\sigma)$, $i=1,2,$ satisfy Assumption 2.2 with $b_{1}(t,x)\leq b_{2}(t,x)$, $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous, preserves non-negativity and suppose there exists $\delta\in(0,1/2)$ with (9) holds for $g_{1},g_{2}$. Moreover, assume that $g_{1},g_{2}$ satisfy (5) and are admissible. Then for any two continuous weak solutions $(X^{i},(\Omega^{i},\mathcal{F}^{i},\mathbb{F}^{i},\mathbb{P}^{i}),B^{i})$, $i=1,2,$ of (10), there exists a continuous weak solution $((\widehat{X}^{1},\widehat{X}^{2}),(\widehat{\Omega},\widehat{\mathcal{F}},\widehat{\mathbb{F}},\widehat{\mathbb{P}}),\widehat{B})$ of the two dimensional stochastic Volterra equation
(23) $\begin{pmatrix}\widehat{X}^{1}_{t}\\ \widehat{X}^{2}_{t}\end{pmatrix}=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(s,\widehat{X}^{1}_{s})\\ b_{2}(s,\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(s,\widehat{X}^{1}_{s})\\ \sigma(s,\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}\widehat{B}_{s}$
such that $\mathcal{L}(\widehat{X}^{i})=\mathcal{L}(X^{i})$, $i=1,2,$ and ^$\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$.
Proof.
Let us define $b_{1,n},b_{2,n},\sigma_{n}$ as in (11). Since $b_{1}\leq b_{2}$, it is clear that this approximation inherits the ordering condition $b_{1,n}\leq b_{2,n}$. Let $(X^{n})_{n\in\mathbb{N}},(Y^{n})_{n\in\mathbb{N}}$ be the family of processes constructed for $(g_{i},K,b_{i,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. By Theorem 2.4, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},t>0]=1$. Following the arguments presented in the proof of Theorem 2.6, we firstly deduce that $((X^{n}-g_{1},Y^{n}-g_{2}))_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+};\mathbb{R}^{2})$. Secondly, we extend the spaces $C(\mathbb{R}_{+})_{0}\coloneqq C(\mathbb{R}_{+};\mathbb{R})_{0}$, $C_{\kappa}(\mathbb{R}_{+})\coloneqq C_{\kappa}(\mathbb{R}_{+};\mathbb{R})$ in an obvious way to $C(\mathbb{R}_{+};\mathbb{R}^{2})_{0}$, $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with $\kappa>\delta$ and infer that $((X^{n},Y^{n}))_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with weak limit denoted by $\overline{\mathbb{P}}.$ To prove the desired comparison result it is essential to observe that $\overline{\mathbb{P}}$ has inherited the ordering structure. Indeed, let $(x,y)=((x,y)(t))_{t\geq 0}$ be the coordinate process
$(x,y)(t)\colon C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})\longrightarrow\mathbb{R}^{2},\quad w=(w_{1},w_{2})\longmapsto(w_{1}(t),w_{2}(t)).$
Let $A\coloneqq\{w=(w_{1},w_{2})\in C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})\mid w_{2}(t)-w_{1}(t)\geq 0,t>0\}$. Writing
$A_{t}\coloneqq\{(w_{1},w_{2})\in C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})\mid w_{2}(t)-w_{1}(t)\geq 0\}=\pi_{t}^{-1}([0,\infty))$
where $\pi_{t}\colon(C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2}),\varrho)\longrightarrow\mathbb{R}$, $\pi_{t}(w_{1},w_{2})=w_{2}(t)-w_{1}(t)$ is a continuous function, we conclude that $A_{t}$ is closed in $((C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2}),\varrho)$. Hence also $A=\bigcap_{t>0}A_{t}$ is a closed subset of $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$. An application of the Portmanteau theorem yields
¯$\overline{\mathbb{P}}[x(t)\leq y(t),t\geq 0]=\overline{\mathbb{P}}[A]\geq\limsup_{n\longrightarrow\infty}\mathbb{P}_{n}[A]=\limsup_{n\longrightarrow\infty}\mathbb{P}[X_{t}^{n}\leq Y_{t}^{n},t>0]=1.$
Finally, we show that $\overline{\mathbb{P}}$ determines a weak solution to the 2-dimensional stochastic Volterra equation (23). Similarly to the proof of Theorem 2.6, we define the continuous $\mathbb{P}$-martingales
$M^{n}_{t}\coloneqq\int_{0}^{t}\begin{pmatrix}\sigma_{n}(s,X^{n}_{s})\\ \sigma_{n}(s,Y^{n}_{s})\end{pmatrix}\,\mathrm{d}B_{s}=\begin{pmatrix}\int_{[0,t]}X^{n}_{t-s}-g_{1}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{1,n}(s,X^{n}_{s})\,\mathrm{d}s\\ \int_{[0,t]}Y^{n}_{t-s}-g_{2}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{2,n}(s,X^{n}_{s})\,\mathrm{d}s\end{pmatrix},$
where $L$ denotes the resolvent of the first kind of $K$. Moreover, we observe that $M^{n}\Longrightarrow\overline{\mathbb{P}}\circ\mathcal{M}^{-1}$, where $\mathcal{M}$ is on $\{(w_{1},w_{2})\in C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2}):(w_{1}-g_{1},w_{2}-g_{2})\in C(\mathbb{R}_{+};\mathbb{R}^{2})_{0}\}$ given by
(24) $\mathcal{M}_{t}(w_{1},w_{2})\coloneqq\begin{pmatrix}\int_{[0,t)}w_{1}(t-s)-g_{1}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{1}(s,w_{1}(s))\,\mathrm{d}s\\ \int_{[0,t)}w_{2}(t-s)-g_{2}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{2}(s,w_{2}(s))\,\mathrm{d}s\\ \end{pmatrix}$
Furthermore, we can show that $(\mathcal{M}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale for the natural filtration generated by the coordinate process on $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ and its quadratic variation is given by
$\langle\mathcal{M}\rangle_{t}=\int_{0}^{t}\begin{pmatrix}\sigma(s,x(s))^{2}&\sigma(s,x(s))\sigma(s,y(s))\\ \sigma(s,x(s))\sigma(s,y(s))&\sigma(s,y(s))^{2}\end{pmatrix}\mathrm{d}s.$
By martingale representation theorem [34, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a two-dimensional Brownian motion $(W^{1},W^{2})$ and a $\mathbb{R}^{2\times 2}$-valued predictable process $\alpha$ such that
$\mathcal{M}(t)=\int_{0}^{t}\alpha_{s}\,\mathrm{d}\!\begin{pmatrix}W^{1}_{s}\\ W^{2}_{s}\end{pmatrix}\qquad\text{ with }\qquad\alpha_{s}\alpha_{s}^{\top}=\begin{pmatrix}\sigma(s,\widehat{X}_{s})^{2}&\sigma(s,\widehat{X}_{s})\sigma(s,\widehat{Y}_{s})\\ \sigma(s,\widehat{X}_{s})\sigma(s,\widehat{Y}_{s})&\sigma(s,\widehat{Y}_{s})^{2}\end{pmatrix}$
where $\widehat{X},\widehat{Y}$ denote the processes $x,y$ considered on the enlargement such that $\operatorname{Law}(\widehat{X})=\operatorname{Law}(x)$, $\operatorname{Law}(\widehat{Y})=\operatorname{Law}(y)$. Note that any other choice of $\widetilde{\alpha}$ that satisfies $\widetilde{\alpha}\widetilde{\alpha}^{\top}=\alpha\alpha^{\top}$ gives the same law of $\mathcal{M}$ under $\overline{\mathbb{P}}$. Let us define
$\widehat{\alpha}_{t}=\frac{1}{\sqrt{2}}\begin{pmatrix}\sigma(s,\widehat{X}_{t})&\sigma(s,\widehat{X}_{t})\\ \sigma(s,\widehat{Y}_{t})&\sigma(s,\widehat{Y}_{t})\end{pmatrix}\ \text{ and }\ \widehat{\mathcal{M}}(t)=\int_{0}^{t}\widehat{\alpha}_{s}\,\mathrm{d}\!\begin{pmatrix}W^{1}_{s}\\ W^{2}_{s}\end{pmatrix}.$
Then ^$\langle\widehat{\mathcal{M}}\rangle_{t}=\int_{0}^{t}\widehat{\alpha}_{s}\widehat{\alpha}_{s}^{\top}\,\mathrm{d}s=\int_{0}^{t}\alpha_{s}\alpha_{s}^{\top}\,\mathrm{d}s=\langle\mathcal{M}\rangle_{t}$. Hence $\widehat{\mathcal{M}}$ and $\mathcal{M}$ have the same law. Consequently, by (24) and [4, Lemma 2.6],
$\begin{pmatrix}\widehat{X}_{t}\\ \widehat{Y}_{t}\end{pmatrix}$ $=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(s,\widehat{X}_{s})\\ b_{2}(s,\widehat{Y}_{s})\end{pmatrix}\mathrm{d}s+\frac{1}{\sqrt{2}}\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(s,\widehat{X}_{s})\,\mathrm{d}(W^{1}_{s}+W^{2}_{s})\\ \sigma(s,\widehat{Y}_{s})\,\mathrm{d}(W^{1}_{s}+W^{2}_{s})\end{pmatrix}$
$=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(s,\widehat{X}_{s})\\ b_{2}(s,\widehat{Y}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(s,\widehat{X}_{s})\\ \sigma(s,\widehat{Y}_{s})\end{pmatrix}\mathrm{d}\widehat{B}_{s},$
where $\widehat{B}=\frac{1}{\sqrt{2}}(W^{1}+W^{2})$ is a one-dimensional Brownian motion. This completes the proof. ∎
As an application of the above theorem, we prove the existence of a non-negative continuous weak solution for a rich class of SVEs with singular initial conditions that include the rough Cox-Ingersoll-Ross process.
Theorem 2.8.
Let $(g,K,b,\sigma)$ satisfy Assumption 2.2 with $b(t,0)\geq 0$ and $\sigma(t,0)=0$ for $t\geq 0$, $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous, preserves non-negativity. Suppose that $g$ is such that
$g(t)=\widetilde{g}(t)+\int_{0}^{t}K(t-s)h(s)\,\mathrm{d}s$
with $0\leq h\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ and $0\leq\widetilde{g}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C((0,\infty))$ is non-decreasing, (5) holds, and there exists $\delta\in(0,1/2)$ with (9). Then
(25) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}B_{s}$
admits a non-negative continuous weak solution.
Proof.
The existence of a continuous weak solution $X$ follows from Theorem 2.6. Moreover, $Y_{t}=0$, $t>0$, clearly solves
$Y_{t}=\int_{0}^{t}K(t-s)(b(s,Y_{s})-b(s,0))\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,Y_{s})\,\mathrm{d}B_{s}.$
Since the pair $(0,g)$ is admissible, by Theorem 2.7 there exists a continuous weak solution $(\widehat{Y},\widehat{X})$ of the associated two-dimensional Volterra equation (23) such that $\mathcal{L}(\widehat{Y})=\mathcal{L}(Y)$, $\mathcal{L}(\widehat{X})=\mathcal{L}(X)$ and ^$\widehat{\mathbb{P}}[\widehat{Y}_{t}\leq\widehat{X}_{t},t>0]=1.$ We claim that ^$\widehat{\mathbb{P}}[\widehat{X}_{t}\geq 0,t>0]=1.$ Indeed, we have $11\geq\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cup\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]\geq\widehat{\mathbb{P}}[\widehat{Y}_{t}\geq 0,t>0]=$ 1. Consequently,
11 11 =^[$=\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cup\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]$
=^$=\widehat{\mathbb{P}}[\widehat{Y}_{t}>0,t>0]+\widehat{\mathbb{P}}[\widehat{Y}_{t}\leq\widehat{X}_{t},t>0]-\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cap\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]$
which yields
11 11 =^[$=\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cap\{\widehat{Y}_{t}\leq\widehat{X}_{t},t>0\}]$
=^[$=\widehat{\mathbb{P}}[\{\widehat{Y}_{t}\geq 0,t>0\}\cap\{0\leq\widehat{X}_{t},t>0\}]$
^$\leq\widehat{\mathbb{P}}[0\leq\widehat{X}_{t},t>0].$
Thus, since $\mathcal{L}(\widehat{X})=\mathcal{L}(X)$, it is the desired non-negative continuous weak solution of (25). ∎
Below we illustrate this result by an application to the rough Cox-Ingersoll-Ross process.
Example 2.9 (rough Cox-Ingersoll-Ross process).
Let $\lambda,\theta,\sigma>0$, and $\alpha\in(1/2,1)$. Then for each $g$ satisfying the condition of Theorem 2.8, there exists a continuous nonnegative weak solution of
$X_{t}=g(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\lambda\left(\theta-X_{s}\right)\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\sqrt{X_{s}}\,\mathrm{d}B_{s}.$
Moreover, this process satisfies the conditions of Theorem 2.7.
Note that the non-negativity or the comparison principle holds on a newly constructed filtered probability space and subsequently does not require pathwise uniqueness for the rough Cox-Ingersoll-Ross process that is still an open problem, see also [30, 32] for related results. Moreover, we do not prove the uniqueness in law for the rough Cox-Ingersoll-Ross process constructed above. The latter would certainly follow from an extension of the affine transformation formula given in [28] applied to this setting. Since such a possible research direction goes beyond the scope of comparison principles studied in this work, we leave the details for the interested reader. However, it is relatively easy to prove the uniqueness of the time-marginals. Indeed, let $X^{g}$ be the process given by Example 2.9. Then it is easy to see that
(26) $\mathbb{E}[X^{g}_{t}]=g(t)-\lambda\int_{0}^{t}E_{\lambda}(t-s)g(s)\,\mathrm{d}s+\lambda\theta\int_{0}^{t}E_{\lambda}(s)\,\mathrm{d}s$
where $E_{\lambda}$ is the unique solution of $E_{\lambda}+\lambda K_{\alpha}\ast E_{\lambda}=K_{\alpha}$. Let $Y^{g}$ be another rough Cox-Ingersoll-Ross process with the same $g$. Since $(g,g)$ is admissible, by Theorem 2.7 we find a coupling $(\widehat{X}^{g},\widehat{Y}^{g})$ that satisfies the comparison principle, say ^$\widehat{\mathbb{P}}[\widehat{X}_{t}^{g}\leq\widehat{Y}_{t}^{g}]=1$. Hence we obtain
$W_{1}(X^{g}_{t},Y_{t}^{g})\leq\widehat{\mathbb{E}}\left[|\widehat{Y}_{t}^{g}-\widehat{X}_{t}^{g}|\right]=0$
where $W_{1}$ denotes the Wasserstein-1-distance and the last relation follows from (26). Similar arguments based on the comparison principle for Markov processes, have been used in [19, 20, 22] to prove the convergence to equilibrium including a rate of convergence. Unfortunately, due to the absence of the Markov property, an immediate extension (at least without the use of Markovian lifts) does not carry over.
3. Regular Case
In this section, we study the case where $K$ does not preserve non-negativity. The latter includes, e.g., the fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>1$. For such kernels $K$ and without further assumptions on the drift coefficients, the comparison principle does not hold as demonstrated in the next example.
Example 3.1.
Let $X^{i}$ with $i=1,2$ be the unique solution of
$X_{t}^{i}=x_{i}\frac{t^{\beta-1}}{\Gamma(\beta)}-\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}X_{s}^{i}\,\mathrm{d}s+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\,\mathrm{d}B_{s}$
where $\alpha>1$, $\beta>0$ and $x_{1},x_{2}\in\mathbb{R}$. Recall that the comparison principle holds if and only if $E_{\alpha,\beta}(-t^{\alpha})\geq 0$ for all $t\geq 0$. If $0<\beta<\alpha$ and $\alpha>1$, then $E_{\alpha,\beta}(-t^{\alpha})$ necessarily takes negative values and hence the comparison principle fails.
This example illustrates that, to obtain a comparison principle, further conditions need to be imposed. In the above example, the drift $b(x)=-x$ is non-increasing.
Theorem 3.2.
[…]
Moreover, suppose that the ordering condition $b_{1}(x)\leq b_{2}(x)$ holds for every $x\in\mathbb{R}$ and that at least one of the functions $b_{1},b_{2}$ is monotonically non-decreasing. Let
(27) $X^{i}_{t}=g_{i}(t)+\int_{0}^{t}K(t-s)b_{i}(X^{i}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(X^{i}_{s})\,\mathrm{d}B_{s}$
for $i=1,2$ be the unique strong solution. Then $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\;t>0]=1$.
Proof.
Let us first note that the equations of interest have weak solutions $X^{1},X^{2}$ with sample paths in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ due to [1, Theorem 1.2]. Moreover, it follows from [1, Theorem 6.1 (ii)] that $X^{1}-g_{1},X^{2}-g_{2}$ have a modification with continuous sample paths. Hence, $X^{1},X^{2}$ are continuous on $(0,\infty)$. Additionally, these solutions satisfy for any $T>0,i=1,2$, $\int_{0}^{T}\mathbb{E}[|X_{t}^{i}|^{2}]\,\mathrm{d}t<\infty$ c.f. Proposition A.1. The Yamada-Watanabe type of approach as given in [3, Proposition B.3] yields pathwise uniqueness and hence the Yamada-Watanabe-Engelbert theorem (see Kurtz [29, Theorem 1.5]) implies strong existence. Without loss of generality, we may assume that $b_{1}$ is non-decreasing. Moreover, since $K$ is non-decreasing and $g_{1},g_{2}$ are admissible,
$g_{2}(t)-g_{1}(t)=\widetilde{g}_{2}(t)-\widetilde{g}_{1}(t)+\int_{0}^{t}K(t-s)(h_{2}(s)-h_{1}(s))\,\mathrm{d}s\geq 0$
is non-decreasing on $(0,\infty)$. Since $K\in W_{\mathrm{loc}}^{1,2}(\mathbb{R}_{+})$, the classical and stochastic version of Fubini’s theorem [37] and [12, Theorem 8.2], imply that
$X^{1}_{t}-X^{2}_{t}$ $=K(0)\int_{0}^{t}\sigma(X^{1}_{s})-\sigma(X^{2}_{s})\,\mathrm{d}B_{s}+\int_{0}^{t}\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{1}_{r})-b_{2}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\quad+\int_{0}^{t}\int_{0}^{s}K^{\prime}(s-r)\left(\sigma(X^{1}_{r})-\sigma(X^{2}_{r})\right)^{2}\mathrm{d}B_{r}\,\mathrm{d}s+K(0)\int_{0}^{t}b_{1}(X^{1}_{s})-b_{2}(X^{2}_{s})\,\mathrm{d}s$
$\quad+(g_{1}-g_{2})(t)$
which is a continuous semimartingale. Hence we may follow the same approximation procedure as presented in [27, Chapter VI, Theorem 1.1]. That is, let $(a_{n})_{n\in\mathbb{N}_{0}}\subset(0,1]$ be a sequence with $a_{0}=1$, $a_{n}<a_{n-1}$ and $\lim_{n\longrightarrow\infty}a_{n}=0$ such that $\int_{a_{n}}^{a_{n-1}}x^{-2\eta}\,\mathrm{d}x=n.$ Then, we can construct a sequence $(\varphi_{n})_{n\in\mathbb{N}}\subset C^{2}(\mathbb{R})$ of twice continuously differentiable functions such that
  • $\varphi_{n}^{\prime}(x)=0$ for $x\leq a_{k}$, $\varphi_{n}^{\prime}(x)=1$ for $x\geq a_{n-1}$ and $\varphi_{n}^{\prime}(x)\in[0,1]$ for $a_{n}<x<a_{n-1},$
  • $\varphi_{n}^{\prime\prime}(x)\in[0,\frac{2}{n}x^{-2\eta}]$ for $a_{n}<x<a_{n-1},$
and $\varphi_{n}(x)\nearrow x_{+}$ as $n\longrightarrow\infty$. Here, $x_{+}\coloneqq\max\{x,0\}$ denotes the positive part of $x$. Applying Itô ’s formula yields $\varphi_{n}(X^{1}_{t}-X^{2}_{t})=I_{1}+\dots+I_{6},$ where
$I_{1}$ $=K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})(\sigma(X^{1}_{s})-\sigma(X^{2}_{s}))\,\mathrm{d}B_{s},$
$I_{2}$ $=\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{1}_{r})-b_{2}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s,$
$I_{3}$ $=\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(\sigma(X^{1}_{r})-\sigma(X^{2}_{r}))\,\mathrm{d}B_{r}\,\mathrm{d}s,$
$I_{4}$ $=K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})(b_{1}(X^{1}_{s})-b_{2}(X^{2}_{s}))\,\mathrm{d}s,$
$I_{5}$ $=\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\,\mathrm{d}(g_{1}-g_{2})(s),$
$I_{6}$ $=\frac{K(0)^{2}}{2}\int_{0}^{t}\varphi_{n}^{\prime\prime}(X^{1}_{s}-X^{2}_{s})(\sigma(X^{1}_{s})-\sigma(X^{2}_{s}))^{2}\,\mathrm{d}s.$
It is a routine matter to show $\mathbb{E}[I_{1}]=\mathbb{E}[I_{3}]=0.$ Since $g_{2}-g_{1}$ is non-decreasing and $\varphi_{n}^{\prime}\geq 0,$ we deduce $I_{5}\leq 0.$ Moreover, by using the construction of $\varphi_{n},$
$I_{6}\lesssim\int_{0}^{t}\varphi_{n}^{\prime\prime}(X^{1}_{s}-X^{2}_{s})|X^{1}_{s}-X^{2}_{s}|^{2\eta}\,\mathrm{d}s\lesssim\frac{1}{n}.$
By using $K(0)>0$, $b_{1}(t,x)\leq b_{2}(t,x)$ and the Lipschitz continuity of $b_{1}$, we obtain
$I_{4}$ $=K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\mathbbm{1}_{\{X^{1}_{s}\geq X^{2}_{s}\}}(b_{1}(X^{1}_{s})-b_{1}(X^{2}_{s}))\,\mathrm{d}s$
$\quad+K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\mathbbm{1}_{\{X^{1}_{s}\geq X^{2}_{s}\}}(b_{1}(X^{2}_{s})-b_{2}(X^{2}_{s}))\,\mathrm{d}s$
$\lesssim K(0)\int_{0}^{t}(X^{1}_{s}-X^{2}_{s})_{+}\,\mathrm{d}s.$
In the same spirit and by using $K^{\prime}\geq 0$ and $b_{1}$ is non-decreasing, we obtain
$I_{2}$ $=\int_{0}^{t}\varphi^{\prime}_{n}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{1}_{r})-b_{1}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\quad+\int_{0}^{t}\varphi^{\prime}_{n}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{2}_{r})-b_{2}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\leq\int_{0}^{t}\varphi^{\prime}_{n}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)\mathbbm{1}_{\{X^{1}_{r}\geq X^{2}_{r}\}}(b_{1}(X^{1}_{r})-b_{1}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\int_{0}^{s}K^{\prime}(s-r)(X^{1}_{r}-X^{2}_{r})_{+}\,\mathrm{d}r\,\mathrm{d}s$
$=\int_{0}^{t}\int_{r}^{t}K^{\prime}(s-r)(X^{1}_{r}-X^{2}_{r})_{+}\,\mathrm{d}s\,\mathrm{d}r$
$=\int_{0}^{t}(K(t-r)-K(0))(X^{1}_{r}-X^{2}_{r})_{+}\,\mathrm{d}r.$
Moreover, $\mathbb{E}[\varphi_{n}(X^{1}_{t}-X^{2}_{t})]$ converges to $\mathbb{E}[(X^{1}_{t}-X^{2}_{t})_{+}]$ as $n\longrightarrow\infty$, thanks to the monotone convergence theorem. Consequently, we arrive at
$\mathbb{E}[(X^{1}_{t}-X^{2}_{t})_{+}]\lesssim\int_{0}^{t}K(t-s)\mathbb{E}\left[(X^{1}_{s}-X^{2}_{s})_{+}\right]\,\mathrm{d}s.$
Finally, Gronwall’s inequality for Volterra equations see e.g. [10, Lemma A.1] implies $\mathbb{E}[(X^{1}_{t}-X^{2}_{t})_{+}]=0$ and by the continuity of the sample paths we conclude $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\,t>0]=1$. ∎
It should be possible to extend the above also to the case of time-dependent coefficients. However, as the current formulation is sufficient to cover all cases of interest, we leave such an extension to the interested reader.
Remark 3.3.
The proof shows that the comparison principle also holds for the kernel $\widetilde{K}\coloneqq-K$ under the new assumption that $b_{1}(x)\geq b_{2}(x)$ and at least one of $b_{1},b_{2}$ is non-increasing.
The fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>3/2$ satisfies the assumptions stated in Theorem 3.2. The case $\alpha>1$ can be now obtained by approximation similar to Theorem 2.7. Since for such cases, the uniqueness is not guaranteed, we show that for any two weak solutions, there exists a coupling that satisfies the comparison principle.
Theorem 3.4.
Let $(g_{i},K,b_{i},\sigma)$, $i=1,2,$ satisfy Assumption 2.2 with $b_{1}\leq b_{2}$ and $b_{1}$ or $b_{2}$ is nondecreasing. Suppose that $K\in C^{1}((0,\infty))$ with $K^{\prime}\geq 0.$ Let $g_{1},g_{2}$ be admissible such that both satisfy (9) for some $\delta\in[0,1/2)$. Then for any two weak continuous solutions $(X^{i},(\Omega^{i},\mathcal{F}^{i},\mathbb{F}^{i},\mathbb{P}^{i}),B^{i})$, $i=1,2,$ a weak continuous solution of (27), there exists a weak continuous solution $((\widehat{X}^{1},\widehat{X}^{2}),$ $(\widehat{\Omega},\widehat{\mathcal{F}},\widehat{\mathbb{F}},\widehat{\mathbb{P}}),\widehat{B})$ of the two dimensional stochastic Volterra equation
$\begin{pmatrix}\widehat{X}^{1}_{t}\\ \widehat{X}^{2}_{t}\end{pmatrix}=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(\widehat{X}^{1}_{s})\\ b_{2}(\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(\widehat{X}^{1}_{s})\\ \sigma(\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}\widehat{B}_{s}$
such that $\mathcal{L}(\widehat{X}^{i})=\mathcal{L}(X^{i})$, $i=1,2,$ and ^$\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$.
Proof.
We choose $(b_{1,n})_{n\in\mathbb{N}}$, $(b_{2,n})_{n\in\mathbb{N}}$,$(\sigma_{n})_{n\in\mathbb{N}}$ as in the proof of Theorem 2.7 (without time dependence) and can assume that $b_{1,n}$ or $b_{2,n}$ is non-decreasing. Moreover, define $K_{n}(t)=K(t+\frac{1}{n}).$ Clearly, $K_{n}\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and $K_{n}\longrightarrow K$ in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+}).$ Let $(X^{n})_{n\in\mathbb{N}}$, $(Y^{n})_{n\in\mathbb{N}}$ be the family of continuous processes constructed by $(g_{1},K_{n},b_{1,n},\sigma_{n})$ and $(g_{2},K_{n},b_{2,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. Then, by Theorem 3.2, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},\;t>0].$ Using Propsition A.2, we obtain for some $p\geq 2$ suitably chosen
$\sup_{n\in\mathbb{N}}\mathbb{E}[|(X^{n}_{t}-g_{1}(t))-(X^{n}_{s}-g_{2}(s))|^{p}]+\sup_{n\in\mathbb{N}}\mathbb{E}[|(Y^{n}_{t}-g_{1}(t))-(Y^{n}_{s}-g_{2}(s))|^{p}]\leq c|t-s|^{1+\theta},$
where $\theta>0$ and $c$ is a constant that only depends on $p,C,T,K_{1}$. Define for every $n\in\mathbb{N}$ the continuous semimartingale
$Z^{n}_{t}=\int_{0}^{t}\begin{pmatrix}b_{1,n}(X^{n}_{s})\\ b_{2,n}(Y^{n}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}\begin{pmatrix}\sigma_{n}(X^{n}_{s})\\ \sigma_{n}(Y^{n}_{s})\end{pmatrix}\mathrm{d}B_{s}.$
Then we also obtain for some $p\geq 2$ suitably chosen
$\sup_{n\in\mathbb{N}}\mathbb{E}[|Z^{n}_{t}-Z^{n}_{s}|^{p}]\leq c|t-s|^{1+\theta},$
where $\theta>0$ and $c$ is a constant that only depends $C,T,K_{1}$ and $p$. Hence, we conclude by Kolmogorov’s tightness criterion that $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ is a tight sequence of continuous processes in $\mathbb{R}^{4}$ and so has a weakly convergent subsequence $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ on the path space $C(\mathbb{R}_{+};\mathbb{R}^{4})$. Define the continuous mapping $\Phi\colon C(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})\longrightarrow C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$ by
$\Phi(x,y,z_{1},z_{2})=(x+g_{1},y+g_{2},z_{1},z_{2}).$
By the continuous mapping theorem, $(X^{n},Y^{n},Z^{n})=\Phi(X^{n}-g_{1},Y^{n}-g_{2},Z)\Longrightarrow\widetilde{\mathbb{P}}$ on $C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$. As in the proof of Theorem 2.7 Step 2, we can show ¯$\overline{\mathbb{P}}[x(t)\leq y(t),\;t>0]=1$ where $(x,y)$ denotes the projection to the first two coordinates of the coordinate process. Note that $C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$ is continuously embedded into $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+};\mathbb{R}^{2})\times D(\mathbb{R}_{+};\mathbb{R}^{2})$ where we have equipped the Skorokhod space $D(\mathbb{R}_{+};\mathbb{R}^{2})$ with the usual Skorokhod topology, see e.g. [11, Section 12]. Hence by passing to a subsequence $(X^{n},Y^{n},Z^{n})\Longrightarrow(\widehat{X},\widehat{Y},\widehat{Z})$ weakly in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+};\mathbb{R}^{2})\times D(\mathbb{R}_{+};\mathbb{R}^{2})$. Using [1, Theorem 1.6] yields that $(\widehat{X},\widehat{Y},\widehat{Z})$ is a weak solution for the data $((g_{1},g_{2}),K,$ $(b_{1},b_{2}),(\sigma,\sigma)))$ and satisfies ^$\widehat{\mathbb{P}}[\widehat{X}_{t}\leq\widehat{Y}_{t},t>0]=1.$
Appendix A Existence and uniqueness for Lipschitz coefficients
Below we prove a result on the existence and uniqueness of solutions for a stochastic Volterra equation with Lipschitz continuous coefficients and a general driving force $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in[2,\infty]$. The following extends [10, Theorem 4.8] to time-dependent coefficients while, for the sake of simplicity, we stay in the finite-dimensional case.
Proposition A.1.
Let $K\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and let $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ be measurable such that
$|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq C|x-y|,$
and
(28) $|b(t,x)|+|\sigma(t,x)|\leq C(1+|x|)$
hold for a.a. $t\geq 0$ and all $x,y\in\mathbb{R}$ with some constant $C>0$. Then for each $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in[2,\infty]$, there exists a unique strong solution of the equation
(29) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}s.$
This solution satisfies $X\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L_{\mathrm{loc}}^{q}(\mathbb{R}_{+}))$ for each $p\in[2,q]$ with $p<\infty$. To be more precise, $\|X\|_{L^{p}(\Omega;L^{q}([0,T]))}\leq c$, where $c>0$ is a constant that only depends on $K,T,p,C$ [^3] The dependence on $C$ only comes through the growth estimate (28)..
Proof.
Below we follow the steps outlined in [10, Theorem 4.8]. Fix $\lambda<0$ and define the functions $K_{\lambda}(t)=\mathrm{e}^{\lambda t}K(t)$ and $g_{\lambda}(t)=\mathrm{e}^{\lambda t}g(t)$. Let $X_{t}(g)$ be a solution of (29). Then $Y^{\lambda}_{t}(g)=\mathrm{e}^{\lambda t}X_{t}(g)$ satisfies
(30) $Y_{t}^{\lambda}(g)=g_{\lambda}(t)+\int_{0}^{t}K_{\lambda}(t-s)b_{\lambda}(s,Y_{s}^{\lambda}(g))\,\mathrm{d}s+\int_{0}^{t}K_{\lambda}(t-s)\sigma_{\lambda}(s,Y_{s}^{\lambda}(g))\,\mathrm{d}B_{s},$
where $b_{\lambda}(s,x)=\mathrm{e}^{\lambda s}b(s,\mathrm{e}^{-\lambda s}x)$ and $\sigma_{\lambda}(s,x)=\mathrm{e}^{\lambda s}\sigma(s,\mathrm{e}^{-\lambda s}x)$. Conversely, let $Y_{\lambda}$ by a solution of (30), then $X^{\lambda}_{t}(g)=\mathrm{e}^{-\lambda t}Y_{\lambda}(t;g)$ satisfies (29). Therefore, it suffices to prove existence and uniqueness for (30). We solve (30) by a fixed point argument for the case $p=q$. Fix $T>0$, then $g|_{[0,T]}\in L^{q}([0,T]).$ For a given $\lambda<0$, define
$\mathcal{T}_{\lambda}(X;g)(t)\coloneqq g_{\lambda}(t)+\int_{0}^{t}K_{\lambda}(t-s)b_{\lambda}(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K_{\lambda}(t-s)\sigma_{\lambda}(s,X_{s})\,\mathrm{d}B_{s}.$
Then (30) is equivalent to $X^{\lambda}=\mathcal{T}_{\lambda}(X^{\lambda}(g))$. Below we show that $\mathcal{T_{\lambda}}(\cdot,g)$ is a contraction on $L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$ when $\lambda<0$ is small enough. For brevity, we let $\|\cdot\|_{p}=\|\cdot\|_{L^{p}(\Omega;L^{p}([0,T])}$. Take $X\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$, then
$\|\mathcal{T}_{\lambda}(X;g)\|_{p}\leq\|g\|_{L^{p}([0,T])}+\|K_{\lambda}\ast b_{\lambda}(\cdot,X)\|_{p}+\|K_{\lambda}\ast\sigma_{\lambda}(\cdot,X)\,\mathrm{d}B\|_{p}.$
For the drift we find, using Young’s inequality
$\|K_{\lambda}\ast b_{\lambda}(\cdot,X)\|_{p}$ $=\left\|\|K_{\lambda}\ast b_{\lambda}(\cdot,X)\|_{L^{p}([0,T])}\right\|_{L^{p}(\Omega)}$
$\leq\|K_{\lambda}\|_{L^{1}([0,T])}\,\left\|\|b_{\lambda}(\cdot,X_{s})\|_{L^{p}([0,T])}\right\|_{L^{p}(\Omega)}$
$\leq C\|K_{\lambda}\|_{L^{1}([0,T])}\|1+X\|_{p}.$
For the stochastic convolution, we find using BDG and Young’s inequality
$\|K_{\lambda}\ast\sigma_{\lambda}(\cdot,X)\,\mathrm{d}B\|_{p}$ $\leq c_{p}\left(\mathbb{E}\left[\|K_{\lambda}^{2}\ast\sigma_{\lambda}(\cdot,X)^{2}\|_{L^{p/2}([0,T])}^{p/2}\right]\right)^{1/p}$
$\leq c_{p}\|K^{2}_{\lambda}\|_{L^{1}([0,T])}^{1/2}\left(\mathbb{E}\left[\|\sigma_{\lambda}(\cdot,X)^{2}\|_{L^{p/2}([0,T])}^{p/2}\right]\right)^{1/p}$
$\leq c_{p}\|K_{\lambda}\|_{L^{2}([0,T])}\left(\mathbb{E}\left[\|C(1+|X|)\|_{L^{p}([0,T])}^{p}\right]\right)^{1/p}$
$\leq Cc_{p}\|K_{\lambda}\|_{L^{2}([0,T])}\|1+X\|_{p}.$
Hence, $\mathcal{T}_{\lambda}(\cdot;g)$ leaves the space $L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$ invariant. In the same way, if $X,Y\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$, we obtain
$\|\mathcal{T}_{\lambda}(X;g)-\mathcal{T}_{\lambda}(Y;g)\|_{p}$ $\leq\|K_{\lambda}\ast(b_{\lambda}(\cdot,X)-b_{\lambda}(\cdot,Y))\|_{p}+\|K_{\lambda}\ast(\sigma_{\lambda}(\cdot,X)-\sigma_{\lambda}(\cdot,Y))\|_{p}$
$\leq C\max\{\|K_{\lambda}\|_{L^{1}([0,T])},c_{p}\|K_{\lambda}\|_{L^{2}([0,T])}\}\|X-Y\|_{p}.$
Since by dominated convergence $\|K_{\lambda}\|_{L^{1}([0,T])},\ \|K_{\lambda}\|_{L^{2}([0,T])}\longrightarrow 0$ as $\lambda\longrightarrow-\infty$, we can choose $\lambda<0$ sufficiently small such that $\mathcal{T}_{\lambda}(\cdot;g)$ is a contraction and, hence, has a unique fixed point $Y_{\lambda}(\cdot;g)$ which is the unique solution of (30). The same fixpoint argument can be executed in the space $L^{p}(\Omega.\mathcal{F},\mathbb{P};L^{\infty}([0,T]))$, provided that $g\in L_{\mathrm{loc}}^{\infty}(\mathbb{R}_{+})$. Using similar bounds as above, it is easy to derive, for $p\in[2,q]$ and a.a. $t\geq 0$
$\mathbb{E}[|X_{t}|^{p}]\leq|g(t)|^{p}+A_{0}+A_{1}\int_{0}^{t}K(t-s)^{2}\mathbb{E}[|X_{s}|^{p}]\,\mathrm{d}s,$
where $A_{0},A_{1}$ are constants that only depends on $K,C,T,p$. Using the Volterra type Gronwall inequality (see e.g. [10, Lemma A.1] yields for a.a. $t\geq 0$
(31) $\mathbb{E}[|X_{t}|^{p}]\leq|g(t)|^{p}+A_{0}+\int_{0}^{t}R(t-s)(|g(s)|^{p}+A_{0})\,\mathrm{d}s,$
where $R\in L^{1}([0,T])$ denotes the resolvent of the second kind of $A_{1}K^{2}$. Since $|g|^{p}\in L^{q/p}([0,T])$, the right-hand side of (31) belongs to $L^{q/p}([0,T]).$ Thus, $\|X\|_{L^{p}(\Omega;L^{q}([0,T]))}\leq c$, where $c$ is a constant that only depends on $K,C,T,p$. ∎
Proposition A.2.
Suppose that $(g,K,b,\sigma)$ satisfies Assumption 2.2. Let $X$ be any solution of (29).
  1. (i)
    Under assumption 2.2.(i), $X-g$ has, for each $\theta<\gamma$, a continuous modification with locally $\theta$-Hölder continuous sample paths.
  2. (ii)
    Under assumption 2.2.(ii), $X-g$ has, for each $\theta<\vartheta$ with
    $\vartheta\coloneqq\gamma+\frac{1}{2}\frac{\eta}{2+\eta}-\frac{2\xi}{q}\left(1+\frac{1}{\eta}\right),$
    a continuous modification with locally $\theta$-Hölder continuous sample paths.
Proof.
Firstly, let us consider the case (ii). Since $q\geq\xi\left(1+\frac{2}{\eta}\right)$, we find $p\coloneqq\frac{\eta q}{(2+\eta)\xi}\geq 2$. Then, for $0<s<t\leq T,$
$|(X_{t}-g(t))-(X_{s}-g(s))|^{p}$
$\quad\lesssim\left|\int_{s}^{t}K(t-r)b(r,X_{r})\,\mathrm{d}r\right|^{p}+\left|\int_{0}^{s}(K(t-r)-K(s-r))b(r,X_{r})\,\mathrm{d}r\right|^{p}$
$\qquad+\left|\int_{s}^{t}K(t-r)\sigma(r,X_{r})\,\mathrm{d}B_{r}\right|^{p}+\left|\int_{0}^{s}(K(t-r)-K(s-r))\sigma(r,X_{r})\,\mathrm{d}B_{r}\right|^{p}$
$\quad\eqqcolon I_{1}+I_{2}+I_{3}+I_{4}$
An application of Jensen’s and Hölder’s inequality yields
(32) $\mathbb{E}[I_{1}]$ $\leq\left(\int_{s}^{t}|K(t-r)|\,\mathrm{d}r\right)^{p-1}\mathbb{E}\left[\int_{s}^{t}|K(t-r)||b(r,X_{r})|^{p}\,\mathrm{d}r\right]$
$\leq C^{p}\|K\|_{L^{1}([0,t-s])}^{p-1}\mathbb{E}\left[\int_{s}^{t}|K(t-r)|(1+|X_{r}|)^{\xi p}\,\mathrm{d}r\right]$
$\leq C^{p}\|K\|_{L^{2+\eta}([0,t-s])}^{p-1}(t-s)^{\frac{1+\eta}{2+\eta}(p-1)}\|K\|_{L^{2+\eta}([0,t-s])}\mathbb{E}\left[\left(\int_{s}^{t}(1+|X_{r}|)^{\xi p\frac{2+\eta}{1+\eta}}\,\mathrm{d}r\right)^{\frac{1+\eta}{2+\eta}}\right]$
$\leq C^{p}c(T,\eta)^{p}(t-s)^{\gamma p+\frac{1+\eta}{2+\eta}(p-1)}\int_{0}^{T}\mathbb{E}\left[(1+|X_{r}|)^{q}\right]\mathrm{d}r$
where the finiteness of the expectation follows from Proposition A.1. Analogously we obtain from Itô ’s isometry, Jensen’s and Hölder’s inequality
$\mathbb{E}[I_{3}]$ $=\mathbb{E}\left[\left(\int_{s}^{t}K(t-r)^{2}\sigma(r,X_{r})^{2}\,\mathrm{d}r\right)^{\frac{p}{2}}\right]$
(33) $\leq C^{p}\left(\int_{s}^{t}K(t-r)^{2}\,\mathrm{d}r\right)^{\frac{p}{2}-1}\mathbb{E}\left[\int_{s}^{t}K(t-r)^{2}(1+|X_{r}|)^{\xi p}\,\mathrm{d}r\right]$
$\leq C^{p}\|K\|_{L^{2+\eta}([0,t-s])}^{p-2}(t-s)^{\frac{\eta}{2+\eta}\left(\frac{p}{2}-1\right)}\|K\|_{L^{2+\eta}([0,t-s])}^{2}\mathbb{E}\left[\left(\int_{s}^{t}(1+|X_{r}|)^{\xi p\frac{2+\eta}{\eta}}\,\mathrm{d}r\right)^{\frac{2}{2+\eta}}\right]$
$\leq C^{p}c(T,\eta)^{p}(t-s)^{\gamma p+\frac{\eta}{2+\eta}\left(\frac{p}{2}-1\right)}\int_{0}^{T}\mathbb{E}\left[(1+|X_{r}|)^{q}\right]$
The same estimate can be derived for $\mathbb{E}[I_{2}],\mathbb{E}[I_{3}].$ Thus, using $p\geq 2$, gives $(1+\eta)(p-1)\geq\eta\left(\frac{p}{2}-1\right)$ and hence we obtain
$\left(\mathbb{E}[|(X_{t}-g(t))-(X_{s}-g(s))|^{p}]\right)^{\frac{1}{p}}\leq c(t-s)^{\frac{1}{p}+\gamma+\frac{\eta}{2+\eta}\left(\frac{1}{2}-\frac{1}{p}\right)-\frac{1}{p}}=c(t-s)^{\frac{1}{p}+\vartheta},$
where $c>0$ is a constant that only depends on $p$,$\eta$, $C$, $T$ but not on $s$ or $t$, and the last equality follows from the particular form of $p$. Note that $\vartheta>0$ if and only if
$q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}$
which is satisfied by Assumption 2.2. The Hölder continuity now follows from the Kolmogorov-Chentsov theorem. For the case of Assumption 2.2.(i), we let $p\geq 2$ be arbitrary. By following similar arguments to (32) and (33) and using $\operatorname*{ess\,sup}_{r\in[0,T]}\mathbb{E}[|X_{r}|^{p}]<\infty$ due to Proposition A.1, we find
$\mathbb{E}[|(X_{t}-g(t))-(X_{s}-g(s)|^{p}]\lesssim(t-s)^{p\gamma}.$
This proves the assertion. ∎
If $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$, then the above result essentially coincides with [4]. For the case of Assumption 2.2.(ii), it gives a new way of proving the existence of continuous solutions.
Appendix B Splitting Method for Stochastic Volterra Equations
Given any solution $X$ of
$X_{t}=g(t)+\int_{0}^{t}K(t-s)[b(s,X_{s})\,\mathrm{d}s+\sigma(s,X_{s})\,\mathrm{d}B_{s}],$
where $g\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})\cap C((0,\infty))$ and $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ are measurable such that there exists $C\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ with
(34) $|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq C(t)|x-y|,\quad|b(t,x)|^{2}+|\sigma(t,x)|^{2}\leq C^{2}(t)(1+|x|^{2}),$
for a.a. $t\geq 0$ and $x,y\in\mathbb{R}$. Let $T>0$, $N\in\mathbb{N}$ be fixed and $t_{k}=kT/N,$ $k\in\{0,1,\dots,N\},$ be an equidistant grid of $[0,T]$ with step size $T/N.$ Let us denote by $(\widehat{X}_{t})_{t\in(0,T]}$ its approximation
$\widehat{X}_{t}$ $=g(t)+\sum_{k=1}^{N}\mathbbm{1}_{[t_{k},T]}(t)\left(K(t-t_{k})\int_{t_{k-1}}^{t_{k}}[b(s,\xi^{k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{k}_{s})\,\mathrm{d}B_{s}]\right)$
(35) $=g(t)+\int_{0}^{t_{\eta(t)}}K(t-t_{\eta(s)})[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}],$
where $\eta(t)=k$ if $t\in[t_{k},t_{k+1})$ and $\xi_{t}\coloneqq\sum_{k=1}^{N}\mathbbm{1}_{[t_{k-1},t_{k})}(t)\xi_{t}^{k}$ with
$\xi^{k}_{t}=\widehat{X}_{t_{k}-}+\int_{t_{k-1}}^{t}K(0_{+})[b(s,\xi^{k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{k}_{s})\,\mathrm{d}B_{s}],\quad t\in[t_{k-1},t_{k}).$
Denote by
$\omega_{f,T}(\delta)\coloneqq\sup_{s,t\in[0,T]:|s-t|\leq\delta}|f(s)-f(t)|$
the $\delta$-modulus of continuity of a function $f\colon[0,T]\longrightarrow\mathbb{R}$ and by $\|f\|_{[0,T]}\coloneqq\sup_{t\in[0,T]}|f(t)|$ its sup-norm. Below we extend the approximation [6, Lemma 3.1, Proposition 3.1] to the case of time-dependent coefficients and possibly singular $g$. As a first step, we prove a refined bound on the approximation where $g$ is regular.
Proposition B.1.
Suppose that $K$ is nonnegative, nonincreasing and continuous on $\mathbb{R}_{+}$, and that $g\in C(\mathbb{R}_{+})$ Then, for every $T>0$, it holds that
(36) $\sup_{t\in[0,T]}\mathbb{E}[|\widehat{X}_{t}|^{2}+|\xi_{t}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}$
and
$\sup_{t\in[0,T]}\left(\mathbb{E}[|\widehat{X}_{t}-\xi_{t}|^{2}]+\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]\right)\\ \lesssim\omega^{2}_{g,T}\left(\frac{T}{N}\right)+\left(1+\|g\|_{[0,T]}^{2}\right)\left(\omega^{2}_{K,T}\left(\frac{T}{N}\right)+\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s\right).$
Proof.
Let $k\in\{1,\dots,N\}$ and $t\in[t_{k-1},t_{k})$. By using (2),
$|\xi_{t}|^{2}\lesssim|\widehat{X}_{t_{k}-}|^{2}+\left|\int_{t_{k-1}}^{t}b(s,\xi_{s})\,\mathrm{d}s\right|^{2}+\left|\int_{t_{k-1}}^{t}\sigma(s,\xi_{s})\,\mathrm{d}B_{s}\right|^{2}.$
An application of Jensen’s inequality, Itô ’s isometry and the growth condition on $b,\sigma$ yields
(37) $\mathbb{E}[|\xi_{t}|^{2}]\lesssim\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]+\int_{t_{k-1}}^{t}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s.$
Thus by Gronwall’s inequality, $\mathbb{E}[|\xi_{t}|^{2}]\lesssim\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]+\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s$ for $t\in[t_{k-1},t_{k})$. By letting $t\nearrow t_{k}$ in (35) and using $\lim_{t\nearrow t_{k}}\eta(t)=k-1$, we obtain
$\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]$ $\lesssim g(t_{k})^{2}+\int_{0}^{t_{k-1}}K(t-t_{\eta(s)})^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$
$\lesssim\|g\|_{[0,T]}^{2}+\int_{0}^{t_{k-1}}C(s)^{2}\,\mathrm{d}s+\sum_{j=1}^{k-1}\int_{t_{j-1}}^{t_{j}}\left(\mathbb{E}[|\widehat{X}_{t_{j}-}|^{2}]+\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s\right)C(s)^{2}\,\mathrm{d}s$
$\lesssim 1+\|g\|_{[0,T]}^{2}+\sum_{j=1}^{k-1}\mathbb{E}[|\widehat{X}_{t_{j}-}|^{2}]\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s$
where we have used $\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s\leq\|C\|_{L^{2}([0,T]}^{2}$. By virtue of the discrete Gronwall inequality [13] and the standard inequality $1+x\leq\mathrm{e}^{x}$, we obtain $\sup_{1\leq k\leq N}\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}<\infty.$ Consequently, $\sup_{t\in[0,T]}\mathbb{E}[|\xi_{t}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}$ and so by construction of the approximation we arrive at (36). Let $t\in[t_{k-1},t_{k})$. We have $|\widehat{X}_{t}-\xi_{t}|^{2}\lesssim|\widehat{X}_{t}-\widehat{X}_{t_{k}-}|^{2}+|\widehat{X}_{t_{k}-}-\xi_{t}|^{2}$. For the second term, using (37) we arrive at
$\mathbb{E}[|\widehat{X}_{t_{k}-}-\xi_{t}|^{2}]\lesssim\int_{t_{k-1}}^{t}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s\lesssim\left(1+\|g\|_{[0,T]}^{2}\right)\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s.$
For the first difference, we have
$\widehat{X}_{t}-\widehat{X}_{t_{k-}}=g(t)-g(t_{k})+\int_{0}^{t_{k-1}}(K(t-t_{\eta(s)})-K(t_{k}-t_{\eta(s)}))[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}]$
and consequently
$\mathbb{E}[|\widehat{X}_{t}-\widehat{X}_{t_{k}-}|^{2}]$
$\quad\lesssim|g(t)-g(t_{k})|^{2}+\int_{0}^{t_{k-1}}(K(t-t_{\eta(s)})-K(t_{k}-t_{\eta(s)}))^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{t}|^{2}])\,\mathrm{d}s$
$\quad\lesssim\omega^{2}_{g,T}\left(\frac{T}{N}\right)+\left(1+\|g\|_{[0,T]}^{2}\right)\omega^{2}_{K,T}\left(\frac{T}{N}\right).$
Finally, it remains to bound the difference
$\widehat{X}_{t}-X_{t}$ $=\int_{0}^{t}K(t-s)[b(s,\xi_{s})-b(s,X_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})-\sigma(s,X_{s})\,\mathrm{d}B_{s}]$
$\quad+\int_{0}^{t_{\eta(t)}}(K(t-t_{\eta(s)})-K(t-s))[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}]$
$\quad-\int_{t_{\eta(t)}}^{t}K(t-s)[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}].$
Using the linear growth and Lipschitz condition (34), we obtain
$\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]$ $\lesssim\int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\xi_{s}-X_{s}|^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t_{\eta(t)}}(K(t-t_{\eta(s)})-K(t-s))^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$
$\quad+\int_{t_{\eta(t)}}^{t}K(t-s)^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s.$
For the first term we use $|\xi_{s}-X_{s}|^{2}\lesssim|\xi_{s}-\widehat{X}_{s}|^{2}+|\widehat{X}_{s}-X_{s}|^{2}$ to bound
$\ \int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\xi_{s}-X_{s}|^{2}]\,\mathrm{d}s$
$\lesssim\int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\xi_{s}-\widehat{X}_{s}|^{2}]\,\mathrm{d}s+\int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\widehat{X}_{s}-X_{s}|^{2}]\,\mathrm{d}s$
$\lesssim\omega^{2}_{g,T}\left(\frac{T}{N}\right)+\left(1+\|g\|_{[0,T]}^{2}\right)\left(\omega^{2}_{K,T}\left(\frac{T}{N}\right)+\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s\right)+\int_{0}^{t}C(s)^{2}\mathbb{E}[|\widehat{X}_{s}-X_{s}|^{2}]\,\mathrm{d}s.$
For the second term, we use (37) to find that
$\int_{0}^{t_{\eta(t)}}(K(t-t_{\eta(s)})-K(t-s))^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$ $\lesssim\left(1+\|g\|_{[0,T]}^{2}\right)\omega^{2}_{K,T}\left(\frac{T}{N}\right).$
Finally, the last term can be bounded by
$\int_{t_{\eta(t)}}^{t}K(t-s)^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$ $\lesssim\left(1+\|g\|_{[0,T]}^{2}\right)\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s.$
Hence, collecting all estimates, Gronwall’s inequality yields the assertion. ∎
Next, we prove the convergence of the approximation also in the case where $g$ may be singular in $t=0$, provided that the singularity is not too rough.
Proposition B.2.
Let $g\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})\cap C((0,\infty))$ be such that there exists $\delta>0$ with
$\int_{0}^{T}\overline{g}(s)^{2+\delta}C(s)^{2}\,\mathrm{d}s<\infty$
where $\overline{g}(s)\coloneqq\sup_{s\leq t\leq T}|g(t)|$. Then, for every $t\in(0,T]$,
$\lim_{N\longrightarrow\infty}\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]=0.$
Proof.
Fix $\varepsilon>0$ and define $g_{\varepsilon}\colon\mathbb{R}_{+}\longrightarrow\mathbb{R}$ by $g_{\varepsilon}(t)=g(t+\varepsilon).$ Moreover, let $X^{\varepsilon}$ be the strong solution of
$X^{\varepsilon}_{t}=g_{\varepsilon}(t)+\int_{0}^{t}K(t-s)[b(s,X^{\varepsilon}_{s})\,\mathrm{d}s+\sigma(s,X^{\varepsilon}_{s})\,\mathrm{d}B_{s}].$
Denote by $\widehat{X}^{\varepsilon}$ its approximation and by $\xi^{\varepsilon}$ the process that arises in its construction. Then $|\widehat{X}_{t}-X_{t}|\leq|\widehat{X}_{t}-\widehat{X}^{\varepsilon}_{t}|+|\widehat{X}^{\varepsilon}_{t}-X^{\varepsilon}_{t}|+|X^{\varepsilon}_{t}-X_{t}|$. Proposition B.1 gives
$\mathbb{E}[|\widehat{X}^{\varepsilon}_{t}-X^{\varepsilon}_{t}|^{2}]\lesssim\omega^{2}_{g_{\varepsilon},T}\left(\frac{T}{N}\right)+\left(1+\|g_{\varepsilon}\|_{[0,T]}^{2}\right)\left(\omega^{2}_{K,T}\left(\frac{T}{N}\right)+\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s\right),$
and it is a routine to show $\mathbb{E}[|X^{\varepsilon}_{t}-X_{t}|^{2}]\lesssim|g_{\varepsilon}(t)-g(t)|^{2}.$ Furthermore, we claim that
(38) $\mathbb{E}[|\widehat{X}_{t}-\widehat{X}^{\varepsilon}_{t}|^{2}]\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{\eta(t)}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s,\qquad t\in(0,T].$
Indeed, the claim is certainly true on $(0,t_{1})$ due to $\widehat{X}_{t}^{\varepsilon}=g_{\varepsilon}(t)$ and $\widehat{X}_{t}=g(t)$. Suppose (38) were true on $(0,t_{k})$ and let $t\in[t_{k},t_{k+1})$ then using (34) we arrive at
$\mathbb{E}[|\widehat{X}_{t}-\widehat{X}^{\varepsilon}_{t}|^{2}]$
$\quad\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{k}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\mathbb{E}[|\xi_{s}-\xi^{\varepsilon}_{s}|^{2}]\,\mathrm{d}s$
$\quad\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{k-1}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s+\int_{t_{k-1}}^{t_{k}}C(s)^{2}\mathbb{E}[|\xi_{s}-\xi^{\varepsilon}_{s}|^{2}]\,\mathrm{d}s.$
Finally, an application of Gronwall’s inequality to estimate $\mathbb{E}[|\xi_{s}-\xi^{\varepsilon}_{s}|^{2}]\lesssim|g(t_{k})-g_{\varepsilon}(t_{k})|^{2}$ yields (38). Consequently, the following less sharp version of (38) holds true
$\mathbb{E}[|\widehat{X}_{t}-\widehat{X}_{t}^{\varepsilon}|^{2}]$ $\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{N}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s$
$=|g(t)-g_{\varepsilon}(t)|^{2}+\int_{0}^{T}G^{N}_{\varepsilon}(s)C^{2}(s)\,\mathrm{d}s,$
where $G^{N}_{\varepsilon}(s)\coloneqq\sum_{j=1}^{N}\mathbbm{1}_{[t_{j-1},t_{j})}(s)|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}$. By continuity of $g$ on $(0,T]$, one can show that $G^{N}_{\varepsilon}(t)\overset{N\longrightarrow\infty}{\longrightarrow}|g(t)-g_{\varepsilon}(t)|^{2}$ for all $t\in(0,T]$ and fixed $\varepsilon>0$. Moreover, it holds that
$\int_{0}^{T}|G^{N}_{\varepsilon}(s)|^{1+\delta/2}C(s)^{2}\,\mathrm{d}s$ $=\sum_{j=1}^{N}\int_{t_{j-1}}^{t_{j}}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2+\delta}C(s)^{2}\,\mathrm{d}s$
$\lesssim\sum_{j=1}^{N}\int_{t_{j-1}}^{t_{j}}|g(t_{j})|^{2+\delta}C(s)^{2}\,\mathrm{d}s+\|g_{\varepsilon}\|_{[0,T]}^{2+\delta}\int_{0}^{T}C(s)^{2}\,\mathrm{d}s$
$\leq\int_{0}^{T}\overline{g}(s)^{2+\delta}C(s)^{2}\,\mathrm{d}s+\|g_{\varepsilon}\|_{[0,T]}^{2+\delta}\int_{0}^{T}C(s)^{2}\,\mathrm{d}s.$
Thus, $(G^{N}_{\varepsilon})_{N\in\mathbb{N}}$ is uniformly integrable. Moreover,
$\lim_{N\longrightarrow\infty}\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s=0.$
Indeed, $t\longmapsto\int_{0}^{t}C(s)^{2}\,\mathrm{d}s$ is an absolutely continuous function. Thus, for every $\widetilde{\varepsilon}>0$ there exists $\widetilde{\delta}>0$ such that for any $a,b>0$ with $b-a<\widetilde{\delta}$ holds $\int_{a}^{b}C(s)^{2}\,\mathrm{d}s<\widetilde{\varepsilon}$. Choosing $N$ large enough yields $t_{k}-t_{k-1}=\frac{1}{N}<\widetilde{\delta}$ and so $\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s<\widetilde{\varepsilon}$, $k\in\{1,\dots,N\}$. Collecting all estimates gives
$\limsup_{N\longrightarrow\infty}\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\int_{0}^{T}|g(s)-g_{\varepsilon}(s)|^{2}C(s)^{2}\,\mathrm{d}s.$
The right-hand side tends to zero as $\varepsilon\downarrow 0$ due to dominated convergence, which completes the proof. ∎
Finally, we prove an auxiliary result for kernels that preserve positivity.
Lemma B.3.
Suppose $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is a non-negativity preserving kernel with $K(0_{+})>0$ and let $f\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ be non-decreasing. Let $N\in\mathbb{N}$, $0<t_{1}<\dots<t_{N}$ and $x_{1},\dots,x_{N}\in\mathbb{R}$ be such that, for any $k\in\{1,\dots,N\},$
(39) $f(t_{k})+\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})\geq 0.$
Then, it follows
$f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}x_{k}K(t-t_{k})\geq 0,\quad t>0.$
Proof.
We define recursively $\widetilde{x}_{1}\coloneqq\frac{-f(t_{1})}{K(0_{+})}$ and, for $k\in\{1,\dots,N\},$
$\widetilde{x}_{k}\coloneqq\frac{-1}{K(0_{+})}\left(f(t_{k})+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}K(t_{k}-t_{\ell})\right).$
By construction we have for any $k\in\{1,\dots,N\}$
(40) $f(t_{k})+\sum_{\ell=1}^{k}\widetilde{x}_{\ell}K(t_{k}-t_{\ell})=0.$
Moreover, define $\delta_{k}\coloneqq x_{k}-\widetilde{x}_{k},\,k\in\{1,\dots,N\}.$ Then, we have by (40) and (39)
$\sum_{\ell=1}^{k}\delta_{\ell}K(t_{k}-t_{\ell})=\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})+f(t_{k})\geq 0,\quad k\in\{1,\dots,N\},$
and since $K$ preserves non-negativity, we obtain $\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\delta_{k}K(t-t_{k})\geq 0$ for $t>0$. We have
$f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}x_{k}K(t-t_{k})=f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\widetilde{x}_{k}K(t-t_{k})+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\delta_{k}K(t-t_{k})$
and hence, it suffices to show that
(41) $f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\widetilde{x}_{k}K(t-t_{k})\geq 0,\quad t>0.$
By induction over $k$ we show that (41) holds on $[t_{k},t_{k+1})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k+1}\leq 0$. For $t\in(0,t_{1})$, (41) is evident and by assumption $\widetilde{x}_{1}=\frac{-f(t_{1})}{K(0_{+})}\leq 0$. Suppose (41) holds on $[t_{k-1},t_{k})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k}\leq 0.$ Then, we find for $t\in[t_{k},t_{k+1})$
$f(t)+\sum_{\ell=1}^{k}\widetilde{x}_{\ell}K(t-t_{\ell})$ $\geq f(t)+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}K(t-t_{\ell})-\left(f(t_{k})+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}K(t_{k}-t_{\ell})\right)$
$=f(t)-f(t_{k})+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}\left(K(t-t_{\ell})-K(t_{k}-t_{\ell})\right)$
$\geq f(t)-f(t_{k})$
$\geq 0,$
where we used that $K(t-t_{\ell})-K(t_{k}-t_{\ell})\leq 0$ since $K\geq 0$ is non-increasing. Replacing $t$ by $t_{k+1}$ in the above calculation yields $\widetilde{x}_{k+1}\leq 0.$