### v2

# Comparison Principles for Stochastic Volterra equations

Abstract.
In this work, we establish a comparison principle for stochastic Volterra equations with respect to the initial condition and the drift $b$ applicable to a wide class of Volterra kernels and input curves $g$ that may be singular at zero. The latter appear, e.g., in the study of Markovian lifts for such Volterra equations. For completely monotone kernels, our result holds without any further restrictions, while for not completely monotone kernels, it is shown that such a principle fails unless the drift is additionally monotone. As a side-product of our results, we also complement the literature on the weak existence of continuous nonnegative solutions, which covers the rough Cox-Ingersoll-Ross process with singular initial conditions.
O.C. is funded by the Irish Research Council grant GOIPG/2023/3129 and was previously supported by the ERASMUS Mobility project during which part of this research was carried out.
Keywords: stochastic Volterra process; comparison principle; rough Cox-Ingersoll-Ross process; splitting method
1. Introduction
Let $B$ be a standard one-dimensional Brownian motion, $0\neq K\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ a Volterra convolution kernel, $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ continuous with linear growth in the spatial variable, and $g\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. In this work, we study stochastic Volterra equations (SVEs) of the form
(1) $X_{t}=g(t)+\int_{0}^{t}K(t-s)b(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,X_{s})\,\mathrm{d}B_{s}.$
In many applications, one assumes that the Volterra kernel is given by the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,2)$. Such equations have gained increased attention in mathematical Finance for the modelling of rough volatility, see [8, 14, 15, 16, 19].
The weak existence of a solution was established in [4] whenever $b,\sigma$ are continuous with linear growth, $g$ is constant and $K$ satisfies an additional regularity condition given below. An extension to processes with jumps and more general $g$ is discussed in [1]. When $b$ and $\sigma$ are Lipschitz continuous, the existence of a unique strong solution was shown in [2, Theorem A.1], see also [9, 28]. For non-Lipschitz coefficients, the pathwise uniqueness of solutions is more subtle and yet not fully understood. For the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$, the method proposed in [25] gives the existence and uniqueness of strong solutions whenever $\sigma$ is $\eta$-Hölder continuous and $\alpha\eta>\frac{1}{2}$. An extension of this method towards general drifts and non-convolution equations was recently discussed in [27]. Most recently, in [22], the author proves the weak existence and uniqueness in law for (1) provided that $b,\sigma$ are uniformly continuous, $\sigma$ is non-degenerate, and some technical condition on the modulus of continuity holds. In contrast to the existing literature, see also [26, 27], in this work we focus on continuous solutions where $g$ is not regular but belongs to $g\in L_{\mathrm{loc}}^{p}(\mathbb{R}_{+})\cap C((0,\infty))$ with an additional growth condition at the origin. Such functions $g$ that are singular in $t=0$ appear, e.g., from Markovian lifts where the initial condition does not belong to the domain of the corresponding projection operator, see e.g. [22, Theorem 3.3]. Thus, our Theorem 2.7 naturally complements the results obtained therein.
In the main part of this work, we study comparison principles for solutions of (1). More precisely, given initial data $(g_{i},K,b_{i},\sigma)$ with $i=1,2$, we provide sufficient conditions such that the corresponding solutions $X^{1},X^{2}$ satisfy
$\mathbb{P}[X_{t}^{1}\leq X_{t}^{2},\ t\geq 0]=1.$
While such comparison principles are well-known for classical ordinary differential equations (see [33]) and for Markovian stochastic equations (see [23] for the continuous case and [18] for the case of continuous-state branching processes), much less is known for their Volterra counterparts. For instance, comparison principles for deterministic Volterra equations have been studied in [21, Chapter 12.1], while the first results on stochastic Volterra equations were obtained in [17, 31] for general drifts and restrictive assumptions on the diffusion which exclude our convolution setting (1).
Based on the splitting method, Alfonsi studied in [7] a numerical approximation of stochastic Volterra equations applicable to $g(t)=x$ constant, globally Lipschitz continuous coefficients $b,\sigma$, and $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. The latter work includes, to our knowledge, the first general comparison principle for equations of the form (1). This splitting method has also been recently applied in [6] to prove the existence and uniqueness of nonnegative solutions of stochastic Volterra equations with jumps for non-Lipschitz continuous coefficients and regular Volterra kernels $K\in C^{2}(\mathbb{R}_{+})$. The reader interested in other numerical schemes for stochastic Volterra processes may consult [3, 5] and the references therein. In this work, we first extend Alfonsi’s numerical approximation [7] to non-constant $g$ with a particular focus on possibly singular functions, and time-dependent coefficients. Based on this approximation, we obtain a comparison principle that covers such cases, including those that are not necessarily Lipschitz continuous. Afterwards, we focus on kernels $K$ that are not completely monotone, hence going beyond the framework discussed in [7].
The first key obstacle when dealing with stochastic Volterra equations stems from the absence of the semimartingale property. Indeed, while the classical methods described in [23, Chapter VI] are deeply based on the Itô formula, in our setting an analogous approach is only applicable when $K$ is sufficiently regular in the sense that $K\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$. The latter excludes, e.g., the Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,3/2)$. To address this issue, we extend the splitting method proposed in [7] to time-dependent coefficients and approximate the solution via solutions of classical stochastic differential equations. Observing that the approximation satisfies the comparison principle, we may then pass to the limit. This approach covers $K$ being either non-negativity preserving with $K(0)>0$ or completely monotone. In particular, it covers the fractional kernel with $\alpha\in(1/2,1)$. On the other side, if $K$ is neither non-negativity preserving nor completely monotone, a comparison principle is, in general, not valid as demonstrated in Example 3.1. The latter is, e.g., the case for the fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>1$. Such a failure of the comparison principle is not surprising since solutions of (1) are typically not Markov processes and hence it is not clear if the past trajectory entering the equation via $b,\sigma$ violates the ordering of solutions. To exclude such a possibility, we suppose that the drift is additionally monotone. Then, we first consider the case of regular kernels $K\in W_{\mathrm{loc}}^{1,2}(\mathbb{R}_{+})$ for which a semimartingale representation is available and a modification of the classical comparison principle for SDEs can be used to deduce the comparison principle. Afterwards, we extend these results to general $K\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})\cap C^{1}((0,\infty))$ with $K^{\prime}\geq 0$ or $K^{\prime}\leq 0$ by approximation.
This work is organised as follows. In Section 2 we treat the case of completely monotone kernels, where we first prove a comparison principle for Lipschitz continuous $b,\sigma$. Then we prove the existence of continuous weak solutions for continuous $b,\sigma$ with at most linear growth, and finally derive the corresponding comparison principle in this case. As an application, we conclude Section 2 with a result on the weak existence of continuous nonnegative solutions of (1). Afterwards, in section 3, we study similar results for the case of not completely monotone kernels. Some technical results on the existence, uniqueness, and sample path regularity of solutions to stochastic Volterra equations are collected in Appendix A, while the convergence of the splitting method is studied in Appendix B.
2. Non-negativity preserving case
In this section, we study the case where the Volterra kernel preserves nonnegativity. As a first step, let us introduce the minimal conditions on the input data $(g,K,b,\sigma)$ that will be used for the construction of continuous weak solutions.
Assumption 2.1.
There exist $C>0$, $\xi\in[0,1]$ such that
$|b(t,x)|+|\sigma(t,x)|\leq C(1+|x|)^{\xi}$
for a.a. $t\geq 0$ and $x\in\mathbb{R}$. Moreover, one of the following two cases holds:
  1. (i)
    $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$ or $\xi=0$ with $g\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. Moreover, $K\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and there exist $\gamma\in(0,1/2]$ and for each $T>0$ a constant $c(T)>0$ such that
    $\|K\|_{L^{2}([0,h])}+\|K(\cdot+h)-K(\cdot)\|_{L^{2}([0,T])}\leq c(T)h^{\gamma},\quad h\in[0,1].$
  2. (ii)
    $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$ for some $q\in(2,\infty)$, $\xi\in(0,1]$, there exists $\eta>0$ such that $K\in L^{2+\eta}_{\mathrm{loc}}(\mathbb{R}_{+})$ and there exist $\gamma\in(0,1/2]$ and for each $T>0$ a constant $c(T,\eta)$ such that for $h\in[0,1]$
    $\|K\|_{L^{2+\eta}([0,h])}+\|K(\cdot+h)-K(\cdot)\|_{L^{2+\eta}([0,T])}\leq c(T,\eta)h^{\gamma}.$
    Moreover, the value $q$ satisfies
    (2) $q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}.$
The above assumption allows us to assume without loss of generality that for any solution $X$ of (1), XgX-g is continuous, c.f. Proposition A.2. Note that such a condition is satisfied by the fractional Riemann-Liouville kernel as demonstrated in the next example.
Example 2.2.
Let $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$. Then $2+\eta<\frac{1}{1-\alpha}$, $\gamma=\alpha-1+\frac{1}{2+\eta}$. Hence (2) takes the particular form $q>2\xi\frac{1+1/\eta}{\alpha-1/2}$. By letting $\eta\nearrow\frac{2\alpha-1}{1-\alpha}$, we see that this condition is satisfied whenever
$q>\frac{\xi\alpha}{\left(\alpha-\frac{1}{2}\right)^{2}}.$
Next, to prove the desired comparison principles, let us introduce some concepts that allow us to compare the solutions. In this section, we focus on Volterra kernels that preserve nonnegativity in the sense of the following definition.
Definition 2.3.
A Volterra kernel $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ preserves non-negativity if it satisfies one of the following conditions:
  1. (a)
    $K$ is completely monotone with $K(0_{+})\coloneqq\lim_{t\downarrow 0}K(t)=+\infty$.
  2. (b)
    $K$ is nonincreasing, $K(0_{+})\in(0,\infty)$ and for each $N\in\mathbb{N}$, and all $x_{1},\dots,x_{N}\in\mathbb{R}$ and $0\leq t_{1}<\dots<t_{N}$ it holds that
    (3) $\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})\geq 0,\ \ \forall k\in\{1,\dots,N\}\Longrightarrow\sum_{\ell=1}^{N}\mathbbm{1}_{\{t_{\ell}\leq t\}}x_{\ell}K(t-t_{\ell})\geq 0,\quad\forall t\geq 0.$
Here we use the convention $\sum_{\varnothing}\coloneqq 0$. This condition was recently introduced and studied in [7], and allows us to show that the Volterra kernel preserves the order structure subject to the discretisation of the process via the splitting method. Finally, to obtain the comparison principle for two different input data $(g_{1},K,b_{1},\sigma)$ and $(g_{2},K,b_{2},\sigma)$, we use the following definition.
Definition 2.4.
Let $(g_{i},K,b_{i},\sigma)$ with $i=1,2$ satisfy Assumption 2.1, and suppose that $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is continuous and preserves non-negativity. We say that $(g_{1},K,b_{1},\sigma)$ and $(g_{2},K,b_{2},\sigma)$ are comparable, if the following conditions hold:
  1. (i)
    There exist $h_{1},h_{2}\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and $\widetilde{g_{1}},\widetilde{g}_{2}\in C((0,\infty))$ such that
    (4) $g_{i}(t)=\widetilde{g}_{i}(t)+\int_{0}^{t}K(t-s)h_{i}(s)\,\mathrm{d}s,\qquad i=1,2,$
    and $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing on $(0,\infty)$.
  2. (ii)
    $b_{1}(t,x)+h_{1}(t)\leq b_{2}(t,x)+h_{2}(t)$ holds for each $t\geq 0$ and $x\in\mathbb{R}$.
This class of comparable input data is sufficiently rich to cover the most frequently used examples from applications, as shown at the end of this section. The following is our first result for non-negativity preserving kernels.
Theorem 2.5.
Suppose that $(g_{i},K,b_{i},\sigma)$ with $i=1,2$ are comparable in the sense of Definition 2.4. Assume that there exists $\delta>0$ with
(5) $\int_{0}^{T}\overline{g}_{i}(t)^{2+\delta}(1+h_{i}(t)^{2})\,\mathrm{d}t$ $<\infty,\qquad i=1,2,$
where $\overline{g}_{i}(t)=\sup_{s\in[t,T]}|g_{i}(s)|$, and that there exists $C>0$ such that for a.a. $t\geq 0$ and $x,y\in\mathbb{R}$,
(6) $|b_{1}(t,x)-b_{1}(t,y)|+|b_{2}(t,x)-b_{2}(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq C|x-y|.$
Let $X^{1},X^{2}$ be the unique strong solutions of (1) with input data $(g_{i},K,b_{i},\sigma)$, $i=1,2$. Then $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\,t>0]=1$.
Proof.
First note that equation (1) has unique strong solutions $X^{1},X^{2}$ with sample paths in $L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ due to Proposition A.1. Moreover, given Assumption 2.1, it follows from Proposition A.2 that $X^{1}-g_{1}$, $X^{2}-g_{2}$ have a modification with continuous sample paths. Hence $X^{1},X^{2}$ are continuous on $(0,\infty)$. Finally, define
$\widetilde{b}_{i}(t,x)=b_{i}(t,x)+h_{i}(t),\quad t\in[0,T].$
Hence $\widetilde{b}_{1}\leq\widetilde{b}_{2}$ and (6) still holds with the same $C$, while the linear growth constant becomes $|\widetilde{b}_{1}(t,x)|+|\widetilde{b}_{2}(t,x)|\leq\widetilde{C}(t)(1+|x|)$ with $\widetilde{C}(t)=C+|h_{1}(t)|+|h_{2}(t)|$, so that $\widetilde{C}\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$. The proof is divided into three steps.
Step 1: We introduce a similar splitting method to Alfonsi [7], but now applied for non-constant possibly singular at zero $g$ and time-dependent coefficients. Suppose first that $K$ is continuous with $K(0_{+})>0$. Let $T>0$, $N\in\mathbb{N}$ be fixed and $t_{k}=kT/N,$ $k\in\{0,1,\dots,N\},$ be an equidistant grid of $[0,T]$ with step size $T/N.$ Recursively, we define the process $(\widehat{X}^{i}_{t})_{t\in(0,T]}$ by setting
$\widehat{X}^{i}_{t}$ $=\widetilde{g}_{i}(t)+\sum_{k=1}^{N}\mathbbm{1}_{[t_{k},T]}(t)\left(K(t-t_{k})\int_{t_{k-1}}^{t_{k}}[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}]\right)$
where $(\xi^{i,k}_{t})_{t\in[t_{k-1},t_{k})}$, $k\in\{1,\dots,N\},$ denotes the strong solution of the auxiliary SDE
$\xi^{i,k}_{t}=\widehat{X}_{t_{k}-}^{i}+\int_{t_{k-1}}^{t}K(0_{+})[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}].$
Under the given conditions, it follows from Proposition B.2 that
(7) $\mathbb{E}[|\widehat{X}^{i}_{t}-X^{i}_{t}|^{2}]\longrightarrow 0,\qquad N\longrightarrow\infty.$
Thus it suffices to prove that $\widehat{X}_{t}^{1}\leq\widehat{X}_{t}^{2}$ a.s. for $t\in(0,T]$.
Step 2: We consider the case where $K$ is continuous and non-negativity preserving with $K(0_{+})\in(0,\infty).$ Let $T>0$ be fixed. For $N\in\mathbb{N}$, we define the corresponding approximations $\widehat{X}^{1},\widehat{X}^{2}$ and denote by $\xi^{1},\xi^{2}$ the auxiliary processes that arise in their construction. Note that we have by continuity of $K$ and $g$
$\xi^{i,k}_{t_{k}-}$ $=\widetilde{g}(t_{k})+\sum_{\ell=1}^{k-1}K(t_{k}-t_{\ell})\int_{t_{\ell-1}}^{t_{\ell}}[\widetilde{b}_{i}(s,\xi^{i,\ell}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,\ell}_{s})\,\mathrm{d}B_{s}]$
$\quad+K(t_{k}-t_{k})\int_{t_{k-1}}^{t_{k}}[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}]=\widehat{X}^{i}_{t_{k}}$
and similarly
$\widehat{X}^{i}_{t_{k}}-\widehat{X}^{i}_{t_{k}-}$
$\quad=\widetilde{g}_{i}(t_{k})+\sum_{\ell=1}^{k}K(t_{k}-t_{\ell})\int_{t_{\ell-1}}^{t_{\ell}}[\widetilde{b}_{i}(s,\xi^{i,\ell}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,\ell}_{s})\,\mathrm{d}B_{s}]$
$\qquad-\widetilde{g}_{i}(t_{k})-\sum_{\ell=1}^{k-1}K(t_{k}-t_{\ell})\int_{t_{\ell-1}}^{t_{\ell}}[\widetilde{b}_{i}(s,\xi^{i,\ell}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,\ell}_{s})\,\mathrm{d}B_{s}]$
$=K(0_{+})\int_{t_{k-1}}^{t_{k}}[\widetilde{b}_{i}(s,\xi^{i,k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{i,k}_{s})\,\mathrm{d}B_{s}].$
Hence, we can represent them in compact form as
$\widehat{X}^{i}_{t}=\widetilde{g}_{i}(t)+\sum_{\ell=1}^{N}\mathbbm{1}_{[t_{\ell},T]}(t)\frac{\widehat{X}^{i}_{t_{\ell}}-\widehat{X}^{i}_{t_{\ell}-}}{K(0_{+})}K(t-t_{\ell}),\qquad t\in(0,T].$
We show by induction on $k$ that $\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t}$ on $(0,t_{k}]$. For $k=1,$ we have per assumption $\widehat{X}^{1}_{t}=\widetilde{g}_{1}(t)\leq\widetilde{g}_{2}(t)=\widehat{X}^{2}_{t}$ for $t\in(0,t_{1})$. Using the Comparison Principle for SDEs, see [23, Chapter VI, Theorem 1.1], we obtain $\widehat{X}^{1}_{t_{1}}=\xi_{t_{1}-}^{1,1}\leq\xi_{t_{1}-}^{2,1}=\widehat{X}_{t_{1}}^{2}$. Suppose that the result is true for some $k\geq 1$. Then, we have for all $j\in\{1,\dots,k\}$, by construction
$0\leq\widehat{X}_{t_{j}}^{2}-\widehat{X}_{t_{j}}^{1}=(\widetilde{g}_{2}-\widetilde{g}_{1})(t_{j})+\sum_{\ell=1}^{j}\frac{(\widehat{X}_{t_{\ell}}^{2}-\widehat{X}_{t_{\ell}-}^{2})-(\widehat{X}_{t_{\ell}}^{1}-\widehat{X}_{t_{\ell}-}^{1})}{K(0_{+})}K(t_{j}-t_{\ell}).$
Since $K$ preserves non-negativity and $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing, we obtain by Lemma B.3 that $\widehat{X}_{t}^{2}-\widehat{X}_{t}^{1}\geq 0$ holds for $t\in[t_{k},t_{k+1})$. Using again the comparison result for SDEs yields
$\widehat{X}_{t_{k+1}}^{1}=\xi^{1,k+1}_{t_{k+1}-}\leq\xi^{2,k+1}_{t_{k+1}-}=\widehat{X}^{2}_{t_{k+1}}.$
By induction, we get for $k=N$ that $\mathbb{P}[\widehat{X}^{1}_{T}\leq\widehat{X}^{2}_{T}]=1$. By (7) we conclude $\mathbb{P}[X^{1}_{T}\leq X^{2}_{T}]=1$. Since $T>0$ was arbitrary and $X$ has continuous sample paths on $(0,\infty)$, we obtain the desired comparison principle.
Step 3: Let us now prove the assertion when $K(0_{+})=+\infty$ and $K$ is completely monotone. Using the Bernstein theorem, see [30, Theorem 4.8], we find a Borel measure $\mu$ on $\mathbb{R}_{+}$ such that $K(t)=\int_{\mathbb{R}_{+}}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho)$ for $t\in(0,\infty)$. Define, for $H>0$, the approximation $K^{H}\colon\mathbb{R}_{+}\longrightarrow\mathbb{R}_{+}$ by $K^{H}(t)\coloneqq\int_{[0,H]}\mathrm{e}^{-\rho t}\,\mu(\mathrm{d}\rho)$. Then $0\leq K^{H}(t)\leq K^{H}(0)$ and $|(K^{H})^{\prime}(t)|\leq HK^{H}(0)$. Hence, it satisfies Assumption 2.1 with $\gamma=1/(2+\eta)$ and therefore
$X^{i,H}_{t}=g_{i}(t)+\int_{0}^{t}K^{H}(t-s)b_{i}(s,X^{i,H}_{s})\,\mathrm{d}s+\int_{0}^{t}K^{H}(t-s)\sigma(s,X^{i,H}_{s})\,\mathrm{d}B_{s}$
admits a unique, strong continuous solution for $i=1,2$. Since completely monotone kernels are non-negativity preserving and satisfy (3), see [7, Theorem 2.3], we may apply steps 1 and 2 to obtain $\mathbb{P}[X^{1,H}_{T}\leq X^{2,H}_{T}]=1$ for each $T>0$. Next, we prove that
$\lim_{H\longrightarrow\infty}\mathbb{E}[|X^{i}_{T}-X^{i,H}_{T}|^{2}]=0$
which completes the proof. Indeed, for fixed $t>0$, we have by Jensen’s inequality and Itô’s isometry
$\mathbb{E}[|X^{i}_{t}-X_{t}^{i,H}|^{2}]$ $\lesssim\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[b(s,X^{i}_{s})^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t}K^{H}(t-s)^{2}\mathbb{E}[|b_{i}(s,X_{s})-b_{i}(s,X^{i,H}_{s})|^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[\sigma(s,X^{i}_{s})^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t}K^{H}(t-s)^{2}\mathbb{E}[|\sigma(s,X_{s})-\sigma(s,X^{i,H}_{s})|^{2}]\,\mathrm{d}s$
$\lesssim\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[(1+|X_{s}|)^{2\xi}]\,\mathrm{d}s$
$\qquad+\int_{0}^{t}K^{H}(t-s)^{2}\mathbb{E}[|X_{s}-X^{i,H}_{s}|^{2}]\,\mathrm{d}s$
where we have used the linear growth and Lipschitz condition. For the first term, we find by using Hölder’s inequality
$\int_{0}^{t}|K(t-s)-K^{H}(t-s)|^{2}\mathbb{E}[(1+|X_{s}|)^{2\xi}]\,\mathrm{d}s$
$\quad\leq\left(\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s\right)^{\frac{2}{2+\eta}}\mathbb{E}\left[\left(\int_{0}^{T}(1+|X_{s}|)^{2\xi\left(1+\frac{2}{\eta}\right)}\,\mathrm{d}s\right)^{\frac{\eta}{2+\eta}}\right]$
$\quad\leq\left(\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s\right)^{\frac{2}{2+\eta}}\int_{0}^{T}\mathbb{E}\left[(1+|X_{s}|)^{2\xi\left(1+\frac{2}{\eta}\right)}\right]\,\mathrm{d}s$
where the second term is finite due to Assumption 2.1
$q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}\geq 2\xi\frac{1+\frac{1}{\eta}}{\frac{1}{2}+\frac{1}{2}\frac{\eta}{2+\eta}}=2\xi\left(1+\frac{2}{\eta}\right)$
and Proposition A.1. Applying the Volterra-type Gronwall inequality yields
$\mathbb{E}[|X_{t}-X^{i,H}_{t}|^{2}]\\ \leq\left(\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s\right)^{\frac{2}{2+\eta}}\left(\int_{0}^{T}\mathbb{E}\left[(1+|X_{s}|)^{2\xi\left(1+\frac{2}{\eta}\right)}\right]\mathrm{d}s\right)\left(1+\int_{0}^{t}R(s)\,\mathrm{d}s\right),$
where $R$ denotes the resolvent of the second kind of $c(K^{H})^{2}$, and $c$ is some large enough constant. Finally, the first term can be bounded by
$\int_{0}^{T}|K(s)-K^{H}(s)|^{2+\eta}\,\mathrm{d}s$ $\leq\int_{0}^{T}\left(\int_{[H,\infty)}\mathrm{e}^{-\rho s}\,\mu(\mathrm{d}\rho)\right)^{2+\eta}\,\mathrm{d}s$
$\leq\int_{0}^{T}\mathrm{e}^{-sH(2+\eta)/2}K(s/2)^{2+\eta}\,\mathrm{d}s.$
By dominated convergence, the latter tends to zero as $H\longrightarrow\infty$, which proves the assertion. ∎
Note that the case where $g_{i}(t)=x_{i}$ for $i=1,2$ was already covered in [7]. Since the fractional Riemann-Liouville kernel satisfies Assumption 2.1 and is completely monotone, Theorem 2.5 can be applied to this case.
Example 2.6.
Let $X^{1},X^{2}$ be fractional Ornstein-Uhlenbeck processes given as the unique strong solutions of
$X^{i}_{t}=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}(b_{i}+\beta X^{i}_{s})\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\,\mathrm{d}B_{s},$
where $b_{1}\leq b_{2}$, $\beta,\sigma\in\mathbb{R}$, $\alpha\in(1/2,1)$, $\gamma>0$, and $x_{1}\leq x_{2}$. If $\beta\geq 0$, then the comparison principle holds. If $\beta<0$, then the comparison principle holds if and only if $\gamma\geq\alpha$.
Proof.
It follows from [4] that $X^{i}$ is explicitly given by
$X_{t}^{i}=x_{i}t^{\gamma-1}E_{\alpha,\gamma}(\beta t^{\alpha})+\int_{0}^{t}s^{\alpha-1}E_{\alpha,\alpha}(\beta s^{\alpha})b_{i}\,\mathrm{d}s+\int_{0}^{t}(t-s)^{\alpha-1}E_{\alpha,\alpha}(\beta(t-s)^{\alpha})\,\mathrm{d}B_{s}$
where $E_{\alpha,\gamma}(z)=\sum_{n=0}^{\infty}\frac{z^{n}}{\Gamma(\alpha n+\gamma)}$ denotes the two-parameter Mittag-Leffler function. Hence
$X_{t}^{2}-X_{t}^{1}=(x_{2}-x_{1})t^{\gamma-1}E_{\alpha,\gamma}(\beta t^{\alpha})+\int_{0}^{t}s^{\alpha-1}E_{\alpha,\alpha}(\beta s^{\alpha})(b_{2}-b_{1})\,\mathrm{d}s.$
Since $E_{\alpha,\alpha}(\beta t^{\alpha})\geq 0$ for any $\beta,t$, the comparison principle is satisfied if and only if $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ for all $t\geq 0$. If $\beta\geq 0$, then $E_{\alpha,\gamma}(\beta t^{\alpha})\geq 0$ and nothing needs to be shown. Suppose that $\beta<0$. If $\gamma\geq\alpha$, then $t\longmapsto E_{\alpha,\gamma}(-t^{\alpha})$ is completely monotone, see [20, Section 4.9.2], and hence the comparison principle holds. Conversely, suppose that $\gamma<\alpha$. Then $(0,\infty)\ni t\longmapsto t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ is integrable and its Laplace transform is given by $\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\mathrm{e}^{-st}\,\mathrm{d}s=\frac{s^{\alpha-\gamma}}{s^{\alpha}+|\beta|}$ for $s\geq 0$. Evaluating this in $s=0$ gives
$\int_{0}^{\infty}t^{\gamma-1}E_{\alpha,\gamma}(-|\beta|t^{\alpha})\,\mathrm{d}s=0$
and hence $(0,\infty)\ni t\longmapsto E_{\alpha,\gamma}(-|\beta|t^{\alpha})$ needs to attain negative values. Consequently, the comparison principle fails to hold. ∎
While the above proof is direct, one may also seek to apply Theorem 2.5 instead. Indeed, consider the fractional Riemann-Liouville kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha\in(1/2,1)$ and define $g_{i}(t)=x_{i}\frac{t^{\gamma-1}}{\Gamma(\gamma)}$ where we additionally assume that $\gamma\geq\alpha$. Then $\widetilde{g}_{i}(t)=0$ and $h_{i}(t)=x_{i}\frac{t^{(\gamma-\alpha)-1}}{\Gamma(\gamma-\alpha)}$. In particular, condition (ii) from Definition 2.4 reduces to
$(b_{2}-b_{1})+(x_{2}-x_{1})\left(\beta+\frac{t^{(\gamma-\alpha)-1}}{\Gamma(\gamma-\alpha)}\right)\geq 0.$
Hence, Theorem 2.5 is applicable for $\beta\geq 0$, while for $\beta<0$ it is generally not applicable.
Next, let us examine cases where the coefficients $b_{1},b_{2},\sigma$ are not Lipschitz continuous. Since, in this generality, the weak existence of solutions is not guaranteed from the literature, below we complement the latter by proving the weak existence of continuous solutions when the function $g$ may be singular in $t=0$. Our proof below employs some tightness arguments. However, in contrast to existing arguments such as [4, 26], here we need to introduce an auxiliary weighted space of continuous functions to control the order of singularity of the function $g$.
Theorem 2.7.
Let $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ be continuous, nonincreasing, and not identically zero. If $(g,K,b,\sigma)$ satisfies Assumption 2.1 and $g$ is continuous on $(0,\infty)$ such that there exists $\delta\in(0,1/2)$ such that for every $T>0$
(8) $\sup_{t\in(0,T]}t^{\delta}|g(t)|<\infty,$
then the stochastic Volterra equation (1) admits a continuous weak solution.
Proof.
Following [26, Lemma 3.6], we define for $f\in\{b,\sigma\}$ a sequence of functions
(9) $f_{n}(t,x)=\psi_{n}(x)\int_{\mathbb{R}}f(t,x-y)\varphi_{n}(y)\,\mathrm{d}y,\quad n\in\mathbb{N},$
where $\psi_{n}\colon\mathbb{R}\longrightarrow[0,1]$ with $\operatorname{supp}\psi_{n}\subset[-(n+1),n+1]$ and $\psi_{n}(x)=1$ for $x\in[-n,n]$, where $\varphi(y)=\frac{1}{c_{n}}(1-y^{2})^{n}\mathbbm{1}_{[-1,1]}(y)$ with $c_{n}=\int_{[-1,1]}(1-y^{2})^{n}\,\mathrm{d}y$. Then $(f_{n})_{n\in\mathbb{N}}$ is a sequence of Lipschitz continuous functions in the second variable and $\lim_{n\to\infty}f_{n}(t,\cdot)=f(t,\cdot)$ uniformly on compacts for all $t\geq 0$. Moreover, the following growth condition is satisfied uniformly in $n$,
(10) $|b_{n}(t,x)|+|\sigma_{n}(t,x)|\leq C(1+|x|)^{\xi},\quad\text{a.a. }t\geq 0,x\in\mathbb{R}.$
Let $(X^{n})_{n\in\mathbb{N}}$ be the family of processes constructed for $(g,K,b_{n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. Thanks to Proposition A.2 and (10), one readily checks that for $T>0$ and all $s,t\in[0,T]$,
$\sup_{n\in\mathbb{N}}\mathbb{E}[{|(X_{t}^{n}-g(t))-(X_{s}^{n}-g(s))|^{p}}]\leq c|t-s|^{1+\theta}$
holds for some $\theta>0$, $p\geq 2$ and a constant $c$ that only depends on $C$, $T$, $K$, pp. By applying Kolmogorov’s tightness criterion (see e.g. [29, Chapter XIII, Theorem 1.8]), we conclude that $(X^{n}-g)_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+})$ with limit denoted by $\overline{\mathbb{P}}_{g}$. By abuse of notation, we denote this subsequence also by $(X^{n}-g)_{n\in\mathbb{N}}$. Let $C(\mathbb{R}_{+})_{0}\subset C(\mathbb{R}_{+})$ be the space of continuous functions that vanish at $t=0$. Since $\mathcal{L}(X^{n}-g)(C(\mathbb{R}_{+})_{0})=1$ it follows by the Portmanteau theorem that
$1=\limsup_{n\to\infty}\mathcal{L}(X^{n}-g)(C(\mathbb{R})_{0})\leq\overline{\mathbb{P}}_{g}(C(\mathbb{R}_{+})_{0}).$
Next, to prove weak convergence of $(X^{n})_{n\in\mathbb{N}}$, let $C_{\kappa}([0,T])$ with $\kappa>0$ be the Banach space of continuous functions $f\colon(0,T]\longrightarrow\mathbb{R}$ with finite norm $\|f\|_{\kappa,T}\coloneqq\sup_{t\in(0,T]}t^{\kappa}|f(t)|$ such that $\lim_{t\downarrow 0}t^{\kappa}f(t)$ exists. Note that $\iota\colon C_{\kappa}([0,T])\longrightarrow C([0,T])$ defined by $\iota(f)(t)=t^{\kappa}f(t)$ satisfies $\|\iota(f)\|_{[0,T]}=\|f\|_{\kappa,T}$, and hence is an isometric isomorphism, whence $C_{\kappa}([0,T])$ is a polish space. Finally, we let $C_{\kappa}(\mathbb{R}_{+})\coloneqq\bigcap_{k\in\mathbb{N}}C_{\kappa}([0,k])$ be equipped with the metric
$\varrho(f,g)\coloneqq\sum_{k=1}^{\infty}2^{-k}(\|f-g\|_{\kappa,k}\land 1).$
By assumption (8), $g\in C_{\kappa}(\mathbb{R}_{+})$ for $\kappa>\delta$. Hence $\Phi\colon C(\mathbb{R}_{+})_{0}\longrightarrow C_{\kappa}(\mathbb{R}_{+})$ given by $\Phi(f)=f+g$ defines a continuous mapping. Thus, $(\Phi(X^{n}-g))_{n\in\mathbb{N}}=(X^{n})_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+})$ with weak limit denoted by $\overline{\mathbb{P}}$. Note that $\overline{\mathbb{P}}_{g},\overline{\mathbb{P}}$ are related by
(11) $\overline{\mathbb{P}}_{g}\circ\Phi^{-1}=\overline{\mathbb{P}}.$
It remains to show that $\overline{\mathbb{P}}$ determines a weak solution of (1).
Let $L$ be the resolvent of the first kind of KK111i.e. a locally finite (positive) measure $L$ such that $L\ast K=K\ast L=1$. The latter exists since $K\neq 0$ is non-negative and non-increasing, see [21, Theorem 5.5.5].. Let $(M_{t}^{n})_{t\geq 0}$ be the continuous $\mathbb{P}$-martingale for the filtration $\mathbb{F}$ given by
$M_{t}^{n}=\int_{0}^{t}\sigma_{n}(s,X^{n}_{s})\,\mathrm{d}B_{s}=\int_{[0,t]}(X^{n}_{t-s}-g(t-s))\,L(\mathrm{d}s)-\int_{0}^{t}b_{n}(s,X^{n}_{s})\,\mathrm{d}s$
where the second equality follows from [4, Lemma 2.6]. Moreover, since (10) holds uniformly in $n\in\mathbb{N}$, it follows that $(M_{t}^{n})_{n\in\mathbb{N}}$ is uniformly integrable due to
$\sup_{n\in\mathbb{N}}\mathbb{E}\left[\sup_{t\in[0,T]}|M_{t}^{n}|^{2}\right]$ $\leq\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}\left[\sigma_{n}(s,X_{s}^{n})^{2}\right]\mathrm{d}s$
$\lesssim C\left(T+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}\left[|X_{s}^{n}|^{2}\right]\mathrm{d}s\right)<\infty.$
Define a continuous mapping $\mathcal{M}\colon C(\mathbb{R}_{+})_{0}\longrightarrow C(\mathbb{R}_{+})_{0}$ by
$\mathcal{M}_{t}(w)=\int_{[0,t]}w(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b(s,\Phi(w)(s))\,\mathrm{d}s,$
where the continuity is guaranteed by [21, Corollary 3.6.2 (iii)]. Since $X^{n}-g\in C(\mathbb{R}_{+})_{0}$, by continuous mapping theorem combined with (11), we obtain
$M^{n}=\mathcal{M}(X^{n}-g)\Longrightarrow\overline{\mathbb{P}}_{g}\circ\mathcal{M}^{-1}=\overline{\mathbb{P}}\circ(\mathcal{M}^{\Phi})^{-1}$
where $\mathcal{M}^{\Phi}\colon\Phi(C(\mathbb{R}_{+})_{0})\longrightarrow C(\mathbb{R}_{+})$ is given by
$\mathcal{M}_{t}^{\Phi}(w)\coloneqq\mathcal{M}_{t}(\Phi^{-1}(w))=\int_{[0,t]}(w(t-s)-g(t-s))\,L(\mathrm{d}s)-\int_{0}^{t}b(s,w(s))\,\mathrm{d}s.$
Next, we show that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale for the natural filtration generated by coordinate process on $C_{\kappa}(\mathbb{R}_{+})$, and determine its quadratic variation. The continuity follows from $\Phi(C(\mathbb{R}_{+})_{0})=\{f\in C_{\kappa}(\mathbb{R}_{+}):f-g\in C(\mathbb{R}_{+})_{0}\}$, [21, Corollary 3.6.2 (iii)] and $\overline{\mathbb{P}}(\Phi(C(\mathbb{R}_{+})_{0}))=\overline{\mathbb{P}}_{g}(C(\mathbb{R}_{+})_{0})=1.$ Let $s<t$, $m\in\mathbb{N}$, $f\colon\mathbb{R}^{m}\longrightarrow\mathbb{R}$ bounded and continuous, and $0\leq s_{1}\leq\dots\leq s_{m}\leq s$. Let $(x(t))_{t\geq 0}$ denote the coordinate process $x(t)\colon C_{\kappa}(\mathbb{R}_{+};\mathbb{R})\longrightarrow\mathbb{R},\ \ w\longmapsto w(t)$. Applying [11, Theorem 3.5] yields
(12) $\mathbb{E}^{\overline{\mathbb{P}}}\left[f(x(s_{1}),\dots,x(s_{m}))\left(\mathcal{M}_{t}^{\Phi}-\mathcal{M}_{s}^{\Phi}\right)\right]=\lim_{n\to\infty}\mathbb{E}[f(X^{n}_{s_{1}},\dots,X^{n}_{s_{m}})(M_{t}^{n}-M_{s}^{n})]=0.$
A monotone class argument implies that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a $\overline{\mathbb{P}}$-martingale. To determine its quadratic variation, let us write
$\langle M^{n}\rangle=\int_{0}^{\cdot}\left(\sigma_{n}(s,X_{s}^{n})^{2}-\sigma(s,X_{s}^{n})^{2}\right)\,\mathrm{d}s+\int_{0}^{\cdot}\sigma(s,X_{s}^{n})^{2}\,\mathrm{d}s.$
Since $X^{n}\Longrightarrow\overline{\mathbb{P}}$, the continuous mapping theorem implies that
$\int_{0}^{\cdot}\sigma(s,X_{s}^{n})^{2}\,\mathrm{d}s\Longrightarrow\mathcal{L}_{\overline{\mathbb{P}}}\left(\int_{0}^{\cdot}\sigma(s,x(s))\,\mathrm{d}s\right).$
Thus, by Slutsky’s theorem, it suffices to show that the first term converges to zero in probability. Let $T>0$, then
$\ \int_{0}^{T}\mathbb{E}\left[\left|\sigma_{n}(s,X_{s}^{n})^{2}-\sigma(s,X_{s}^{n})^{2}\right|\right]\mathrm{d}s$
$\lesssim\left(\int_{0}^{T}\mathbb{E}\left[\sigma_{n}(s,X_{s}^{n})^{2}+\sigma(s,X_{s}^{n})^{2}\right]\mathrm{d}s\right)^{1/2}\left(\int_{0}^{T}\mathbb{E}\left[\left(\sigma_{n}(s,X_{s}^{n})-\sigma(s,X_{s}^{n})\right)^{2}\right]\mathrm{d}s\right)^{1/2}$
$\lesssim\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}[|X_{s}^{n}|^{2}]\,\mathrm{d}s\right)^{1/2}\left(\int_{0}^{T}\mathbb{E}\left[(\sigma_{n}(s,X_{s}^{n})-\sigma(s,X^{n}_{s}))^{2}\right]\mathrm{d}s\right)^{1/2}$
$\lesssim R^{-\lambda/2}\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}[|X_{s}^{n}|^{2}]\mathrm{d}s\right)^{1/2}\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}\left[|X_{s}^{n}|^{2+\lambda}\right]\mathrm{d}s\right)^{1/2}$
$\qquad+\left(1+\sup_{n\in\mathbb{N}}\int_{0}^{T}\mathbb{E}[|X_{s}^{n}|^{2}]\,\mathrm{d}s\right)^{1/2}\left(\int_{0}^{T}\|\sigma_{n}(s,\cdot)-\sigma(s,\cdot)\|_{[-R,R]}^{2}\,\mathrm{d}s\right)^{1/2}$
where we have used (10) and the inequality
$\mathbb{E}[|\sigma_{n}(s,X^{n}_{s})-\sigma(s,X_{s}^{n})|^{2}]$ $\lesssim\mathbb{E}\left[\mathbbm{1}_{\{|X^{n}_{s}|>R\}}\left(|\sigma_{n}(s,X^{n}_{s})|^{2}+|\sigma(s,X_{s}^{n})|^{2}\right)\right]$
$\qquad\qquad+\|\sigma_{n}(s,\cdot)-\sigma(s,\cdot)\|_{[-R,R]}^{2}$
$\lesssim\frac{C}{R^{\lambda}}\left(1+\mathbb{E}[|X^{n}_{s}|^{2+\lambda}]\right)+\|\sigma_{n}(s,\cdot)-\sigma(s,\cdot)\|_{[-R,R]}^{2},$
where $\lambda=q-2>0$ if $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+}),q\in(2,\infty),$ and $\lambda=1$ if $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$. Thus, letting first $n\longrightarrow\infty$ and then $R\longrightarrow\infty$, proves the assertion. Arguing as in (12) yields that $(\mathcal{M}^{\Phi})^{2}-\int_{0}^{\cdot}\sigma(s,x(s))^{2}\,\mathrm{d}s$ is an $\overline{\mathbb{P}}$-martingale. Consequently, we have shown that $(\mathcal{M}^{\Phi}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale with quadratic variation
$\langle\mathcal{M}^{\Phi}\rangle_{t}=\int_{0}^{t}\sigma(s,x(s))^{2}\,\mathrm{d}s.$
By martingale representation theorem [29, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a Brownian motion $W$ and a predictable process $\alpha$ such that $\alpha_{s}^{2}=\sigma(s,\widehat{X}_{s})^{2}$ and $\mathcal{M}^{\Phi}_{t}=\int_{0}^{t}\alpha_{s}\,\mathrm{d}W_{s}$ where $\widehat{X}$ denotes the processes $x$ considered on the enlargement such that $\mathcal{L}(\widehat{X})=\mathcal{L}(x)$. Consequently, by [4, Lemma 2.6]
$\widehat{X}_{t}$ $=g(t)+\int_{0}^{t}K(t-s)b(s,\widehat{X}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma(s,\widehat{X}_{s})\,\mathrm{d}W_{s}$
which completes the proof. ∎
Remark that $g$ given as in (8) automatically satisfies $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in(0,1/\delta)$, and hence condition (2) is satisfied whenever
$2\xi\delta<\frac{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}{1+\frac{1}{\eta}}.$
As another consequence, we may also prove the existence of nonnegative continuous weak solutions. The latter complements the results obtained in [2].
Corollary 2.8.
Let $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ be continuous and preserve nonnegativity, and let $(g,K,b,\sigma)$ satisfy Assumption 2.1. Moreover, suppose that there exists a sequence of globally Lipschitz continuous functions $b_{n},\sigma_{n}\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ with property (10) such that $b_{n}(t,\cdot)\longrightarrow b(t,\cdot)$, $\sigma_{n}(t,\cdot)\longrightarrow\sigma(t,\cdot)$ locally uniformly for each $t\geq 0$, and it holds that
(13) $b_{n}(t,0)\geq 0\ \text{ and }\ \sigma_{n}(t,0)=0,\qquad t\geq 0,\ n\in\mathbb{N}.$
If $g=\widetilde{g}+K\ast h$ is such that $0\leq\widetilde{g}\in C((0,\infty))\cap L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ is nondecreasing, $0\leq h\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$, (5),(8) hold for some $\delta\in(0,1/2)$, then the stochastic Volterra equation (1) admits a nonnegative continuous weak solution.
Proof.
Let $X^{n}$ be the unique strong solution of (1) with $b,\sigma$ replaced by $b_{n},\sigma_{n}$, i.e.
$X_{t}^{n}=g(t)+\int_{0}^{t}K(t-s)b_{n}(s,X_{s}^{n})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma_{n}(s,X_{s}^{n})\,\mathrm{d}B_{s}.$
By the convergence arguments given in the proof of Theorem 2.7, it suffices to prove that $\mathbb{P}[X^{n}_{t}\geq 0,\,t>0]=1$. To prove the latter, we apply Theorem 2.5 to the pair of processes $X^{n}$ and $Y\equiv 0$. Indeed, note that $Y\equiv 0$ is the unique solution of
$Y_{t}^{n}=\widehat{g}_{n}(t)+\int_{0}^{t}K(t-s)b_{n}(s,Y^{n}_{s})\,\mathrm{d}s+\int_{0}^{t}K(t-s)\sigma_{n}(s,Y^{n}_{s})\,\mathrm{d}B_{s}$
where $\widehat{g}_{n}(t)=-\int_{0}^{t}K(t-s)b_{n}(s,0)\,\mathrm{d}s$. It is easy to check that $(g,K,b_{n},\sigma_{n})$ and $(\widehat{g}_{n},K,b_{n},\sigma_{n})$ are comparable in the sense of Definition 2.4. Hence Theorem 2.5 gives $\mathbb{P}[X^{n}\geq Y=0,\,t>0]=1$, which proves the assertion. ∎
Below, we illustrate this existence result by an application to a general one-dimensional nonnegative process with a power-law diffusion coefficient. The latter contains the rough Cox-Ingersoll-Ross process as a particular case.
Example 2.9.
Let $\lambda,\theta,\sigma>0$, $\gamma_{1},\gamma_{2}\in(0,1]$, and $\alpha\in(1/2,1)$. Take $g$ of the form
$g(t)=\widetilde{g}(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}h(s)\,\mathrm{d}s$
such that $0\leq\widetilde{g}\in C((0,\infty))\cap L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$ is nondecreasing, $0\leq h\in L_{\mathrm{loc}}^{2}(\mathbb{R}_{+})$, and (5), (8) hold for some $\delta\in(0,1/2)$. If Assumption 2.1 is satisfied for
$X_{t}=g(t)+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\lambda\left(\theta-X_{s}\right)^{\gamma_{1}}\,\mathrm{d}s+\sigma\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)},X_{s}^{\gamma_{2}}\,\mathrm{d}B_{s},$
then it admits a continuous nonnegative weak solution.
Proof.
The assertion follows from Corollary 2.8, once we have verified condition (13). Here, we may take $\sigma_{n}(x)=n^{1-\gamma}|x|\mathbbm{1}_{\{|x|\leq\frac{1}{n}\}}+|x|^{\gamma}\mathbbm{1}_{\{|x|>\frac{1}{n}\}}$ for $\sigma(x)=|x|^{\gamma}$ and similarly for the drift coefficient. Then it is easy to see that (13) holds, which proves the assertion. ∎
Next, we turn to the comparison principle beyond Lipschitz continuous coefficients. Since a comparison principle as formulated in Theorem 2.5 would imply pathwise uniqueness of (1), it is in this generality out of reach. However, by approximation, we show that there always exists a couple of solutions $(X^{1},X^{2})$ defined on a joint filtered probability space that satisfies the comparison principle, i.e., the existence of a monotone coupling.
Theorem 2.10.
[…]
(14) $\begin{pmatrix}\widehat{X}^{1}_{t}\\ \widehat{X}^{2}_{t}\end{pmatrix}=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(s,\widehat{X}^{1}_{s})\\ b_{2}(s,\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(s,\widehat{X}^{1}_{s})\\ \sigma(s,\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}\widehat{B}_{s}$
such that $\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$. In particular, its marginals are continuous weak solutions of (1).
Proof.
Let us define $b_{1,n},b_{2,n},\sigma_{n}$ as in (9). Since $b_{1}\leq b_{2}$, it is clear that this approximation inherits the ordering condition $b_{1,n}\leq b_{2,n}$. Let $(X^{n})_{n\in\mathbb{N}},(Y^{n})_{n\in\mathbb{N}}$ be the family of processes constructed for $(g_{i},K,b_{i,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. By Theorem 2.5, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},t>0]=1$.
Following the arguments presented in the proof of Theorem 2.7, we firstly deduce that $((X^{n}-g_{1},Y^{n}-g_{2}))_{n\in\mathbb{N}}$ is a tight sequence of processes and hence has a weakly convergent subsequence on the path space $C(\mathbb{R}_{+};\mathbb{R}^{2})$. Secondly, we extend the spaces $C(\mathbb{R}_{+})_{0}\coloneqq C(\mathbb{R}_{+};\mathbb{R})_{0}$, $C_{\kappa}(\mathbb{R}_{+})\coloneqq C_{\kappa}(\mathbb{R}_{+};\mathbb{R})$ in an obvious way to $C(\mathbb{R}_{+};\mathbb{R}^{2})_{0}$, $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with $\kappa>\delta$ and infer that $((X^{n},Y^{n}))_{n\in\mathbb{N}}$ is weakly convergent on $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ with weak limit denoted by $\overline{\mathbb{P}}.$ To prove the desired comparison result, it is essential to observe that $\overline{\mathbb{P}}$ has inherited the ordering structure. Indeed, let $(x,y)=((x,y)(t))_{t\geq 0}$ be the coordinate process
$(x,y)(t)\colon C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})\longrightarrow\mathbb{R}^{2},\quad w=(w_{1},w_{2})\longmapsto(w_{1}(t),w_{2}(t)).$
Let $A\coloneqq\{w=(w_{1},w_{2})\in C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})\mid w_{2}(t)-w_{1}(t)\geq 0,t>0\}$. Writing
$A_{t}\coloneqq\{(w_{1},w_{2})\in C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})\mid w_{2}(t)-w_{1}(t)\geq 0\}=\pi_{t}^{-1}([0,\infty))$
where $\pi_{t}\colon(C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2}),\varrho)\longrightarrow\mathbb{R}$, $\pi_{t}(w_{1},w_{2})=w_{2}(t)-w_{1}(t)$ is a continuous function, we conclude that $A_{t}$ is closed in $((C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2}),\varrho)$. Hence also $A=\bigcap_{t>0}A_{t}$ is a closed subset of $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$. An application of the Portmanteau theorem yields
$\overline{\mathbb{P}}[x(t)\leq y(t),t\geq 0]=\overline{\mathbb{P}}[A]\geq\limsup_{n\to\infty}\mathbb{P}_{n}[A]=\limsup_{n\to\infty}\mathbb{P}[X_{t}^{n}\leq Y_{t}^{n},t>0]=1.$
Finally, we show that $\overline{\mathbb{P}}$ determines a weak solution to the 2-dimensional stochastic Volterra equation (14). Similarly to the proof of Theorem 2.7, we define the continuous $\mathbb{P}$-martingales
$M^{n}_{t}\coloneqq\int_{0}^{t}\begin{pmatrix}\sigma_{n}(s,X^{n}_{s})\\ \sigma_{n}(s,Y^{n}_{s})\end{pmatrix}\,\mathrm{d}B_{s}=\begin{pmatrix}\int_{[0,t]}X^{n}_{t-s}-g_{1}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{1,n}(s,X^{n}_{s})\,\mathrm{d}s\\ \int_{[0,t]}Y^{n}_{t-s}-g_{2}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{2,n}(s,X^{n}_{s})\,\mathrm{d}s\end{pmatrix}$,
where $L$ denotes the resolvent of the first kind of $K$. Moreover, we observe that $M^{n}\Longrightarrow\overline{\mathbb{P}}\circ\mathcal{M}^{-1}$, where $\mathcal{M}$ is on $\{(w_{1},w_{2})\in C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2}):(w_{1}-g_{1},w_{2}-g_{2})\in C(\mathbb{R}_{+};\mathbb{R}^{2})_{0}\}$ given by
(15) $\mathcal{M}_{t}(w_{1},w_{2})\coloneqq\begin{pmatrix}\int_{[0,t)}w_{1}(t-s)-g_{1}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{1}(s,w_{1}(s))\,\mathrm{d}s\\ \int_{[0,t)}w_{2}(t-s)-g_{2}(t-s)\,L(\mathrm{d}s)-\int_{0}^{t}b_{2}(s,w_{2}(s))\,\mathrm{d}s\\ \end{pmatrix}$
Furthermore, we can show that $(\mathcal{M}_{t})_{t\geq 0}$ is a continuous $\overline{\mathbb{P}}$-martingale for the natural filtration generated by the coordinate process on $C_{\kappa}(\mathbb{R}_{+};\mathbb{R}^{2})$ and its quadratic variation is given by
$\langle\mathcal{M}\rangle_{t}=\int_{0}^{t}\begin{pmatrix}\sigma(s,x(s))^{2}&\sigma(s,x(s))\sigma(s,y(s))\\ \sigma(s,x(s))\sigma(s,y(s))&\sigma(s,y(s))^{2}\end{pmatrix}\mathrm{d}s.$
By martingale representation theorem [29, Chapter V, Theorem 3.9], there exists, possibly on an enlargement of the probability space, a two-dimensional Brownian motion $(W^{1},W^{2})$ and a $\mathbb{R}^{2\times 2}$-valued predictable process $\alpha$ such that
$\mathcal{M}(t)=\int_{0}^{t}\alpha_{s}\,\mathrm{d}\!\begin{pmatrix}W^{1}_{s}\\ W^{2}_{s}\end{pmatrix}\qquad\text{ with }\qquad\alpha_{s}\alpha_{s}^{\top}=\begin{pmatrix}\sigma(s,\widehat{X}_{s})^{2}&\sigma(s,\widehat{X}_{s})\sigma(s,\widehat{Y}_{s})\\ \sigma(s,\widehat{X}_{s})\sigma(s,\widehat{Y}_{s})&\sigma(s,\widehat{Y}_{s})^{2}\end{pmatrix}$
where $\widehat{X},\widehat{Y}$ denote the processes $x,y$ considered on the enlargement such that $\operatorname{Law}(\widehat{X})=\operatorname{Law}(x)$, $\operatorname{Law}(\widehat{Y})=\operatorname{Law}(y)$. Note that any other choice of $\widetilde{\alpha}$ that satisfies $\widetilde{\alpha}\widetilde{\alpha}^{\top}=\alpha\alpha^{\top}$ gives the same law of $\mathcal{M}$ under $\overline{\mathbb{P}}$. Let us define
$\widehat{\alpha}_{t}=\frac{1}{\sqrt{2}}\begin{pmatrix}\sigma(s,\widehat{X}_{t})&\sigma(s,\widehat{X}_{t})\\ \sigma(s,\widehat{Y}_{t})&\sigma(s,\widehat{Y}_{t})\end{pmatrix}\ \text{ and }\ \widehat{\mathcal{M}}(t)=\int_{0}^{t}\widehat{\alpha}_{s}\,\mathrm{d}\!\begin{pmatrix}W^{1}_{s}\\ W^{2}_{s}\end{pmatrix}$.
Then $\langle\widehat{\mathcal{M}}\rangle_{t}=\int_{0}^{t}\widehat{\alpha}_{s}\widehat{\alpha}_{s}^{\top}\,\mathrm{d}s=\int_{0}^{t}\alpha_{s}\alpha_{s}^{\top}\,\mathrm{d}s=\langle\mathcal{M}\rangle_{t}$. Hence $\widehat{\mathcal{M}}$ and $\mathcal{M}$ have the same law. Consequently, by (15) and [4, Lemma 2.6],
$\begin{pmatrix}\widehat{X}_{t}\\ \widehat{Y}_{t}\end{pmatrix}$ $=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(s,\widehat{X}_{s})\\ b_{2}(s,\widehat{Y}_{s})\end{pmatrix}\mathrm{d}s+\frac{1}{\sqrt{2}}\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(s,\widehat{X}_{s})\,\mathrm{d}(W^{1}_{s}+W^{2}_{s})\\ \sigma(s,\widehat{Y}_{s})\,\mathrm{d}(W^{1}_{s}+W^{2}_{s})\end{pmatrix}$
$=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(s,\widehat{X}_{s})\\ b_{2}(s,\widehat{Y}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(s,\widehat{X}_{s})\\ \sigma(s,\widehat{Y}_{s})\end{pmatrix}\mathrm{d}\widehat{B}_{s},$
where $\widehat{B}=\frac{1}{\sqrt{2}}(W^{1}+W^{2})$ is a one-dimensional Brownian motion. This completes the proof.
Also here, let us note that the model given in Example 2.9 satisfies the conditions of Theorem 2.10, and hence fulfils the comparison principle.
3. Regular Case
In this section, we study the case where $K$ does not preserve non-negativity. The latter includes, e.g., the fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>1$. For such kernels $K$ and without further assumptions on the drift coefficients, the comparison principle does not hold as demonstrated in the next example.
Example 3.1.
Let $X^{i}$ with $i=1,2$ be the unique solution of
$X_{t}^{i}=x_{i}\frac{t^{\beta-1}}{\Gamma(\beta)}-\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}X_{s}^{i}\,\mathrm{d}s+\int_{0}^{t}\frac{(t-s)^{\alpha-1}}{\Gamma(\alpha)}\,\mathrm{d}B_{s}$
where $\alpha>1$, $\beta>0$ and $x_{1},x_{2}\in\mathbb{R}$. Recall that by Example 2.6, the comparison principle holds if and only if $E_{\alpha,\beta}(-t^{\alpha})\geq 0$ for all $t\geq 0$. If $0<\beta<\alpha$ and $\alpha>1$, then $E_{\alpha,\beta}(-t^{\alpha})$ necessarily takes negative values and hence the comparison principle fails.
This example illustrates that to obtain a comparison principle, further conditions need to be imposed. In the above example, the drift b(x)=xb(x)=-x is non-increasing.
Theorem 3.2.
Suppose $K\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$ such that $K(0)\geq 0$ and $K^{\prime}\geq 0$ [^2] By abuse of terminology, we denote by $K$ its continuous representative and by $K^{\prime}$ its weak derivative.. Let $b_{1},b_{2},\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ be such that there exist constants $C>0$, $\eta\in[1/2,1]$ such that
$|b_{1}(x)-b_{1}(y)|+|b_{2}(x)-b_{2}(y)|\leq C|x-y|\quad\text{and}\quad|\sigma(x)-\sigma(y)|\leq C|x-y|^{\eta}$
hold for all $x,y\in\mathbb{R}$. Moreover, suppose that the following conditions hold:
  1. (i)
    the ordering condition $b_{1}(x)\leq b_{2}(x)$ holds for every $x\in\mathbb{R}$;
  2. (ii)
    at least one of the functions $b_{1}(\cdot),b_{2}(\cdot)$ is monotonically non-decreasing;
  3. (iii)
    Let $g_{1},g_{2}$ be given as in (4) with $\widetilde{g_{1}},\widetilde{g}_{2}\in C((0,\infty))$ such that $\widetilde{g}_{2}-\widetilde{g}_{1}\geq 0$ is non-decreasing on $(0,\infty)$, and $h_{1},h_{2}\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ satisfy $h_{1}\leq h_{2}$.
Then the unique strong solutions $X^{1},X^{2}$ of (1) with $g,b$ replaced by $g_{i},b_{i}$ with $i=1,2$ satisfy $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\;t>0]=1$.
Proof.
Let us first note that the equations of interest have weak solutions $X^{1},X^{2}$ with sample paths in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ due to [1, Theorem 1.2]. Moreover, it follows from [1, Theorem 6.1 (ii)] that $X^{1}-g_{1},X^{2}-g_{2}$ have a modification with continuous sample paths. Hence, $X^{1},X^{2}$ are continuous on $(0,\infty)$. Additionally, these solutions satisfy for any $T>0,i=1,2$, $\int_{0}^{T}\mathbb{E}[|X_{t}^{i}|^{2}]\,\mathrm{d}t<\infty$ c.f. Proposition A.1. The Yamada-Watanabe type of approach as given in [3, Proposition B.3] yields pathwise uniqueness and hence the Yamada-Watanabe-Engelbert theorem (see Kurtz [24, Theorem 1.5]) implies strong existence.
Without loss of generality, we may assume that $b_{1}$ is non-decreasing. Moreover, since $K$ is non-decreasing by assumption (iii), we get that
$g_{2}(t)-g_{1}(t)=\widetilde{g}_{2}(t)-\widetilde{g}_{1}(t)+\int_{0}^{t}K(t-s)(h_{2}(s)-h_{1}(s))\,\mathrm{d}s\geq 0$
is non-decreasing on $(0,\infty)$. Since $K\in W_{\mathrm{loc}}^{1,2}(\mathbb{R}_{+})$, the classical and stochastic version of Fubini’s theorem [32] and [12, Theorem 8.2], imply that
$X^{1}_{t}-X^{2}_{t}$ $=K(0)\int_{0}^{t}\sigma(X^{1}_{s})-\sigma(X^{2}_{s})\,\mathrm{d}B_{s}+\int_{0}^{t}\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{1}_{r})-b_{2}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\quad+\int_{0}^{t}\int_{0}^{s}K^{\prime}(s-r)\left(\sigma(X^{1}_{r})-\sigma(X^{2}_{r})\right)^{2}\mathrm{d}B_{r}\,\mathrm{d}s+K(0)\int_{0}^{t}b_{1}(X^{1}_{s})-b_{2}(X^{2}_{s})\,\mathrm{d}s$
$\quad+(g_{1}-g_{2})(t)$
which is a continuous semimartingale. Hence we may follow the same approximation procedure as presented in [23, Chapter VI, Theorem 1.1]. That is, let $(a_{n})_{n\in\mathbb{N}_{0}}\subset(0,1]$ be a sequence with $a_{0}=1$, $a_{n}<a_{n-1}$ and $\lim_{n\longrightarrow\infty}a_{n}=0$ such that $\int_{a_{n}}^{a_{n-1}}x^{-2\eta}\,\mathrm{d}x=n.$ Then, we can construct a sequence $(\varphi_{n})_{n\in\mathbb{N}}\subset C^{2}(\mathbb{R})$ of twice continuously differentiable functions such that
  • $\varphi_{n}^{\prime}(x)=0$ for $x\leq a_{k}$, $\varphi_{n}^{\prime}(x)=1$ for $x\geq a_{n-1}$ and $\varphi_{n}^{\prime}(x)\in[0,1]$ for $a_{n}<x<a_{n-1},$
  • $\varphi_{n}^{\prime\prime}(x)\in[0,\frac{2}{n}x^{-2\eta}]$ for $a_{n}<x<a_{n-1},$
and $\varphi_{n}(x)\nearrow x_{+}$ as $n\longrightarrow\infty$. Here, $x_{+}\coloneqq\max\{x,0\}$ denotes the positive part of $x$. Applying Itô ’s formula yields $\varphi_{n}(X^{1}_{t}-X^{2}_{t})=I_{1}+\dots+I_{6},$ where
$I_{1}$ $=K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})(\sigma(X^{1}_{s})-\sigma(X^{2}_{s}))\,\mathrm{d}B_{s},$
$I_{2}$ $=\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{1}_{r})-b_{2}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s,$
$I_{3}$ $=\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(\sigma(X^{1}_{r})-\sigma(X^{2}_{r}))\,\mathrm{d}B_{r}\,\mathrm{d}s,$
$I_{4}$ $=K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})(b_{1}(X^{1}_{s})-b_{2}(X^{2}_{s}))\,\mathrm{d}s,$
$I_{5}$ $=\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\,\mathrm{d}(g_{1}-g_{2})(s),$
$I_{6}$ $=\frac{K(0)^{2}}{2}\int_{0}^{t}\varphi_{n}^{\prime\prime}(X^{1}_{s}-X^{2}_{s})(\sigma(X^{1}_{s})-\sigma(X^{2}_{s}))^{2}\,\mathrm{d}s.$
It is a routine matter to show $\mathbb{E}[I_{1}]=\mathbb{E}[I_{3}]=0.$ Since $g_{2}-g_{1}$ is non-decreasing and $\varphi_{n}^{\prime}\geq 0,$ we deduce $I_{5}\leq 0.$ Moreover, by using the construction of $\varphi_{n},$
$I_{6}\lesssim\int_{0}^{t}\varphi_{n}^{\prime\prime}(X^{1}_{s}-X^{2}_{s})|X^{1}_{s}-X^{2}_{s}|^{2\eta}\,\mathrm{d}s\lesssim\frac{1}{n}.$
By using $K(0)>0$, $b_{1}(x)\leq b_{2}(x)$ and the Lipschitz continuity of $b_{1}$, we obtain
$I_{4}$ $=K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\mathbbm{1}_{\{X^{1}_{s}\geq X^{2}_{s}\}}(b_{1}(X^{1}_{s})-b_{1}(X^{2}_{s}))\,\mathrm{d}s$
$\quad+K(0)\int_{0}^{t}\varphi_{n}^{\prime}(X^{1}_{s}-X^{2}_{s})\mathbbm{1}_{\{X^{1}_{s}\geq X^{2}_{s}\}}(b_{1}(X^{2}_{s})-b_{2}(X^{2}_{s}))\,\mathrm{d}s$
$\lesssim K(0)\int_{0}^{t}(X^{1}_{s}-X^{2}_{s})_{+}\,\mathrm{d}s.$
In the same spirit and by using $K^{\prime}\geq 0$ and $b_{1}$ is non-decreasing, we obtain
$I_{2}$ $=\int_{0}^{t}\varphi^{\prime}_{n}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{1}_{r})-b_{1}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\quad+\int_{0}^{t}\varphi^{\prime}_{n}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)(b_{1}(X^{2}_{r})-b_{2}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\leq\int_{0}^{t}\varphi^{\prime}_{n}(X^{1}_{s}-X^{2}_{s})\int_{0}^{s}K^{\prime}(s-r)\mathbbm{1}_{\{X^{1}_{r}\geq X^{2}_{r}\}}(b_{1}(X^{1}_{r})-b_{1}(X^{2}_{r}))\,\mathrm{d}r\,\mathrm{d}s$
$\lesssim\int_{0}^{t}\int_{0}^{s}K^{\prime}(s-r)(X^{1}_{r}-X^{2}_{r})_{+}\,\mathrm{d}r\,\mathrm{d}s$
$=\int_{0}^{t}\int_{r}^{t}K^{\prime}(s-r)(X^{1}_{r}-X^{2}_{r})_{+}\,\mathrm{d}s\,\mathrm{d}r$
$=\int_{0}^{t}(K(t-r)-K(0))(X^{1}_{r}-X^{2}_{r})_{+}\,\mathrm{d}r.$
Moreover, $\mathbb{E}[\varphi_{n}(X^{1}_{t}-X^{2}_{t})]$ converges to $\mathbb{E}[(X^{1}_{t}-X^{2}_{t})_{+}]$ as $n\longrightarrow\infty$, thanks to the monotone convergence theorem. Consequently, we arrive at
$\mathbb{E}[(X^{1}_{t}-X^{2}_{t})_{+}]\lesssim\int_{0}^{t}K(t-s)\mathbb{E}\left[(X^{1}_{s}-X^{2}_{s})_{+}\right]\,\mathrm{d}s.$
Finally, Gronwall’s inequality for Volterra equations see e.g. [10, Lemma A.1] implies $\mathbb{E}[(X^{1}_{t}-X^{2}_{t})_{+}]=0$ and by the continuity of the sample paths we conclude $\mathbb{P}[X^{1}_{t}\leq X^{2}_{t},\,t>0]=1$. ∎
Remark 3.3.
The proof shows that the comparison principle also holds for the kernel $\widetilde{K}\coloneqq-K$ under the new assumption that $b_{1}(x)\geq b_{2}(x)$ and at least one of $b_{1}(\cdot),b_{2}(\cdot)$ is non-increasing for each $t\geq 0$.
The fractional kernel $K(t)=\frac{t^{\alpha-1}}{\Gamma(\alpha)}$ with $\alpha>3/2$ satisfies the assumptions stated in Theorem 3.2. The case $\alpha>1$ can now be obtained by approximation similar to Theorem 2.10. Since uniqueness is not guaranteed in such cases, we show the existence of a monotone coupling.
Theorem 3.4.
[…]
$\begin{pmatrix}\widehat{X}^{1}_{t}\\ \widehat{X}^{2}_{t}\end{pmatrix}=\begin{pmatrix}g_{1}(t)\\ g_{2}(t)\end{pmatrix}+\int_{0}^{t}K(t-s)\begin{pmatrix}b_{1}(\widehat{X}^{1}_{s})\\ b_{2}(\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}K(t-s)\begin{pmatrix}\sigma(\widehat{X}^{1}_{s})\\ \sigma(\widehat{X}^{2}_{s})\end{pmatrix}\mathrm{d}\widehat{B}_{s}$
such that $\widehat{\mathbb{P}}[\widehat{X}^{1}_{t}\leq\widehat{X}^{2}_{t},\,t>0]=1$.
Proof.
We choose $(b_{1,n})_{n\in\mathbb{N}}$, $(b_{2,n})_{n\in\mathbb{N}}$$(\sigma_{n})_{n\in\mathbb{N}}$ as in the proof of Theorem 2.10 and can assume that $b_{1,n}$ or $b_{2,n}$ is non-decreasing. Moreover, define $K_{n}(t)=K(t+\frac{1}{n}).$ Clearly, $K_{n}\in W^{1,2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and $K_{n}\longrightarrow K$ in $L^{2}_{\mathrm{loc}}(\mathbb{R}_{+}).$ Let $(X^{n})_{n\in\mathbb{N}}$, $(Y^{n})_{n\in\mathbb{N}}$ be the family of continuous processes constructed by $(g_{1},K_{n},b_{1,n},\sigma_{n})$ and $(g_{2},K_{n},b_{2,n},\sigma_{n})$ on some fixed filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. Then, by Theorem 3.2, $\mathbb{P}[X^{n}_{t}\leq Y^{n}_{t},\;t>0].$ Using Propsition A.2, we obtain for some $p\geq 2$ suitably chosen
$\sup_{n\in\mathbb{N}}\mathbb{E}[|(X^{n}_{t}-g_{1}(t))-(X^{n}_{s}-g_{2}(s))|^{p}]+\sup_{n\in\mathbb{N}}\mathbb{E}[|(Y^{n}_{t}-g_{1}(t))-(Y^{n}_{s}-g_{2}(s))|^{p}]\leq c|t-s|^{1+\theta},$
where $\theta>0$ and $c$ is a constant that only depends on $p,C,T,K_{1}$. Define for every $n\in\mathbb{N}$ the continuous semimartingale
$Z^{n}_{t}=\int_{0}^{t}\begin{pmatrix}b_{1,n}(X^{n}_{s})\\ b_{2,n}(Y^{n}_{s})\end{pmatrix}\mathrm{d}s+\int_{0}^{t}\begin{pmatrix}\sigma_{n}(X^{n}_{s})\\ \sigma_{n}(Y^{n}_{s})\end{pmatrix}\mathrm{d}B_{s}.$
Then we also obtain for some $p\geq 2$ suitably chosen
$\sup_{n\in\mathbb{N}}\mathbb{E}[|Z^{n}_{t}-Z^{n}_{s}|^{p}]\leq c|t-s|^{1+\theta},$
where $\theta>0$ and $c$ is a constant that only depends $C,T,K_{1}$ and $p$. Hence, we conclude by Kolmogorov’s tightness criterion that $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ is a tight sequence of continuous processes in $\mathbb{R}^{4}$ and so has a weakly convergent subsequence $((X^{n}-g_{1},Y^{n}-g_{2},Z^{n}))_{n\in\mathbb{N}}$ on the path space $C(\mathbb{R}_{+};\mathbb{R}^{4})$.
Define the continuous mapping $\Phi\colon C(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})\longrightarrow C_{\delta}(\mathbb{R}_{+};\mathbb{R}^{2})\times C(\mathbb{R}_{+};\mathbb{R}^{2})$ by
$\Phi(x,y,z_{1},z_{2})=(x+g_{1},y+g_{2},z_{1},z_{2}).$
[…]
We expect that the results presented in section 3 can be extended to time-dependent coefficients. However, establishing such an extension would require extending the results in [1] to time-dependent coefficients, which is beyond the scope of this paper, and leave such an extension to the interested reader.
Appendix A Existence and uniqueness for Lipschitz coefficients
Below we prove a result on the existence and uniqueness of solutions for a stochastic Volterra equation with Lipschitz continuous coefficients and a general driving force $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in[2,\infty]$. The following extends [10, Theorem 4.8] to time-dependent coefficients while, for the sake of simplicity, we stay in the finite-dimensional case.
Proposition A.1.
Let $K\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ and let $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ be measurable such that
$|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq C|x-y|,$
and
(16) $|b(t,x)|+|\sigma(t,x)|\leq C(1+|x|)$
hold for a.a. $t\geq 0$ and all $x,y\in\mathbb{R}$ with some constant $C>0$. Then for each $g\in L^{q}_{\mathrm{loc}}(\mathbb{R}_{+})$, $q\in[2,\infty]$, there exists a unique strong solution of the equation (1). This solution satisfies $X\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L_{\mathrm{loc}}^{q}(\mathbb{R}_{+}))$ for each $p\in[2,q]$ with $p<\infty$. To be more precise, $\|X\|_{L^{p}(\Omega;L^{q}([0,T]))}\leq c$, where $c>0$ is a constant that only depends on K,T,p,CK,T,p,C [^3] The dependence on $C$ only comes through the growth estimate (16)..
Proof.
Below we follow the steps outlined in [10, Theorem 4.8]. Fix $\lambda<0$ and define the functions $K_{\lambda}(t)=\mathrm{e}^{\lambda t}K(t)$ and $g_{\lambda}(t)=\mathrm{e}^{\lambda t}g(t)$. Let $X_{t}(g)$ be a solution of (1). Then $Y^{\lambda}_{t}(g)=\mathrm{e}^{\lambda t}X_{t}(g)$ satisfies
(17) $Y_{t}^{\lambda}(g)=g_{\lambda}(t)+\int_{0}^{t}K_{\lambda}(t-s)b_{\lambda}(s,Y_{s}^{\lambda}(g))\,\mathrm{d}s+\int_{0}^{t}K_{\lambda}(t-s)\sigma_{\lambda}(s,Y_{s}^{\lambda}(g))\,\mathrm{d}B_{s},$
where $b_{\lambda}(s,x)=\mathrm{e}^{\lambda s}b(s,\mathrm{e}^{-\lambda s}x)$ and $\sigma_{\lambda}(s,x)=\mathrm{e}^{\lambda s}\sigma(s,\mathrm{e}^{-\lambda s}x)$. Conversely, let $Y_{\lambda}$ by a solution of (17), then $X^{\lambda}_{t}(g)=\mathrm{e}^{-\lambda t}Y_{\lambda}(t;g)$ satisfies (1). Therefore, it suffices to prove existence and uniqueness for (17). We solve (17) by a fixed point argument for the case $p=q$.
Fix $T>0$, then $g|_{[0,T]}\in L^{q}([0,T]).$ For a given $\lambda<0$, define
$\mathcal{T}_{\lambda}(X;g)(t)\coloneqq g_{\lambda}(t)+\int_{0}^{t}K_{\lambda}(t-s)b_{\lambda}(s,X_{s})\,\mathrm{d}s+\int_{0}^{t}K_{\lambda}(t-s)\sigma_{\lambda}(s,X_{s})\,\mathrm{d}B_{s}.$
Then (17) is equivalent to $X^{\lambda}=\mathcal{T}_{\lambda}(X^{\lambda}(g))$. Below we show that $\mathcal{T_{\lambda}}(\cdot,g)$ is a contraction on $L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$ when $\lambda<0$ is small enough. For brevity, we let $\|\cdot\|_{p}=\|\cdot\|_{L^{p}(\Omega;L^{p}([0,T])}$. Take $X\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$, then
$\|\mathcal{T}_{\lambda}(X;g)\|_{p}\leq\|g\|_{L^{p}([0,T])}+\|K_{\lambda}\ast b_{\lambda}(\cdot,X)\|_{p}+\|K_{\lambda}\ast\sigma_{\lambda}(\cdot,X)\,\mathrm{d}B\|_{p}.$
For the drift we find, using Young’s inequality
$\|K_{\lambda}\ast b_{\lambda}(\cdot,X)\|_{p}$ $=\left\|\|K_{\lambda}\ast b_{\lambda}(\cdot,X)\|_{L^{p}([0,T])}\right\|_{L^{p}(\Omega)}$
$\leq\|K_{\lambda}\|_{L^{1}([0,T])}\,\left\|\|b_{\lambda}(\cdot,X_{s})\|_{L^{p}([0,T])}\right\|_{L^{p}(\Omega)}$
$\leq C\|K_{\lambda}\|_{L^{1}([0,T])}\|1+X\|_{p}.$
For the stochastic convolution, we find using BDG and Young’s inequality
$\|K_{\lambda}\ast\sigma_{\lambda}(\cdot,X)\,\mathrm{d}B\|_{p}$ $\leq c_{p}\left(\mathbb{E}\left[\|K_{\lambda}^{2}\ast\sigma_{\lambda}(\cdot,X)^{2}\|_{L^{p/2}([0,T])}^{p/2}\right]\right)^{1/p}$
$\leq c_{p}\|K^{2}_{\lambda}\|_{L^{1}([0,T])}^{1/2}\left(\mathbb{E}\left[\|\sigma_{\lambda}(\cdot,X)^{2}\|_{L^{p/2}([0,T])}^{p/2}\right]\right)^{1/p}$
$\leq c_{p}\|K_{\lambda}\|_{L^{2}([0,T])}\left(\mathbb{E}\left[\|C(1+|X|)\|_{L^{p}([0,T])}^{p}\right]\right)^{1/p}$
$\leq Cc_{p}\|K_{\lambda}\|_{L^{2}([0,T])}\|1+X\|_{p}.$
Hence, $\mathcal{T}_{\lambda}(\cdot;g)$ leaves the space $L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$ invariant. In the same way, if $X,Y\in L^{p}(\Omega,\mathcal{F},\mathbb{P};L^{p}([0,T]))$, we obtain
$\|\mathcal{T}_{\lambda}(X;g)-\mathcal{T}_{\lambda}(Y;g)\|_{p}$ $\leq\|K_{\lambda}\ast(b_{\lambda}(\cdot,X)-b_{\lambda}(\cdot,Y))\|_{p}+\|K_{\lambda}\ast(\sigma_{\lambda}(\cdot,X)-\sigma_{\lambda}(\cdot,Y))\|_{p}$
$\leq C\max\{\|K_{\lambda}\|_{L^{1}([0,T])},c_{p}\|K_{\lambda}\|_{L^{2}([0,T])}\}\|X-Y\|_{p}.$
Since by dominated convergence $\|K_{\lambda}\|_{L^{1}([0,T])},\ \|K_{\lambda}\|_{L^{2}([0,T])}\longrightarrow 0$ as $\lambda\longrightarrow$-$\infty$, we can choose $\lambda<0$ sufficiently small such that $\mathcal{T}_{\lambda}(\cdot;g)$ is a contraction and, hence, has a unique fixed point $Y_{\lambda}(\cdot;g)$ which is the unique solution of (17). The same fixpoint argument can be executed in the space $L^{p}(\Omega.\mathcal{F},\mathbb{P};L^{\infty}([0,T]))$, provided that $g\in L_{\mathrm{loc}}^{\infty}(\mathbb{R}_{+})$. Using similar bounds as above, it is easy to derive, for $p\in[2,q]$ and a.a. $t\geq 0$
$\mathbb{E}[|X_{t}|^{p}]\leq|g(t)|^{p}+A_{0}+A_{1}\int_{0}^{t}K(t-s)^{2}\mathbb{E}[|X_{s}|^{p}]\,\mathrm{d}s,$
where $A_{0},A_{1}$ are constants that only depends on $K,C,T,p$. Using the Volterra type Gronwall inequality (see e.g. [10, Lemma A.1] yields for a.a. $t\geq 0$
(18) $\mathbb{E}[|X_{t}|^{p}]\leq|g(t)|^{p}+A_{0}+\int_{0}^{t}R(t-s)(|g(s)|^{p}+A_{0})\,\mathrm{d}s,$
where $R\in L^{1}([0,T])$ denotes the resolvent of the second kind of $A_{1}K^{2}$. Since $|g|^{p}\in L^{q/p}([0,T])$, the right-hand side of (18) belongs to $L^{q/p}([0,T]).$ Thus, $\|X\|_{L^{p}(\Omega;L^{q}([0,T]))}\leq c$, where $c$ is a constant that only depends on $K,C,T,p$. ∎
Proposition A.2.
Suppose that $(g,K,b,\sigma)$ satisfies Assumption 2.1. Let $X$ be any solution of (1). Then the following assertions hold:
  1. (i)
    Under Assumption 2.1.(i), XgX-g has, for each $\theta<\gamma$, a continuous modification with locally $\theta$-Hölder continuous sample paths.
  2. (ii)
    Under Assumption 2.1.(ii), XgX-g has, for each $\theta<\vartheta$ with
    $\vartheta\coloneqq\gamma+\frac{1}{2}\frac{\eta}{2+\eta}-\frac{2\xi}{q}\left(1+\frac{1}{\eta}\right),$
    a continuous modification with locally $\theta$-Hölder continuous sample paths.
Proof.
Firstly, let us consider the case (ii). Since $q\geq 2\xi\left(1+\frac{2}{\eta}\right)$, we find $p\coloneqq\frac{\eta q}{(2+\eta)\xi}\geq 2$. Then, for $0<s<t\leq T,$
$|(X_{t}-g(t))-(X_{s}-g(s))|^{p}$
$\quad\lesssim\left|\int_{s}^{t}K(t-r)b(r,X_{r})\,\mathrm{d}r\right|^{p}+\left|\int_{0}^{s}(K(t-r)-K(s-r))b(r,X_{r})\,\mathrm{d}r\right|^{p}$
$\qquad+\left|\int_{s}^{t}K(t-r)\sigma(r,X_{r})\,\mathrm{d}B_{r}\right|^{p}+\left|\int_{0}^{s}(K(t-r)-K(s-r))\sigma(r,X_{r})\,\mathrm{d}B_{r}\right|^{p}$
$\quad\eqqcolon I_{1}+I_{2}+I_{3}+I_{4}$
An application of Jensen’s and Hölder’s inequality yields
(19) $\mathbb{E}[I_{1}]$ $\leq\left(\int_{s}^{t}|K(t-r)|\,\mathrm{d}r\right)^{p-1}\mathbb{E}\left[\int_{s}^{t}|K(t-r)||b(r,X_{r})|^{p}\,\mathrm{d}r\right]$
$\leq C^{p}\|K\|_{L^{1}([0,t-s])}^{p-1}\mathbb{E}\left[\int_{s}^{t}|K(t-r)|(1+|X_{r}|)^{\xi p}\,\mathrm{d}r\right]$
$\leq C^{p}\|K\|_{L^{2+\eta}([0,t-s])}^{p-1}(t-s)^{\frac{1+\eta}{2+\eta}(p-1)}\|K\|_{L^{2+\eta}([0,t-s])}\mathbb{E}\left[\left(\int_{s}^{t}(1+|X_{r}|)^{\xi p\frac{2+\eta}{1+\eta}}\,\mathrm{d}r\right)^{\frac{1+\eta}{2+\eta}}\right]$
$\leq C^{p}c(T,\eta)^{p}(t-s)^{\gamma p+\frac{1+\eta}{2+\eta}(p-1)}\int_{0}^{T}\mathbb{E}\left[(1+|X_{r}|)^{q}\right]\mathrm{d}r$
where the finiteness of the expectation follows from Proposition A.1. Analogously, we obtain from Itô’s isometry, Jensen’s and Hölder’s inequality
$\mathbb{E}[I_{3}]$ $=\mathbb{E}\left[\left(\int_{s}^{t}K(t-r)^{2}\sigma(r,X_{r})^{2}\,\mathrm{d}r\right)^{\frac{p}{2}}\right]$
(20) $\leq C^{p}\left(\int_{s}^{t}K(t-r)^{2}\,\mathrm{d}r\right)^{\frac{p}{2}-1}\mathbb{E}\left[\int_{s}^{t}K(t-r)^{2}(1+|X_{r}|)^{\xi p}\,\mathrm{d}r\right]$
$\leq C^{p}\|K\|_{L^{2+\eta}([0,t-s])}^{p-2}(t-s)^{\frac{\eta}{2+\eta}\left(\frac{p}{2}-1\right)}\|K\|_{L^{2+\eta}([0,t-s])}^{2}\mathbb{E}\left[\left(\int_{s}^{t}(1+|X_{r}|)^{\xi p\frac{2+\eta}{\eta}}\,\mathrm{d}r\right)^{\frac{2}{2+\eta}}\right]$
$\leq C^{p}c(T,\eta)^{p}(t-s)^{\gamma p+\frac{\eta}{2+\eta}\left(\frac{p}{2}-1\right)}\int_{0}^{T}\mathbb{E}\left[(1+|X_{r}|)^{q}\right]$
The same estimate can be derived for $\mathbb{E}[I_{2}],\mathbb{E}[I_{3}].$ Thus, using $p\geq 2$, gives $(1+\eta)(p-1)\geq\eta\left(\frac{p}{2}-1\right)$ and hence we obtain
$\left(\mathbb{E}[|(X_{t}-g(t))-(X_{s}-g(s))|^{p}]\right)^{\frac{1}{p}}\leq c(t-s)^{\frac{1}{p}+\gamma+\frac{\eta}{2+\eta}\left(\frac{1}{2}-\frac{1}{p}\right)-\frac{1}{p}}=c(t-s)^{\frac{1}{p}+\vartheta},$
[…]
$q>2\xi\frac{1+\frac{1}{\eta}}{\gamma+\frac{1}{2}\frac{\eta}{2+\eta}}$
which is satisfied by Assumption 2.1. The Hölder continuity now follows from the Kolmogorov-Chentsov theorem.
For the case of Assumption 2.1.(i), we let $p\geq 2$ be arbitrary. By following similar arguments to (19) and (20) and using $\operatorname*{ess\,sup}_{r\in[0,T]}\mathbb{E}[|X_{r}|^{p}]<\infty$ due to Proposition A.1, we find
$\mathbb{E}[|(X_{t}-g(t))-(X_{s}-g(s))|^{p}]\lesssim(t-s)^{p\gamma}.$
This proves the assertion. ∎
If $g\in L^{\infty}_{\mathrm{loc}}(\mathbb{R}_{+})$, then the above result essentially coincides with [4]. For the case of Assumption 2.1.(ii), it gives a new way of proving the existence of continuous solutions.
Appendix B Splitting Method for Stochastic Volterra Equations
Given any solution $X$ of (1) where $g\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})\cap C((0,\infty))$ and $b,\sigma\colon\mathbb{R}_{+}\times\mathbb{R}\longrightarrow\mathbb{R}$ are measurable such that there exists $C\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})$ with
(21) $|b(t,x)-b(t,y)|+|\sigma(t,x)-\sigma(t,y)|\leq C(t)|x-y|,\quad|b(t,x)|^{2}+|\sigma(t,x)|^{2}\leq C^{2}(t)(1+|x|^{2}),$
for a.a. $t\geq 0$ and $x,y\in\mathbb{R}$. Let $T>0$, $N\in\mathbb{N}$ be fixed and $t_{k}=kT/N,$ $k\in\{0,1,\dots,N\},$ be an equidistant grid of $[0,T]$ with step size $T/N.$ Let us denote by $(\widehat{X}_{t})_{t\in(0,T]}$ its approximation
$\widehat{X}_{t}$ $=g(t)+\sum_{k=1}^{N}\mathbbm{1}_{[t_{k},T]}(t)\left(K(t-t_{k})\int_{t_{k-1}}^{t_{k}}[b(s,\xi^{k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{k}_{s})\,\mathrm{d}B_{s}]\right)$
(22) $=g(t)+\int_{0}^{t_{\eta(t)}}K(t-t_{\eta(s)})[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}],$
where $\eta(t)=k$ if $t\in[t_{k},t_{k+1})$ and $\xi_{t}\coloneqq\sum_{k=1}^{N}\mathbbm{1}_{[t_{k-1},t_{k})}(t)\xi_{t}^{k}$ with
$\xi^{k}_{t}=\widehat{X}_{t_{k}-}+\int_{t_{k-1}}^{t}K(0_{+})[b(s,\xi^{k}_{s})\,\mathrm{d}s+\sigma(s,\xi^{k}_{s})\,\mathrm{d}B_{s}],\quad t\in[t_{k-1},t_{k}).$
Denote by
$\omega_{f,T}(\delta)\coloneqq\sup_{s,t\in[0,T]:|s-t|\leq\delta}|f(s)-f(t)|$
the $\delta$-modulus of continuity of a function $f\colon[0,T]\longrightarrow\mathbb{R}$ and by $\|f\|_{[0,T]}\coloneqq\sup_{t\in[0,T]}|f(t)|$ its sup-norm. Below we extend the approximation [7, Lemma 3.1, Proposition 3.1] to the case of time-dependent coefficients and possibly singular $g$. As a first step, we prove a refined bound on the approximation where $g$ is regular.
Proposition B.1.
Suppose that $K$ is nonnegative, nonincreasing and continuous on $\mathbb{R}_{+}$, and that $g\in C(\mathbb{R}_{+})$ Then, for every $T>0$, it holds that
(23) $\sup_{t\in[0,T]}\mathbb{E}[|\widehat{X}_{t}|^{2}+|\xi_{t}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}$
and
$\sup_{t\in[0,T]}\left(\mathbb{E}[|\widehat{X}_{t}-\xi_{t}|^{2}]+\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]\right)\\ \lesssim\omega^{2}_{g,T}\left(\frac{T}{N}\right)+\left(1+\|g\|_{[0,T]}^{2}\right)\left(\omega^{2}_{K,T}\left(\frac{T}{N}\right)+\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s\right).$
Proof.
Let $k\in\{1,\dots,N\}$ and $t\in[t_{k-1},t_{k})$. By using (2),
$|\xi_{t}|^{2}\lesssim|\widehat{X}_{t_{k}-}|^{2}+\left|\int_{t_{k-1}}^{t}b(s,\xi_{s})\,\mathrm{d}s\right|^{2}+\left|\int_{t_{k-1}}^{t}\sigma(s,\xi_{s})\,\mathrm{d}B_{s}\right|^{2}.$
An application of Jensen’s inequality, Itô ’s isometry and the growth condition on $b,\sigma$ yields
(24) $\mathbb{E}[|\xi_{t}|^{2}]\lesssim\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]+\int_{t_{k-1}}^{t}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s.$
Thus by Gronwall’s inequality, $\mathbb{E}[|\xi_{t}|^{2}]\lesssim\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]+\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s$ for $t\in[t_{k-1},t_{k})$. By letting $t\nearrow t_{k}$ in (22) and using $\lim_{t\nearrow t_{k}}\eta(t)=k-1$, we obtain
$\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]$ $\lesssim g(t_{k})^{2}+\int_{0}^{t_{k-1}}K(t-t_{\eta(s)})^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$
$\lesssim\|g\|_{[0,T]}^{2}+\int_{0}^{t_{k-1}}C(s)^{2}\,\mathrm{d}s+\sum_{j=1}^{k-1}\int_{t_{j-1}}^{t_{j}}\left(\mathbb{E}[|\widehat{X}_{t_{j}-}|^{2}]+\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s\right)C(s)^{2}\,\mathrm{d}s$
$\lesssim 1+\|g\|_{[0,T]}^{2}+\sum_{j=1}^{k-1}\mathbb{E}[|\widehat{X}_{t_{j}-}|^{2}]\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s$
where we have used $\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s\leq\|C\|_{L^{2}([0,T]}^{2}$. By virtue of the discrete Gronwall inequality [13] and the standard inequality $1+x\leq\mathrm{e}^{x}$, we obtain $\sup_{1\leq k\leq N}\mathbb{E}[|\widehat{X}_{t_{k}-}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}<\infty.$ Consequently, $\sup_{t\in[0,T]}\mathbb{E}[|\xi_{t}|^{2}]\lesssim 1+\|g\|_{[0,T]}^{2}$ and so by construction of the approximation we arrive at (23).
Let $t\in[t_{k-1},t_{k})$. We have $|\widehat{X}_{t}-\xi_{t}|^{2}\lesssim|\widehat{X}_{t}-\widehat{X}_{t_{k}-}|^{2}+|\widehat{X}_{t_{k}-}-\xi_{t}|^{2}$. For the second term, using (24) we arrive at
$\mathbb{E}[|\widehat{X}_{t_{k}-}-\xi_{t}|^{2}]\lesssim\int_{t_{k-1}}^{t}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s\lesssim\left(1+\|g\|_{[0,T]}^{2}\right)\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s.$
For the first difference, we have
$\widehat{X}_{t}-\widehat{X}_{t_{k-}}=g(t)-g(t_{k})+\int_{0}^{t_{k-1}}(K(t-t_{\eta(s)})-K(t_{k}-t_{\eta(s)}))[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}]$
and consequently
$\mathbb{E}[|\widehat{X}_{t}-\widehat{X}_{t_{k}-}|^{2}]$
$\quad\lesssim|g(t)-g(t_{k})|^{2}+\int_{0}^{t_{k-1}}(K(t-t_{\eta(s)})-K(t_{k}-t_{\eta(s)}))^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{t}|^{2}])\,\mathrm{d}s$
$\quad\lesssim\omega^{2}_{g,T}\left(\frac{T}{N}\right)+\left(1+\|g\|_{[0,T]}^{2}\right)\omega^{2}_{K,T}\left(\frac{T}{N}\right).$
Finally, it remains to bound the difference
$\widehat{X}_{t}-X_{t}$ $=\int_{0}^{t}K(t-s)[b(s,\xi_{s})-b(s,X_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})-\sigma(s,X_{s})\,\mathrm{d}B_{s}]$
$\quad+\int_{0}^{t_{\eta(t)}}(K(t-t_{\eta(s)})-K(t-s))[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}]$
$\quad-\int_{t_{\eta(t)}}^{t}K(t-s)[b(s,\xi_{s})\,\mathrm{d}s+\sigma(s,\xi_{s})\,\mathrm{d}B_{s}].$
Using the linear growth and Lipschitz condition (21), we obtain
$\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]$ $\lesssim\int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\xi_{s}-X_{s}|^{2}]\,\mathrm{d}s$
$\quad+\int_{0}^{t_{\eta(t)}}(K(t-t_{\eta(s)})-K(t-s))^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$
$\quad+\int_{t_{\eta(t)}}^{t}K(t-s)^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s.$
For the first term we use $|\xi_{s}-X_{s}|^{2}\lesssim|\xi_{s}-\widehat{X}_{s}|^{2}+|\widehat{X}_{s}-X_{s}|^{2}$ to bound
$\ \int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\xi_{s}-X_{s}|^{2}]\,\mathrm{d}s$
$\lesssim\int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\xi_{s}-\widehat{X}_{s}|^{2}]\,\mathrm{d}s+\int_{0}^{t}K(t-s)^{2}C(s)^{2}\mathbb{E}[|\widehat{X}_{s}-X_{s}|^{2}]\,\mathrm{d}s$
$\lesssim\omega^{2}_{g,T}\left(\frac{T}{N}\right)+\left(1+\|g\|_{[0,T]}^{2}\right)\left(\omega^{2}_{K,T}\left(\frac{T}{N}\right)+\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s\right)+\int_{0}^{t}C(s)^{2}\mathbb{E}[|\widehat{X}_{s}-X_{s}|^{2}]\,\mathrm{d}s.$
For the second term, we use (24) to find that
$\int_{0}^{t_{\eta(t)}}(K(t-t_{\eta(s)})-K(t-s))^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$ $\lesssim\left(1+\|g\|_{[0,T]}^{2}\right)\omega^{2}_{K,T}\left(\frac{T}{N}\right).$
Finally, the last term can be bounded by
$\int_{t_{\eta(t)}}^{t}K(t-s)^{2}C(s)^{2}(1+\mathbb{E}[|\xi_{s}|^{2}])\,\mathrm{d}s$ $\lesssim\left(1+\|g\|_{[0,T]}^{2}\right)\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s.$
Hence, collecting all estimates, Gronwall’s inequality yields the assertion. ∎
Next, we prove the convergence of the approximation also in the case where $g$ may be singular in $t=0$, provided that the singularity is not too rough.
Proposition B.2.
Let $g\in L^{2}_{\mathrm{loc}}(\mathbb{R}_{+})\cap C((0,\infty))$ be such that there exists $\delta>0$ with
$\int_{0}^{T}\overline{g}(s)^{2+\delta}C(s)^{2}\,\mathrm{d}s<\infty$
where $\overline{g}(s)\coloneqq\sup_{s\leq t\leq T}|g(t)|$. Then, for every $t\in(0,T]$,
$\lim_{N\longrightarrow\infty}\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]=0.$
Proof.
Fix $\varepsilon>0$ and define $g_{\varepsilon}\colon\mathbb{R}_{+}\longrightarrow\mathbb{R}$ by $g_{\varepsilon}(t)=g(t+\varepsilon).$ Moreover, let $X^{\varepsilon}$ be the strong solution of
$X^{\varepsilon}_{t}=g_{\varepsilon}(t)+\int_{0}^{t}K(t-s)[b(s,X^{\varepsilon}_{s})\,\mathrm{d}s+\sigma(s,X^{\varepsilon}_{s})\,\mathrm{d}B_{s}].$
Denote by $\widehat{X}^{\varepsilon}$ its approximation and by $\xi^{\varepsilon}$ the process that arises in its construction. Then $|\widehat{X}_{t}-X_{t}|\leq|\widehat{X}_{t}-\widehat{X}^{\varepsilon}_{t}|+|\widehat{X}^{\varepsilon}_{t}-X^{\varepsilon}_{t}|+|X^{\varepsilon}_{t}-X_{t}|$. Proposition B.1 gives
$\mathbb{E}[|\widehat{X}^{\varepsilon}_{t}-X^{\varepsilon}_{t}|^{2}]\lesssim\omega^{2}_{g_{\varepsilon},T}\left(\frac{T}{N}\right)+\left(1+\|g_{\varepsilon}\|_{[0,T]}^{2}\right)\left(\omega^{2}_{K,T}\left(\frac{T}{N}\right)+\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s\right),$
and it is a routine to show $\mathbb{E}[|X^{\varepsilon}_{t}-X_{t}|^{2}]\lesssim|g_{\varepsilon}(t)-g(t)|^{2}.$ Furthermore, we claim that
(25) $\mathbb{E}[|\widehat{X}_{t}-\widehat{X}^{\varepsilon}_{t}|^{2}]\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{\eta(t)}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s,\qquad t\in(0,T].$
Indeed, the claim is certainly true on $(0,t_{1})$ due to $\widehat{X}_{t}^{\varepsilon}=g_{\varepsilon}(t)$ and $\widehat{X}_{t}=g(t)$. Suppose (25) were true on $(0,t_{k})$ and let $t\in[t_{k},t_{k+1})$ then using (21) we arrive at
$\mathbb{E}[|\widehat{X}_{t}-\widehat{X}^{\varepsilon}_{t}|^{2}]$
$\quad\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{k}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\mathbb{E}[|\xi_{s}-\xi^{\varepsilon}_{s}|^{2}]\,\mathrm{d}s$
$\quad\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{k-1}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s+\int_{t_{k-1}}^{t_{k}}C(s)^{2}\mathbb{E}[|\xi_{s}-\xi^{\varepsilon}_{s}|^{2}]\,\mathrm{d}s.$
Finally, an application of Gronwall’s inequality to estimate $\mathbb{E}[|\xi_{s}-\xi^{\varepsilon}_{s}|^{2}]\lesssim|g(t_{k})-g_{\varepsilon}(t_{k})|^{2}$ yields (25). Consequently, the following less sharp version of (25) holds true
$\mathbb{E}[|\widehat{X}_{t}-\widehat{X}_{t}^{\varepsilon}|^{2}]$ $\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\sum_{j=1}^{N}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}\int_{t_{j-1}}^{t_{j}}C(s)^{2}\,\mathrm{d}s$
$=|g(t)-g_{\varepsilon}(t)|^{2}+\int_{0}^{T}G^{N}_{\varepsilon}(s)C^{2}(s)\,\mathrm{d}s,$
where $G^{N}_{\varepsilon}(s)\coloneqq\sum_{j=1}^{N}\mathbbm{1}_{[t_{j-1},t_{j})}(s)|g(t_{j})-g_{\varepsilon}(t_{j})|^{2}$. By continuity of $g$ on $(0,T]$, one can show that $G^{N}_{\varepsilon}(t)\overset{N\longrightarrow\infty}{\longrightarrow}|g(t)-g_{\varepsilon}(t)|^{2}$ for all $t\in(0,T]$ and fixed $\varepsilon>0$. Moreover, it holds that
$\int_{0}^{T}|G^{N}_{\varepsilon}(s)|^{1+\delta/2}C(s)^{2}\,\mathrm{d}s$ $=\sum_{j=1}^{N}\int_{t_{j-1}}^{t_{j}}|g(t_{j})-g_{\varepsilon}(t_{j})|^{2+\delta}C(s)^{2}\,\mathrm{d}s$
$\lesssim\sum_{j=1}^{N}\int_{t_{j-1}}^{t_{j}}|g(t_{j})|^{2+\delta}C(s)^{2}\,\mathrm{d}s+\|g_{\varepsilon}\|_{[0,T]}^{2+\delta}\int_{0}^{T}C(s)^{2}\,\mathrm{d}s$
$\leq\int_{0}^{T}\overline{g}(s)^{2+\delta}C(s)^{2}\,\mathrm{d}s+\|g_{\varepsilon}\|_{[0,T]}^{2+\delta}\int_{0}^{T}C(s)^{2}\,\mathrm{d}s.$
Thus, $(G^{N}_{\varepsilon})_{N\in\mathbb{N}}$ is uniformly integrable. Moreover,
$\lim_{N\longrightarrow\infty}\sup_{1\leq k\leq N}\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s=0.$
Indeed, $t\longmapsto\int_{0}^{t}C(s)^{2}\,\mathrm{d}s$ is an absolutely continuous function. Thus, for every $\widetilde{\varepsilon}>0$ there exists $\widetilde{\delta}>0$ such that for any $a,b>0$ with $b-a<\widetilde{\delta}$ holds $\int_{a}^{b}C(s)^{2}\,\mathrm{d}s<\widetilde{\varepsilon}$. Choosing $N$ large enough yields $t_{k}-t_{k-1}=\frac{1}{N}<\widetilde{\delta}$ and so $\int_{t_{k-1}}^{t_{k}}C(s)^{2}\,\mathrm{d}s<\widetilde{\varepsilon}$, $k\in\{1,\dots,N\}$. Collecting all estimates gives
$\limsup_{N\longrightarrow\infty}\mathbb{E}[|\widehat{X}_{t}-X_{t}|^{2}]\lesssim|g(t)-g_{\varepsilon}(t)|^{2}+\int_{0}^{T}|g(s)-g_{\varepsilon}(s)|^{2}C(s)^{2}\,\mathrm{d}s.$
The right-hand side tends to zero as $\varepsilon\downarrow 0$ due to dominated convergence, which completes the proof. ∎
Finally, we prove an auxiliary result for kernels that preserve positivity.
Lemma B.3.
Suppose $K\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ is a non-negativity preserving kernel with $K(0_{+})>0$ and let $f\colon(0,\infty)\longrightarrow\mathbb{R}_{+}$ be non-decreasing. Let $N\in\mathbb{N}$, $0<t_{1}<\dots<t_{N}$ and $x_{1},\dots,x_{N}\in\mathbb{R}$ be such that, for any $k\in\{1,\dots,N\},$
(26) $f(t_{k})+\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})\geq 0.$
Then, it follows
$f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}x_{k}K(t-t_{k})\geq 0,\quad t>0.$
Proof.
We define recursively $\widetilde{x}_{1}\coloneqq\frac{-f(t_{1})}{K(0_{+})}$ and, for $k\in\{1,\dots,N\},$
$\widetilde{x}_{k}\coloneqq\frac{-1}{K(0_{+})}\left(f(t_{k})+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}K(t_{k}-t_{\ell})\right).$
By construction we have for any $k\in\{1,\dots,N\}$
(27) $f(t_{k})+\sum_{\ell=1}^{k}\widetilde{x}_{\ell}K(t_{k}-t_{\ell})=0.$
Moreover, define $\delta_{k}\coloneqq x_{k}-\widetilde{x}_{k},\,k\in\{1,\dots,N\}.$ Then, we have by (27) and (26)
$\sum_{\ell=1}^{k}\delta_{\ell}K(t_{k}-t_{\ell})=\sum_{\ell=1}^{k}x_{\ell}K(t_{k}-t_{\ell})+f(t_{k})\geq 0,\quad k\in\{1,\dots,N\},$
and since $K$ preserves non-negativity, we obtain $\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\delta_{k}K(t-t_{k})\geq 0$ for $t>0$. We have
$f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}x_{k}K(t-t_{k})=f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\widetilde{x}_{k}K(t-t_{k})+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\delta_{k}K(t-t_{k})$
and hence, it suffices to show that
(28) $f(t)+\sum_{k=1}^{N}\mathbbm{1}_{\{t_{k}\leq t\}}\widetilde{x}_{k}K(t-t_{k})\geq 0,\quad t>0.$
By induction over $k$ we show that (28) holds on $t_{k},t_{k+1})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k+1}\leq 0$. For $t\in(0,t_{1})$, (28) is evident and by assumption $\widetilde{x}_{1}=\frac{-f(t_{1})}{K(0_{+})}\leq 0$. Suppose (28) holds on $t_{k-1},t_{k})$ and $\widetilde{x}_{1},\dots,\widetilde{x}_{k}\leq 0.$ Then, we find for $t\in[t_{k},t_{k+1})$
$f(t)+\sum_{\ell=1}^{k}\widetilde{x}_{\ell}K(t-t_{\ell})$ $\geq f(t)+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}K(t-t_{\ell})-\left(f(t_{k})+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}K(t_{k}-t_{\ell})\right)$
$=f(t)-f(t_{k})+\sum_{\ell=1}^{k-1}\widetilde{x}_{\ell}\left(K(t-t_{\ell})-K(t_{k}-t_{\ell})\right)$
$\geq f(t)-f(t_{k})$
$\geq 0,$
where we used that $K(t-t_{\ell})-K(t_{k}-t_{\ell})\leq 0$ since $K\geq 0$ is non-increasing. Replacing $t$ by $t_{k+1}$ in the above calculation yields $\widetilde{x}_{k+1}\leq 0.$